### v3

# Codes Correcting Two Bursts of Exactly $b$ Deletions

Abstract
In this paper, we study codes that correct two bursts of deletions, with each burst having length exactly $b$, where $b>1$. Previously, the best known construction, derived using the syndrome compression technique, achieved at most $7\log n+O\left(\log n/\log\log n\right)$ redundant bits. In this work, we develop new idea to construct $q$-ary codes that achieve redundancy at most $6\log n+O(\log\log n)$ for all $b>1$ and $q\geq 2$. At last, for $b=1$, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
Index Terms:
deletion, burst-deletion, error-correcting codes, DNA-based storage
I Introduction
A subset $\mathcal{C}\subseteq\{0,1,\ldots,q-1\}^{n}$ (where $q\geq 2$) is called a $t$-deletion correcting code, if it has the property that if a codeword $\mathbf{x}\in\mathcal{C}$ is corrupted by deleting $t$ symbols to obtain a subsequence $\mathbf{y}\in\{0,1,\ldots,q-1\}^{n-t}$, then one can recover $\mathbf{x}$ from $\mathbf{y}$. The study of deletion correcting codes has a long history, dating back to at least the 1960s [1]. The seminal work in this field is [2], where by proposing a linear-time decoding algorithm, Levenshtein proved that the binary code (Varshamov-Tenengolts code, or VT code for short) constructed in [3] can combat a single deletion error. In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4]. For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|C\right|\right)$ [^1] All logarithms in this paper are to the base $2$, is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lowere bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6]. This implies that codes in [3, 4] have redundancy optimal up to a constant.
Due to applications in DNA-based data storage [7, 8, 9], document exchange [10, 11], multiple-deletion correcting codes with low redundancy have attracted a lot of interest in recent years [10, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19]. For general $t\geq 3$, the smallest redundancy, which is $(4t-1)\log n+o(\log n)$, was achieved by a construction given in [20]. For non-binary alphabets, Sima et al [15] presented a family of $q$-ary $t$-deletion correcting codes with redundancy $4t\log n+o(\log n)$ by using the syndrome compression technique. The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding. A straightforward application of this method will give a $q$-ary $t$-deletion correcting codes with redundancy $(4t-1)\log n+o(\log n)$, which is the best known result in redundancy. When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary $2$-deletion correcting codes with redundancy $5\log(n)+O(\log\log(n))$ [21]. And a following work [22] gave a $q$-ary $2$-deletion correcting code with redundancy $5\log(n)+O(\log\log(n))$ for all $q>2$. When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23]. In addition, they showed that there is a linear time encoder with near-optimal redundancy for their code.
If deletions occur at consecutive positions, we call them a burst of deletions. Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 24], wireless sensor networks and satellite communication devices [25]. A code is called a $b$-burst-deletion correcting code, if it can correct any single burst of exactly $b$ deletions. In 1970, Levenshtein gave a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [26]. For $b\geq 3$, Cheng et al in 2014 constructed a class of binary codes with redundancy $b\log(n/b+1)$ [27], which was later improved to $\log n+(b-1)\log\log n+O(1)$ by Schoeny et al in 2017 [28]. Schoeny’s result was generalized to non-binary alphabets in [29, 30]. The best known redundancy for all $q\geq 2$ is $\log n+O(1)$, which was contributed recently by Sun et al[31]. It was proved in [28, 30] that the redundancy of a $b$-burst-deletion correcting code is at least $\log n+\Omega(1)$. Therefore, the codes in [26] and [31, Theorem 9] have redundancy optimal up to a constant. There are also a lot of works on codes correcting single burst of at most $b$ deletions [26, 28, 32, 33, 21, 34, 35, 31].
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. To the best of our knowledge, there is no explicit results about such codes. A related result can be found in a work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively. Let $t=2$ and $t_{L}=b$. Then their result gives a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes. On the other hand, [17, 19] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique. Specifically, [17] gave a binary code with redundancy at most $7\log n+O(1)$ and [19] gave a binary code with redundancy at most $4\log n+O(\log\log n)$. Note that in a burst of deletions, all deletions occur consecutively. Therefore, intuitively, it is reasonable to deem that there is no big difference between two-deletion correcting codes and codes for correcting two $b$-burst-deletions (where $b>1$). This raises a natural question: for $b>1$, is there a construction which leads to codes which are as good as or even better than the one given by the syndrome compression technique. Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$. Our contributions include:
  • We establish lower and upper bounds on the size (or equivalently, the redundancy) of $2$-$b$-burst-deletion correcting codes;
  • A binary code of length $n$ with redundancy at most $6\log n+(15b-5)\log\log n+O(1)$;
  • A $q$-ary code of length $n$ with redundancy at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q\geq 2$;
  • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
Here, it is assumed that $q$ and $b$ are constants with respect to $n$.
The rest of this paper is organized as follows. In Section II, we introduce some necessary definitions and related results. In Section III, we bound above and below the size (or equivalently, the redundancy) of codes correcting two bursts of exactly $b$ deletions. Section IV and Section V deal with binary and non-binary codes, respectively. In Section VI, we show that our idea in Section V can result in a new construction of non-binary two-deletion correcting code. Finally Section VII concludes this paper.
II Preliminaries
In this section, we introduce some necessary terminologies and notations.
For an integer $q\geq 2$ and a positive integer $n$, denote $\Sigma_{q}=\left\{0,1,\ldots,q-1\right\}$ and $\Sigma_{q}^{n}$ the set of all $q$-ary sequences with $n$ symbols. Let $\mathbf{x}\in\Sigma_{q}^{n}$ be a sequence. Unless otherwise stated, the $i$th coordinate of $\mathbf{x}$ is denoted by $x_{i}$, i.e., $\mathbf{x}=x_{1}\cdots x_{n}$. We call $n$ the length of $\mathbf{x}$ and denote $\left|\mathbf{x}\right|=n$. For a finite set $A$, we denote by $\left|A\right|$ the cardinality of $A$.
For two integers $m$ and $n$ such that $m\leq n$, let $[m,n]$ denote the set $\left\{m,m+1,\ldots,n\right\}$. If $m=1$, denote $[n]=[1,n]$. For a sequence $\mathbf{x}\in\Sigma_{q}^{n}$ and a subset $I=\left\{i_{1},i_{2},\ldots,i_{t}\right\}\subseteq[n]$ where $i_{1}<i_{2}<\cdots<i_{t}$, we define $\mathbf{x}_{I}\triangleq x_{i_{1}}x_{i_{2}}\cdots x_{i_{t}}$. For each $I\subseteq[n]$, we say $\mathbf{x}_{I}$ is a subsequence of $\mathbf{x}$. In particular, if $I$ is an interval of $[n]$ (i.e., $I=[i,j]$ for some $1\leq i\leq j\leq n$), we say $\mathbf{x}_{I}$ is a substring of $\mathbf{x}$. A run in $\mathbf{x}$ is a maximal substring consisting of the same symbols. The number of runs of $\mathbf{x}$ is denoted by $r(\mathbf{x})$. For example, if $\mathbf{x}=100101$, then there are five runs in $\mathbf{x}$: $1$, $00$, $1$, 00 and $1$. So $r\left(\mathbf{x}\right)=5$.
The concatenation of two sequences $\mathbf{x}$ and $\mathbf{y}$ is denoted by $\mathbf{x}\mathbf{y}$. For example, let $\mathbf{x}=102$ and $\mathbf{y}=121$ be two sequences in $\Sigma_{3}^{3}$, then $\mathbf{x}\mathbf{y}=102121\in\Sigma_{3}^{6}$. Let $b$ and $n$ be two positive integers satisfying $b<n$. When a substring of length $b$ is deleted, we refer to it as a deletion-burst of size $b$ or a $b$-burst-deletion; that is to say, from $\mathbf{x}\in\Sigma_{q}^{n}$, we obtain a subsequence $\mathbf{x}_{[n]\setminus[i,i+b-1]}$ for some $11\leq i\leq n-b+$ 1.
In this paper, we focus on codes correcting two $b$-burst-deletions. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$, where $n>2b$. There are two ways to define two $b$-burst-deletions:
  1. (D1)
    the two bursts are caused by two channels: $\mathbf{x}$ passes the first channel, resulting in $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$ and then $\mathbf{z}$ passes the second channel, resulting in $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$;
  2. (D2)
    the two bursts are caused by a single channel: symbols in $\mathbf{x}$ pass a channel sequentially and we receive $\mathbf{y}=\mathbf{x}_{[n]\setminus I_{1}\cup I_{2}}$, where $I_{1}$ and $I_{2}$ are two disjoint intervals of length $b$ in [n].
In fact, (D1) and (D2) are equivalent. Firstly, it is clear that (D1) covers (D2). Next, we show that (D2) also covers (D1).
Remark II.1
There is another possibility: the two bursts might overlap and result in a single burst that is shorter than $2b$. We do not take this situation into account, since it is covered by a more comprehensive problem: correcting two bursts of deletions, where each burst has length at most $b$. Our idea in this paper fails in this situation. We left this problem for future research.
Observation II.1
Let $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}$ is obtained from $\mathbf{x}$ by process (D1). Then there exist two intervals $I_{1}=[j_{1},j_{1}+b-1]$, $I_{2}=[j_{2},j_{2}+b-1]\subseteq[n]$, where $j_{2}-j_{1}\geq b$, such that $\mathbf{y}=\mathbf{x}_{[n]\setminus(I_{1}\cup I_{2})}$. In particular, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by two $b$-burst-deletions, we can always assume that $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two non-overlap substrings of length $b$ from $\mathbf{x}$.
Proof:
By assumption, there is $11\leq i_{1}\leq n-b+$ 1 and $11\leq i_{2}\leq n-2b+$ 1 such that $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$, where $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$. If $i_{2}\geq i_{1}$, let $j_{1}=i_{1}$ and $j_{2}=i_{2}+b$. Then the conclusion follows.
Now suppose $11\leq i_{2}\leq i_{1}-$ 1. If $i_{2}\leq i_{1}-b$, let $j_{1}=i_{2}$ and $j_{2}=i_{1}$. Then the conclusion follows. If $i_{1}-b<i_{2}<i_{1}$, it is clear that $\mathbf{y}=\mathbf{x}_{[n]\setminus[i_{2},i_{2}+2b-1]}$. Let $j_{1}=i_{2}$ and $j_{2}=i_{2}+b$. Then the conclusion follows. ∎
For $t\in\{1,2\}$ and $n>tb$, define
[…]
When $b=1$, we use notation $\mathcal{B}_{t}(\mathbf{x})$ instead of $\mathcal{B}_{t}^{b}\left(\mathbf{x}\right)$.
Definition II.1
Let $\mathcal{C}$ be a subset of $\Sigma_{q}^{n}$ with $\left|\mathcal{C}\right|\geq 2$. Suppose $t\in\{1,2\}$. We call $\mathcal{C}$ a $t$-$b$-burst-deletion correcting code if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for any two distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}$. In particular, if $b=1$, we call $\mathcal{C}$ a $t$-deletion correcting code.
Clearly, if any $\mathbf{x}\in\mathcal{C}$ can be uniquely and efficiently recovered from any given $\mathbf{x}^{\prime}\in\mathcal{B}_{t}^{b}(\mathbf{x})$, then $\mathcal{C}$ is a $t$-$b$-burst-deletion correcting code. Here, “efficiently” means that the time complexity of decoding $\mathbf{x}$ from $\mathbf{y}$ is polynomial in $n$. In this paper, we construct $t$-$b$-burst-deletion correcting codes and show that any codeword can be uniquely and efficiently decoded.
The redundancy of a code $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is defined to be $\rho(\mathcal{C})=\log\left(q^{n}/\left|\mathcal{C}\right|\right)$. All logarithms in this paper are to the base $2$. In addition, we always assume that $q$ and $b$ are fixed with respect to the code-length $n$.
Let $n^{\prime}$ and $n$ be two positive integers satisfying $n^{\prime}<n$. For each sequence $\mathbf{x}\in\Sigma_{q}^{n}$, let $\widetilde{\mathbf{x}}$ be the zero padding of $\mathbf{x}$ to the shortest length that is greater than $n$ and is divisible by $n^{\prime}$, that is, $\widetilde{\mathbf{x}}=\mathbf{x}0^{\left\lceil n/n^{\prime}\right\rceil n^{\prime}-n}$, and then $n^{\prime}\mid\left|\widetilde{\mathbf{x}}\right|$. We can represent $\mathbf{x}$ as an $n^{\prime}\times\left\lceil n/n^{\prime}\right\rceil$ array $A\left(\mathbf{x},n,n^{\prime}\right)=\left[a_{i,j}\right]$, where $a_{i,j}=\tilde{x}_{i+n^{\prime}j}$ for all $1\leq i\leq n^{\prime}$ and $10\leq j\leq\left\lceil n/n^{\prime}\right\rceil-$ 1. In other words, the $i$-th row of $A\left(\mathbf{x},n,n^{\prime}\right)$ is
$A\left(\mathbf{x},n,n^{\prime}\right)_{i}\triangleq\left(\tilde{x}_{i},\tilde{x}_{i+n^{\prime}},\tilde{x}_{i+2n^{\prime}},\ldots,\tilde{x}_{i+n^{\prime}\left(\left\lceil\frac{n}{n^{\prime}}\right\rceil-1\right)}\right).$
We call $A\left(\mathbf{x},n,n^{\prime}\right)$ a matrix representation of $\mathbf{x}$. If $n$ and $n^{\prime}$ are clear from the context, we will denote $A(\mathbf{x},n,n^{\prime})$ by $A(\mathbf{x})$.
For example, let $n=7$, $\mathbf{x}=1011010\in\Sigma_{2}^{7}$ and $n^{\prime}=2$. Then $\widetilde{\mathbf{x}}=10110100\in\Sigma_{2}^{8}$ and
$A\left(\mathbf{x},7,2\right)=\begin{pmatrix}1&1&0&0\\ 0&1&1&0\end{pmatrix}.$
If $n^{\prime}=3$, then $\widetilde{\mathbf{x}}=\mathbf{x}00$ and so
$A\left(\mathbf{x},8,3\right)=\begin{pmatrix}1&1&0\\ 0&0&0\\ 1&1&0\end{pmatrix}.$
In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as an array $A(\mathbf{x},n,b)$. To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$. All results in this paper still hold even if $b\nmid n$, as long as we replace $n/b$ by $\left\lceil n/b\right\rceil$.
The next observation is a straightforward corollary of II.1.
Observation II.2
For each $b<n$ and $b\mid n$, we can represent a sequence $\mathbf{x}\in\Sigma_{2}^{n}$ as a $b\times n/b$ array $A(\mathbf{x})$. Any two $b$-burst-deletions in $\mathbf{x}$ will induce two deletions in each row of $A(\mathbf{x})$. Furthermore, if the positions of the two deletions in $A(\mathbf{x})_{1}$ are $j_{1}$ and $j_{2}$, then for each $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred at coordinate $j_{1}-1$ or $j_{1}$, and the other deletion occurred at coordinate $j_{2}-1$ or $j_{2}$.
II-A Related Results
Recently, there have been two notable developments in non-binary two-deletion correcting codes with low redundancy. In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for all even $q>2$. In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for all $q>2$. In Section VI, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for all $q>2$.
Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available. A related result can be found in the work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively. For $t=2$ and $t_{L}=b$, their result provides a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. Their result was derived using the syndrome compression technique, which was later extended to syndrome compression with pre-coding in [20]. We will apply this extended technique to provide a $q$-ary code with redundancy at most $7\log n+O\left(\log n/\log\log n\right)$, for all $q\geq 2$.
For a subset $\mathcal{E}\subseteq\Sigma_{q}^{n}$ and a sequence $\mathbf{x}\in\mathcal{E}$, define
$\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)=\left\{\mathbf{y}\in\mathcal{E}~{}:~{}\mathbf{y}\neq\mathbf{x}\text{ and }\mathcal{B}_{2}^{b}\left(\mathbf{y}\right)\cap\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)\neq\emptyset\right\}.$
Lemma II.1
[13, 20, 36] Let $\mathcal{E}\subseteq\Sigma_{q}^{n}$ be a code and $N>\max\left\{\left|\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)\right|~{}:~{}\mathbf{x}\in\mathcal{E}\right\}$. Suppose that the function $f:\Sigma_{q}^{n}\rightarrow\{0,1\}^{R(n)}$ (where $R(n)$ is a function of $n$ and $R(n)\geq 2$) satisfies the following property:
  1. $(\textup{{P}}1)$
    if $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}\in\mathcal{N}_{\Sigma_{q}^{n}}\left(\mathbf{x}\right)$, then $f\left(\mathbf{x}\right)\neq f\left(\mathbf{y}\right)$.
Then there exists a function $\bar{f}:\mathcal{E}\rightarrow\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, computable in polynomial time, such that $\bar{f}\left(\mathbf{x}\right)\neq\bar{f}\left(\mathbf{y}\right)$ for any $\mathbf{x}\in\mathcal{E}$ and $\mathbf{y}\in\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$.
Let $\mathcal{E}$ be a $1$-$b$-burst correcting code and $\bar{f}$ be given in Lemma II.1. Then Lemma II.1 asserts that if $\mathbf{x},\mathbf{y}\in\mathcal{E}$ are distinct codewords and $\bar{f}(\mathbf{x})=\bar{f}(\mathbf{y})$, we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$. Therefore, for any $\mathbf{a}\in\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, the code =$\mathcal{E}^{\prime}=\left\{\mathbf{x}\in\mathcal{E}:\bar{f}(\mathbf{x})=a\right\}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, by the pigeonhole principle, there exists an $\mathbf{a}$ such that the redundancy of $\mathcal{E}^{\prime}$ is at most $\rho\left(\mathcal{E}^{\prime}\right)=\rho(\mathcal{E})+2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$ [^2] We select $\mathcal{E}$ to be a $1$-$b$-burst correcting code to obtain a better redundancy. If we take $\mathcal{E}=\Sigma_{q}^{n}$ we get a redundancy of $8\log n+o(\log n)$..
For the choice of $\mathcal{E}$, we have the following result.
Lemma II.2
[31, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n>b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
This lemma gives a $1$-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$. Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|\leq q^{b}n^{3}$. So we let $N=q^{b}n^{3}$. Now the redundancy of $\mathcal{E}^{\prime}$ is at most $7\log n+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$. To conclude our discussion, it remains to find an $f$ satisfying (P1) in Lemma II.1 such that $R(n)=O(\log n)$, which is given in Lemma II.5. The proof of Lemma II.5 is based on Lemmas II.3 and II.4.
The following lemma is a corollary of [17, Theorem 2].
Lemma II.3
[17, Theorem 2] For any integer $n\geq 3$, there exists a function $\xi:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{7\log n+O\left(1\right)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\xi\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}^{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
Lemma II.4
Suppose $q\geq 2$. There is a function $\xi_{1}:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{7\left\lceil\log q\right\rceil\log n+O_{q}(1)}$, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ and $\xi_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q}(1)$ is a constant dependent only on $q$.
Proof:
Any $\mathbf{x}\in\Sigma_{q}^{n}$ can be uniquely represented as an array
$M(\mathbf{x})\triangleq\begin{pmatrix}x_{1,1}&\cdots&x_{1,n}\\ \vdots&\cdots&\vdots\\ x_{\left\lceil\log q\right\rceil,1}&\cdots&x_{\left\lceil\log q\right\rceil,n}\end{pmatrix},$
where $x_{k,i}\in\{0,1\}$ be such that $x_{i}=\sum_{k=1}^{\left\lceil\log q\right\rceil}x_{k,i}2^{k-1}$ for all $1\leq i\leq n$.
Denote the $k$-th row of $M(\mathbf{x})$ by $M(\mathbf{x})_{k}$. Suppose $\mathbf{y}\in\mathcal{B}_{2}(\mathbf{x})$. It is clear that $M(\mathbf{y})_{k}\in\mathcal{B}_{2}\left(M(\mathbf{x})_{k}\right)$ for all $1\leq k\leq\left\lceil\log q\right\rceil$. Let $\xi(\cdot)$ be the function defined in Lemma II.3. For $\mathbf{x}\in\Sigma_{q}^{n}$, define $\xi_{1}(\mathbf{x})\triangleq\left(\xi\left(M(\mathbf{x})_{1}\right),\ldots,\xi\left(M(\mathbf{x})_{\left\lceil\log q\right\rceil}\right)\right)$. Then by Lemma II.3, given $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ and $\xi_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi\left(M(\mathbf{x})_{k}\right)$ is a binary vector of length $7\log n+O(1)$, we can see that $\xi_{1}(\mathbf{x})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log n+O_{q}(1)$. ∎
Lemma II.5
Suppose $b>1$, $n>2b$ and $q\geq 2$. There is a function $\psi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)}$, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$.
Proof:
Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{y})=A(\mathbf{y},n-b,b)$. Since $\mathbf{y}\in\mathcal{B}_{2}^{b}(\mathbf{x})$, we have $A(\mathbf{y})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. Define $\psi(\mathbf{x})\triangleq\left(\xi_{1}\left(A(\mathbf{x})_{1}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b}\right)\right)$. Then by Lemma II.4, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi_{1}(A(\mathbf{x})_{i})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log(n/b)+O_{q}(1)$, we can see that $\psi(\mathbf{x})$ is a binary vector of length $7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)$. ∎
Taking $f=\psi$ gives a function satisfying (P1) as needed.
Before proceeding to subsequent sections, we introduce two useful lemmas. The first one is a well-known result from number theory, which will be frequently used. The second one will be used in Sections IV-A and V-A.
Lemma II.6 (Bertrand’s Postulate/Bertrand–Chebyshev theorem)
For every integer $n\geq 2$, there is always at least one prime $p$ such that $n\leq p<2n$.
Let $N>2$ be an integer. Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$. Let ???? denote an unknown symbol (not in $[0,N-1]$). If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ be such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for all $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two). For our purpose in Sections IV-A and V-A, assume $j\geq i+2$. For a sequence $\mathbf{x}$ over the alphabet $[0,N-1]$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,2)$ and $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
Lemma II.7
Let $Q\geq N$ be the smallest prime. For all $0\leq a_{1},a_{2},b_{1},b_{2}<Q$ and $0\leq c<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
  • (C1)
    $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\equiv a_{1}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\equiv a_{2}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\equiv b_{1}\pmod{Q}$,
    $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\equiv b_{2}\pmod{Q}$;
  • (C2)
    $W(\mathbf{x})\equiv c\pmod{nN^{2}}$, where $W(\mathbf{x})=\textup{Syn}\left(A(\mathbf{x})_{1}\right)+2N\cdot\textup{Syn}\left(A(\mathbf{x})_{2}\right)$.
Then the code $\mathcal{C}$ can correct two bursts of erasures. In particular, there is a function $\mathcal{E}_{1}:[0,N-1]^{n}\rightarrow\Sigma_{2}^{\log n+6\log N+4}$, computable in linear time, such that for any $\mathbf{x}\in[0,N-1]^{n}$, given $\mathcal{E}_{1}(\mathbf{x})$, we can efficiently and uniquely recover $\mathbf{x}$ from $\mathbf{y}$, where $\mathbf{y}$ is any given sequence obtained from $\mathbf{x}$ by two bursts of erasures.
Proof:
Firstly, if the correctness of the code $\mathcal{C}$ is proved, we can define
$\mathcal{E}_{1}(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\pmod{Q},\right.$
$\left.\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\pmod{Q},W(\mathbf{x})\pmod{nN^{2}}\right).$
Here, we view $\mathcal{E}(\mathbf{x})$ as a binary vector. By Lemma II.6, we have $N\leq Q<2N$. Now it is easy to see that the length of $\mathcal{E}(\mathbf{x})$ is at most $\log n+6\log N+4$. It remains to prove the correctness of $\mathcal{C}$.
Suppose $\mathbf{y}$ is obtained from a codeword $\mathbf{x}\in\mathcal{C}$ by two bursts of erasures. Then $A(\mathbf{y})_{i}$ is obtained from $A(\mathbf{x})_{i}$ by two erasures for each $i=1,2$. Suppose the error positions in $A(\mathbf{y})_{1}$ are $i_{1},i_{2}$, and the error positions in $A(\mathbf{y})_{2}$ are $i_{3},i_{4}$, where $i_{1}<i_{2}$ and $i_{3}\leq i_{4}$. Clearly, we have $i_{3}\in\{i_{1},i_{1}-1\}$ and $i_{4}\in\{i_{2},i_{2}-1\}$. This implies that $i_{4}-i_{3}\in\{i_{2}-i_{1},i_{2}-i_{1}-1,i_{2}-i_{1}+1\}$. By definition, the two bursts in $\mathbf{x}$ do not overlap. It follows that $i_{2}\geq i_{1}+2$ if $i_{4}-i_{3}=i_{2}-i_{1}-1$.
To decode $\mathbf{x}$ from $\mathbf{y}$. We first calculate the two erased symbols in each row. Notice that $N>2$ and hence $Q$ is an odd prime. Suppose $\sigma_{1},\sigma_{2}\in[0,N-1]$ satisfy
$\sigma_{1}+\sigma_{2}\equiv t_{1}\pmod{Q},$ (1)
$\sigma_{1}^{2}+\sigma_{2}^{2}\equiv t_{2}\pmod{Q}$
where $t_{1}$ and $t_{2}$ are given. Then we can uniquely recover the (multi)set $\{\sigma_{1},\sigma_{2}\}$ by solving system (1). Now we can determine the erased symbols in $A(\mathbf{y})_{1}$ and $A(\mathbf{y})_{2}$ by (C1). Suppose that the two erased symbols in $A(\mathbf{y})_{1}$ are $\alpha_{1},\alpha_{2}$, and the two erased symbols in $A(\mathbf{y})_{2}$ are $\alpha_{3},\alpha_{3}$. It remains to determine positions of $\alpha_{1}$ and $\alpha_{2}$ in $A(\mathbf{x})_{1}$, and positions of $\alpha_{3}$ and $\alpha_{4}$ in $A(\mathbf{x})_{2}$. We can assume $\alpha_{1}\neq\alpha_{2}$ or $\alpha_{3}\neq\alpha_{4}$, since otherwise, we can uniquely recover $\mathbf{x}$.
There are at most four choices for the positions of $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ and for each choice, we get a candidate for $\mathbf{x}$. We need to show that among these candidates, there is exactly one equals $\mathbf{x}$. Firstly, by the relationship of $\mathbf{x}$ and $\mathbf{y}$, it is clear that there is at least candidate equals $\mathbf{x}$. Next, we will show there is at most one such candidate. To that end, let $(\beta_{1},\beta_{2})$ be a permutation of $(\alpha_{1},\alpha_{2})$, and $(\beta_{3},\beta_{4})$ be a permutation of $(\alpha_{3},\alpha_{4})$, where $(\beta_{1},\beta_{2})\neq(\alpha_{1},\alpha_{2})$ or $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$. Let $\mathbf{u}$ be obtained from $\mathbf{y}$ by assigning $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively. Let $\mathbf{v}$ be obtained from $\mathbf{y}$ by assigning $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ at coordinates $i_{1},i_{2},i_{3},i_{4}$, respectively. By assumption, we have $\mathbf{u}\neq\mathbf{v}$. We need to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ is not a codeword in $\mathcal{C}$. Let $\Delta=W(\mathbf{u})-W(\mathbf{v})$. Then we have
$\Delta=i_{1}(\alpha_{1}-\beta_{1})+i_{2}(\alpha_{2}-\beta_{2})+2N\left(i_{3}(\alpha_{3}-\beta_{3})+i_{4}(\alpha_{4}-\beta_{4})\right).$
To complete the proof, it is sufficient to show that at least one of $\mathbf{u}$ and $\mathbf{v}$ does not satisfies (C2). Suppose on the contrary that $W(\mathbf{u})\equiv c\equiv W(\mathbf{v})\pmod{nN^{2}}$. Then we have $\Delta\equiv 0\pmod{nN^{2}}$.
If $(\beta_{1},\beta_{2})=(\alpha_{1},\alpha_{2})$, it holds that $(\beta_{3},\beta_{4})\neq(\alpha_{3},\alpha_{4})$. Therefore, we conclude that $\alpha_{3}\neq\alpha_{4}$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$. Then it follow that $\left|\Delta\right|=2N(i_{4}-i_{3})\left|\alpha_{3}-\alpha_{4}\right|$ and hence $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)<nN^{2}$, which is a contradiction. Similarly, if $(\beta_{3},\beta_{4})=(\alpha_{3},\alpha_{4})$, we have $0<\left|\Delta\right|\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)<nN^{2}$, which is a contradiction.
Now suppose $\alpha_{1}\neq\alpha_{2}$, $\alpha_{3}\neq\alpha_{4}$, $(\beta_{1},\beta_{2})=(\alpha_{2},\alpha_{1})$ and $(\beta_{3},\beta_{4})=(\alpha_{4},\alpha_{3})$. Then we have $\Delta=(i_{2}-i_{1})(\alpha_{2}-\alpha_{1})+2N(i_{4}-i_{3})(\alpha_{4}-\alpha_{3})$. Without loss of generality, suppose $\alpha_{1}<\alpha_{2}$. If $\alpha_{3}<\alpha_{4}$, then $0<\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(N-1)+\left\lfloor\frac{n}{2}\right\rfloor 2N(N-1)\leq n(N^{2}-1)$, which is a contradiction. Now suppose $\alpha_{3}>\alpha_{4}$. Recall that $i_{4}-i_{3}\in\{i_{2}-i_{1}-1,i_{2}-i_{1},i_{2}-i_{1}+1\}$. When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it is easy to verify that $0<-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor(2N(N-1)-1)+2N(N-1)<nN^{2}$, which is a contradiction. When $i_{4}-i_{3}=i_{2}-i_{1}-1$, we have $i_{2}\geq i_{1}+2$. Then it follows that
$0<2N(\alpha_{3}-\alpha_{4})-2(\alpha_{2}-\alpha_{1})\leq-\Delta\leq\left\lfloor\frac{n}{2}\right\rfloor\left(2N(N-1)-1\right)-2N<nN^{2},$
which is a contradiction. ∎
III Bounds
We could not find any existing upper or lower bounds on the maximum size of a $2$-$b$-burst-deletion code. In this section, we will derive these bounds.
Let $M_{q,n,b}$ be the maximum size of a $2$-$b$-burst-deletion correcting code in $\Sigma_{q}^{n}$, where $n>2b$.
Theorem III.1
$M_{n,q,b}\geq\frac{q^{n-2b}}{\binom{n}{2}^{2}}$.
Proof:
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. It is clear that a subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [37, page 100, Theorem 1] that
$M_{q,n,b}\geq\sum_{\mathbf{x}\in\Sigma_{q}^{n}}\frac{1}{d(\mathbf{x})+1}.$ (2)
For $\mathbf{x}\in\Sigma_{q}^{n}$, by II.1, we conclude that each $\mathbf{y}$ (including $\mathbf{x}$ itself) with $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$ can be obtained as follows: 1) deleting two non-overlap substrings of length $b$ from $\mathbf{x}$, resulting in a sequence $\mathbf{z}\in\Sigma_{q}^{n-2b}$; 2) inserting two sequences of length $b$ into $\mathbf{z}$. Therefore, we have $d(\mathbf{x})+1\leq\binom{n}{2}^{2}q^{2b}$. ∎
In [38, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
Theorem III.2
For $q\geq 2$, let $f(q)=\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. When $n\geq 30$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
$M_{q,n,b}\leq\left(\frac{3b^{2}}{q^{2b-2}(q-1)^{2}}+\frac{(1.121)^{3b}}{n}\right)\frac{q^{n}}{n^{2}}.$
Proof:
Define $m=n/b-1$. Let $\mathcal{C}\subseteq\Sigma_{q}^{n}$ be a $2$-$b$-busrt-deletion correcting code. Set $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and $t=\left(1-\frac{1}{q}-\epsilon\right)m$. We partition $\mathcal{C}$ into two disjoint subsets: $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$, where $\mathcal{C}_{1}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\geq t+2\text{ for some }1\leq i\leq b\right\}$ and $\mathcal{C}_{2}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\leq t+1\text{ for all }1\leq i\leq b\right\}$ (recall that $r(\cdot)$ denotes the number of runs). To derive an upper bound of $\left|\mathcal{C}\right|$, it is sufficient to upper bound $\left|\mathcal{C}_{1}\right|$ and $\left|\mathcal{C}_{2}\right|$.
For any $\mathbf{x}\in\mathcal{C}$, define $\mathcal{A}_{2}^{b}(\mathbf{x})=\left\{A(\mathbf{x}^{\prime}):\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})\right\}$. Let $A(\ell,k)$ be the array obtained by deleting the $\ell$-th and the $k$-th columns of $A(\mathbf{x})$ for $1\leq\ell\neq k\leq n/b$. It is clear that $A(\ell,k)\in\mathcal{A}_{2}^{b}(\mathbf{x})$ and
$\cup_{1\leq\ell\neq k\leq\frac{n}{b}}\left\{A(\ell,k)_{i}\right\}=\mathcal{B}_{2}(A(\mathbf{x})_{i}),$
for all $1\leq i\leq b$. By [39, eq. (11)], we know that
$\binom{r-1}{2}\leq\left|\mathcal{B}_{2}(\mathbf{v})\right|\leq\binom{r+1}{2}$
for any sequence $\mathbf{v}\in\Sigma_{q}^{n}$ with exactly $r$ runs. Since $\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|$ and $\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|=\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|$ for all $\mathbf{x}$, it follows that
$\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|\geq\binom{\max_{1\leq i\leq b}\left\{r\left(A(\mathbf{x})_{i}\right)\right\}-1}{2}\geq\binom{t+1}{2},$ (3)
for all $\mathbf{x}\in\mathcal{C}_{1}$. Since $\mathcal{C}$ is a $2$-$b$-burst-deletion correcting code, we know that $\mathcal{C}_{1}$ is also a $2$-$b$-burst-deletion correcting code. So we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for all distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}_{1}$. Then it follows from Equation 3 that $\left|\mathcal{C}_{1}\right|\binom{t+1}{2}\leq\sum_{\mathbf{x}\in\mathcal{C}_{1}}\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\leq q^{n-2b}$ and hence
$\left|\mathcal{C}_{1}\right|\leq\frac{q^{n-2b}}{\binom{t+1}{2}}\leq\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{t^{2}}$ $=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{\left(1-1/q-\epsilon\right)^{2}\left(n/b-1\right)^{2}}$ (4)
$=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2b^{2}}{(1-1/q-\epsilon)^{2}(1-b/n)^{2}}$
$=\frac{2b^{2}q^{n+2}}{n^{2}q^{2b}(q-1)^{2}}\cdot\frac{1}{\left(1-\frac{q}{q-1}\epsilon\right)^{2}\left(1-b/n\right)^{2}}$
$\leq\frac{3b^{2}q^{n+2}}{n^{2}q^{2b}(q-1)^{2}}$
as long as $\left(1-\frac{q}{q-1}\epsilon\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, which is possible when $n$ is sufficiently large.
Next, we proceed to upper bound $\left|\mathcal{C}_{2}\right|$. To that end, define a set
$\mathcal{C}^{\prime}\triangleq\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(b)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in\Sigma_{q}^{n/b},\;\forall 1\leq i\leq b\\ r\left(\mathbf{x}^{(i)}\right)\leq t+1,\;\forall 1\leq i\leq b\end{array}\right\}.$
Since $r\left(A(\mathbf{x})_{i}\right)\leq t+1$ for any $\mathbf{x}\in\mathcal{C}_{2}$ and $1\leq i\leq b$, we conclude that $\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|$. So it suffices to estimate an upper bound of $\left|\mathcal{C}^{\prime}\right|$.
Since $m=n/b-1$, by definition of $\mathcal{C}^{\prime}$, we have
$\left|\mathcal{C}^{\prime}\right|$ =|$=\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})\leq t+1\right\}\right|^{b}$ (5)
$=\left|\cup_{j=1}^{t+1}\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|^{b}$
$=\left(\sum_{j=1}^{m+1}\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|\right)^{b}$
=$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
where (a) follows from the well-known result (see the proof of [40, Theorem 3.1]):
|$\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|=\binom{m}{j-1}q(q-1)^{j-1}.$
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [41, Proposition 3.3.3] that
$\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\leq q^{mH_{q}(t/m)}=q^{mH_{q}(1-1/q-\epsilon)},$ (6)
where $H_{q}(x)\triangleq x\log_{q}(q-1)-x\log_{q}x-(1-x)\log_{q}(1-x)$ is the $q$-ary entropy function. Here, $\log_{q}x=\frac{\log x}{\log q}$ for any positive real number $x$.
If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. So by Lemma .1, we have.
$H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
Now it follows from Equations 5 and 6 that
$\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}\overset{(b)}{<}(1.121)^{3b}\frac{q^{n}}{n^{3}},$ (7)
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 4 and 7. ∎
Then next corollary is a direct consequence of Theorems III.1 and III.2.
Corollary III.1
Suppose that $q$ and $b$ are fixed. Let $f(q)$ be defined in Theorem III.2. When $n\geq\max\left\{30,\frac{(1.21)^{3b}q^{2b-2}(q-1)^{2}}{b^{2}}\right\}$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
$2\log n+(2b-2)\log q+2\log(q-1)-2\log b-2\leq\rho(\mathcal{C})\leq 4\log n+2b\log q,$
where $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code with maximum size.
IV Binary $2$-$b$-Burst-Deletion Correcting Codes
In this section, we will construct a class of binary $2$-$b$-burst-deletion correcting codes with redundancy at most $7\log n+O(\log\log n)$. First, we introduce some $1$-deletion correcting codes and $2$-deletion correcting codes, which will be useful later.
Lemma IV.1
[21, Lemma 1] For any integer $n\geq 3$, there exists a function $\mathsf{VT}:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{\log(n+1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\mathsf{VT}\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
Definition IV.1 (regularity)
A sequence $\mathbf{x}\in\Sigma_{2}^{n}$ is said to be $d$-regular if each substring of $\mathbf{x}$ of length at least $d\log n$ contains both $00$ and $11$.
For the number of regular sequences, we have the following lemma.
Lemma IV.2
[19, Lemma 11] The number of $d$-regular sequences of length $n$ is at least $2^{n-1}$, as long as $d\geq 7$ and $n$ be such that $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{0.15d-1}\geq 12$. In particular, when $d=7$, it suffices to require $n\geq 9$.
Lemma IV.3
[19, Theorem 7] Suppose $d\geq 7$. There exists a function $\eta:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{4\log n+10\log\log n+O_{d}(1)}$, computable in linear time, such that for any $d$-regular sequence $\mathbf{x}\in\Sigma_{2}^{n}$, given $\eta\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{d}(1)$ denotes a constant depending only on $d$.
If the first row $A(\mathbf{x})_{1}$ is regular, then by correcting the two deletions in $A(\mathbf{x})_{1}$, we can approximate the location of the two deletions, as ensured by the following lemma.
Lemma IV.4
[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$. Then we can
  1. $(1)$
    either find distinct runs $\mathbf{x}_{J_{1}}$ and $\mathbf{x}_{J_{2}}$ of $\mathbf{x}$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, such that $i_{1}\in J_{1}$ and $i_{2}\in J_{2}$.
  2. $(2)$
    or find an interval $J\subseteq[n]$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, of length at most $3d\log n$ such that $i_{1},i_{2}\in J$.
From now on, we assume that $A(\mathbf{x})_{1}$ is $7$-regular. Furthermore, suppose $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$. Then by II.2, we have $A\left(\mathbf{x}^{\prime}\right)_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Lemma IV.3 ensures that when given $\eta\left(A(\mathbf{x})_{1}\right)$, one can uniquely recover $A(\mathbf{x})_{1}$ from $A\left(\mathbf{x}^{\prime}\right)_{1}$. Then by Lemma IV.4, we can know the approximate positions of deletions in $A(\mathbf{x})_{1}$. This further enables us to approximately determine the positions of deletions in $A(\mathbf{x})_{i}$ for all $2\leq i\leq b$. Based on the two cases in Lemma IV.4, we split our discussion into Section IV-A and Section IV-B. Our codes for correcting two $b$-burst-deletions is presented in Theorem IV.1, whose proof is divided in into Lemmas IV.5, IV.6 and IV.7.
IV-A When Case (1) in Lemma IV.4 happens
Let $A(\mathbf{x})_{1,I_{j}}$ be the $j$th run of $A(\mathbf{x})_{1}$, where $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$ (recall that $r\left(A(\mathbf{x})_{1}\right)$ is the number of runs in $A(\mathbf{x})_{1}$) and $I_{j}\subseteq[n/b]$ is the index set of $A(\mathbf{x})_{1,I_{j}}$. For simplicity, denote $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$ and $p_{r\left(A(\mathbf{x})_{1}\right)}=n/b$. Since $A(\mathbf{x})_{1}$ is $7$-regular, we have $p_{j}-p_{j-1}=\left|I_{j}\right|\leq 7\log(n/b)$. Note that $I_{j}$ depends on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $j_{1}<j_{2}$. Then for all $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $\left[p_{j_{1}-1},p_{j_{1}}\right]$ (if $j_{1}=1$, then this deletion occurred in the interval $[1,p_{1}]$) and the other occurred in the interval $\left[p_{j_{2}-1},p_{j_{2}}\right]$. So $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin\left[p_{j_{1}-1},p_{j_{1}}\right]\cup\left[p_{j_{2}-1},p_{j_{2}}\right]$. Specifically, we have
$A(\mathbf{x})_{i,k}=\begin{cases}A(\mathbf{x}^{\prime})_{i,k},\text{ if }k<p_{j_{1}-1},\\ A(\mathbf{x}^{\prime})_{i,k-1},\text{ if }p_{j_{1}}<k<p_{j_{2}-1},\\ A(\mathbf{x}^{\prime})_{i,k-2},\text{ if }k>p_{j_{2}}.\end{cases}$ (8)
for all $2\leq i\leq b$.
Therefore, we need to decode $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ from $A(\mathbf{x}^{\prime})_{i}$, for all $2\leq i\leq b$. Notice that when $j\geq 2$, the interval $[p_{j-1},p_{j}]$ intersects with both $I_{j-1}$ and $I_{j}$. Our process of decoding $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ from $A(\mathbf{x}^{\prime})_{i}$ is divided in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$.
For the first step, for all $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$, define
$f_{j}(\mathbf{x})=\left(A(\mathbf{x})_{2,p_{j}},\ldots,A(\mathbf{x})_{b,p_{j}}\right),$
which can be viewed as an integer in $\left[0,N-1\right]$, where $N=2^{b-1}$. For uniform discussion, we also let $f_{0}(\mathbf{x})=0$. Now define $L(\mathbf{x})=f_{0}(\mathbf{x})f_{1}(\mathbf{x})\cdots f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})\in[0,N-1]^{r\left(A(\mathbf{x})_{1}\right)+1}$. In other word, sequence $L(\mathbf{x})$ is the concatenation of $f_{0}(\mathbf{x}),f_{1}(\mathbf{x}),\ldots,f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})$. To recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$ for all $2\leq i\leq b$, it is sufficient to obtain the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$. By Equation 8, the value of $f_{j}(\mathbf{x})$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
In the following, we always assume $j_{2}-j_{1}\geq 2$, since otherwise, the two deletions in $A(\mathbf{x})_{i}$ occurred in an interval of length at most $\left|I_{j_{1}}\cup I_{j_{2}}\right|+1\leq 14\log(n/b)+1$ and this situation is covered by case (2) in Lemma IV.4. Now we have $j_{2}-1\geq j_{1}-1+2$. Then it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures. To determine the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$, we apply Lemma II.7. Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7. Define
$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
Since $r(A(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+1)+6\log N+4}=2^{\log n+O_{b}(1)}$. Now the following lemma follows from above analysis and Lemma II.7.
Lemma IV.5
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$. Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known. Then given $f(\mathbf{x})$, we can recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Upon knowing the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$, it remains to recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$. Recall that $j_{2}-j_{1}\geq 2$. This implies $p_{j_{2}-1}>p_{j_{1}}$ and thus for all $2\leq i\leq b$, the deletion in $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ does not affect $A(\mathbf{x})_{i,I_{j_{1}}}$. Therefore, we can conclude that $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{1}}}\right)$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{2}}}\right)$, for all $2\leq i\leq b$. So we can apply Lemma IV.1 to decode $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}$, respectively.
For $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$, define
$g_{j}(\mathbf{x})=\left(\mathsf{VT}\left(A(\mathbf{x})_{2,I_{j}}\right),\cdots,\mathsf{VT}\left(A(\mathbf{x})_{b,I_{j}}\right)\right).$
Since $A(\mathbf{x})_{1}$ is $7$-regular, we have $\left|I_{j}\right|\leq 7\log(n/b)$. Then it follows from Lemma IV.1 that
$\mathsf{VT}\left(A(\mathbf{x})_{i,I_{j}}\right)\in\{0,1\}^{\log\left(7\log(n/b)+1\right)}.$
So $g_{j}(\mathbf{x})$ can be viewed as an integer in $[0,N_{2}-1]$, where $N_{2}=2^{(b-1)\log\log(n)+O_{b}(1)}$. Here, $O_{b}(1)$ is a constant dependent only on $b$.
Let $Q_{1}\geq\max\{N_{2},n/b\}$ be the smallest prime. Define
$g^{(0)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}g_{j}(\mathbf{x})\pmod{2N_{2}},$ (9)
$g^{(1)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}jg_{j}(\mathbf{x})\pmod{Q_{1}}$
Lemma IV.6
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$. Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known. If the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are known, then given $g^{(0)}(\mathbf{x})$ and $g^{(1)}(\mathbf{x})$, we can recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Proof:
As discussed above, we have $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{1}}}\right)$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{2}}}\right)$, for all $2\leq i\leq b$. So by Lemma IV.1, to recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $\mathbf{x}^{\prime}$, it is sufficient to obtain the values of $g_{j_{1}}(\mathbf{x})$ and $g_{j_{2}}(\mathbf{x})$.
Since the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are given, the value of $g_{j}(\mathbf{x})$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}g_{j}(\mathbf{x})\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}jg_{j}(\mathbf{x})\right)\pmod{Q_{1}}$. Then it follows from Equation 9 that
$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{2}},$ (10)
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$ (11)
Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 10 that $g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})=\delta^{(0)}$. Combining this with Equation 11, we get
$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{1}},$ (12)
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
Since $1\leq j_{1}<j_{2}\leq n/b\leq Q_{1}$, system (12) can be uniquely solved and the unique solution is
$\left(g_{j_{1}}(\mathbf{x})\pmod{Q_{1}},g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}\right).$
Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}\leq Q_{1}$, we have $g_{j_{1}}(\mathbf{x})=g_{j_{1}}(\mathbf{x})\pmod{Q_{1}}$ and $g_{j_{2}}(\mathbf{x})=g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}$. ∎
IV-B When Case (2) in Lemma IV.4 happens
In this case, we can find an interval $J\subseteq[n/b]$ of length at most $21\log(n/b)$ such that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting two symbols in $A(\mathbf{x})_{1,J}$. Suppose $J=[c,d]$. Then for all $2\leq i\leq b$, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $[c-1,d]$, which is of length at most $121\log n+$ 1.
Let $P=21\log\left(n/b\right)+1$ and
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/bP\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/bP\right\rceil-1.\end{cases}$
Let $\xi(\cdot)$ be the function defined in Lemma II.3. For each $11\leq j\leq\left\lceil n/bP\right\rceil-$ 1, define
$h_{j}\left(\mathbf{x}\right)=\left(\xi\left(A(\mathbf{x})_{2,J_{j}}\right),\xi\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi\left(A(\mathbf{x})_{b,J_{j}}\right)\right).$
Since $\left|J_{j}\right|\leq 2P=42\log\left(n/b\right)+2$ for all $j$, each $h_{j}(\mathbf{x})$ is a binary vector of length at most $7(b-1)\log(42\log\left(n/b\right)+2)+O_{b}(1)$ and so we can view $h_{j}(\mathbf{x})$ as an integer in $\left[0,N_{3}-1\right]$, where $N_{3}=2^{7(b-1)\log(42\log\left(n/b\right)+2)+O_{b}(1)}$. Let
$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\text{ even}}h_{j}(\mathbf{x})\pmod{N_{3}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\text{ odd}}h_{j}(\mathbf{x})\pmod{N_{3}}.\end{array}$ (13)
Lemma IV.7
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting two symbols in $A(\mathbf{x})_{1,J}$, where $J\subseteq[n/b]$ is an interval of length $21\log(n/b)$ and is known to us. Given $h_{e}\left(\mathbf{x}\right)$ and $h_{o}\left(\mathbf{x}\right)$, one can uniquely recover $A(\mathbf{x})_{i}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Proof:
We can find a $j_{0}$ such that $J\subseteq J_{j_{0}}$. Without loss of generality, we assume that $j_{0}$ is even. Then $A(\mathbf{x})_{i,J_{j}}$ and thus $h_{j}(\mathbf{x})$ is known to us for each even $j\neq j_{0}$. So it remains to recover $A(\mathbf{x})_{J_{j_{0}}}$. If $J_{j_{0}}=[c,d]$, it is clear that $A(\mathbf{x}^{\prime})_{i,[c,d-2]}=\mathcal{B}_{2}\left(A(\mathbf{x})_{i,[c,d]}\right)$, for all $2\leq i\leq b$. By Lemma II.3, it is suffices to obtain $h_{j_{0}}(\mathbf{x})$. By Equation 13, we have
$h_{j_{0}}(\mathbf{x})\equiv\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\text{ even},j\neq j_{0}}h_{j}(\mathbf{x})\right)\pmod{N_{3}}.$
Since $0\leq h_{j_{0}}(\mathbf{x})<N_{2}$, it follows that
$h_{j_{0}}(\mathbf{x})=\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\text{ even},j\neq j_{0}}h_{j}(\mathbf{x})\right)\pmod{N_{3}}.$
Now the proof is completed. ∎
Now we are ready to present our main result in this section.
Theorem IV.1
Suppose $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{1}$ as
$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{2}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }7\text{-regular},\\ \eta(A(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $n\geq 9$, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $6\log n+(15b-5)\log\log n+O_{b}(1)$.
Proof:
Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$. Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},n,b)$. Then $A(\mathbf{x}^{\prime})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Since $A(\mathbf{x})_{1}$ is $7$-regular, by Lemma IV.3, we can recover $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ based on $\eta\left(A(\mathbf{x})_{1}\right)$. Now according to Lemmas IV.4, IV.5, IV.6 and IV.7, it is clear that $\mathcal{C}$ is a $2$-$b$-burst-deletion correcting code.
Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$. By Lemma IV.2, there are at least $2^{n-1}$ $7$-regular sequences when $n\geq 9$. It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that
$\left|\mathcal{C}_{1}\right|\geq\frac{2^{n-1}}{N_{0}\cdot N_{1}\cdot(2N_{2})\cdot Q_{1}\cdot(N_{3})^{2}},$
which implies by definition that
$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+\log N_{1}+\log(2N_{2})+\log Q_{1}+2\log N_{3}+1$
$=5\log n+\log Q_{1}+(15b-5)\log\log n+O_{b}(1)$
$\leq 6\log n+(15b-5)\log\log n+O_{b}(1).$
The last inequality follows from the fact $n/b\leq Q_{1}<2n/b$, which is guaranteed by Lemma II.6. ∎
IV-C Another construction when $b=2$
Recall that in Section IV-A, the decoding of $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ is conducted in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$. This is because the interval $\left[p_{j-1},p_{j}\right]$ intersects both $I_{j-1}$ and $I_{j}$. A natural question is whether we can handle deletions in the interval $\left[p_{j-1},p_{j}\right]$ in one step, instead of two steps. Next, we show that this is possible when $b=2$.
Recall that in Section IV-A, once the first row $A(\mathbf{x})_{1}$ is recovered, the two deletions in the second row $A(\mathbf{x})_{2}$ are determined in two non-overlapping intervals $\left[p_{j_{1}-1},p_{j_{1}}\right]$ and $\left[p_{j_{2}-1},p_{j_{2}}\right]$ in $[n/2]$ if we assume $p_{j_{2}-1}>p_{j_{1}}$ (otherwise, the two deletions can be corrected using the method in Section IV-B). In Lemmas 21, 22 and 23 in [42], Lu et al proved the follows result, which is a straightforward corollary of Lemma 5 and Observation 2 in [19].
Lemma IV.8
[42, Lemma 23] Let $m$, $P_{1}$ and $P_{2}$ be positive integers, where $P_{1}+P_{2}<m$. Let $P=\min\{P_{1},P_{2}\}$. There is a function $\mathcal{E}_{2}:\Sigma_{2}^{m}\rightarrow\Sigma_{2}^{2\log m+\log P+O(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{m}$, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two bits at two known non-overlapping intervals of length at most $P_{1}$ and $P_{2}$ respectively, then given $\mathcal{E}_{2}(\mathbf{x})$ and $\mathbf{y}$, one can uniqely and efficently recover $\mathbf{x}$.
Note that $p_{j_{1}}-p_{j_{1}-1},p_{j_{2}}-p_{j_{2}-1}\leq 7\log(n/2)+1$. So, let $P_{1}=P_{2}=7\log(n/2)+1$ and $m=n/2$. Then we can apply the function $\mathcal{E}_{2}$ in Lemma IV.8 to handle interval $\left[p_{j_{1}-1},p_{j_{1}}\right]$ (and $\left[p_{j_{2}-1},p_{j_{2}}\right]$) in one step, introducing only $2\log n+\log\log n+O(1)$ additional redundant bits. Now replacing the six functions $f^{(s,\ell)}$, $g^{(}0)$ and $g^{(}1)$ in the definition of $\mathcal{C}_{1}$ by only one function $\mathcal{E}_{2}$, we have the following corollary.
Corollary IV.1
Let $b=2$ and $n\geq 9$. There is a binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $6\log n+25\log\log n+O(1)$.
V Non-binary $2$-$b$-Burst-Deletion Correcting Codes
Let $q\geq 2$. In this section, we will extend our idea in Section IV to construct $q$-ary codes capable of correcting two $b$-burst-deletions. Surprisingly, our construction in this section gives rise to a binary code which is better than that in Theorem IV.1 when $b\geq 17$.
The first step is to establish a result similar to Lemma IV.4 to approximately determine the error positions. To that end, we associate with any $q$-ary sequence a binary sequence.
For any $x\in\Sigma_{q}$, we can uniquely write it as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<\left\lceil q/2\right\rceil$. Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$. Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$. It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
Remark V.1
Although any $x\in\Sigma_{q}$ can be decomposed as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, it might exist some $u\in\{0,1\}$ and $0\leq v<\left\lceil q/2\right\rceil$ such that $\left\lceil q/2\right\rceil u+v\notin\Sigma_{q}$. In fact, if $q$ is even, then for any $u\in\{0,1\}$ and any $0\leq v<\left\lceil q/2\right\rceil$, we have $\left\lceil q/2\right\rceil u+v\in\Sigma_{q}$. But when $q$ is odd, we have $\left\lceil q/2\right\rceil+\left\lceil q/2\right\rceil-1=q\notin\Sigma_{q}$.
We first generalize the concept of $d$-regular sequence to sequences over general alphabets.
Definition V.1
A sequence $\mathbf{x}\in\Sigma_{q}^{n}$ is said to be $d$-regular if $u(\mathbf{x})$ is $d$-regular. In other words, $\mathbf{x}$ is $d$-regular if and only if any substring of $\mathbf{x}$ of length at least $d\log n$ contains two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$.
[…]
Lemma V.1
Let $q\geq 2$.
  • If $q$ is even, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n}/2\geq q^{n-1}$ for all $d\geq 7$ and $n$ such that $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{0.15d-1}\geq 12$. In particular, when $d=7$, it suffices to require $n\geq 9$.
  • If $q$ is odd, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n-1}$ for all $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. Note that when $d\geq 10$, we have $-1-\frac{d}{2}\log(0.87)>0$ and hence this condition could be satisfied when $n$ is sufficiently large. In particular, when $d=10$, it suffices to require $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
Proof:
Suppose first that $q$ is even. Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. Then it follows from Lemma IV.2 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
Now suppose that $q\geq 3$ is odd. In this case, previous argument does not hold. This is because when $q$ is odd, $u_{x}=1$ and $v_{x}=(q-1)/2$, we have $\left\lceil q/2\right\rceil u_{x}+(q-1)/2=q\notin\Sigma_{q}$. Fortunately, we can apply similar ideas in proofs of [19, Lemma 11] and [22, Lemma 5] to derive our lower bound on $\left|\mathcal{R}_{q,n,d}\right|$. For $m\geq 2$, define
$S_{m}^{L}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }0\leq v_{i},v_{i+1}\leq(q-1)/2\right\},$
$S_{m}^{U}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }(q+1)/2\leq v_{i},v_{i+1}<q\right\}.$
We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$. By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}^{3}\right)=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$. Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$. For $\mathbf{v}\in S_{m}^{L}$, it must holds that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$. So follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
For $\mathbf{v}\in\Sigma_{q}^{m}$, let $\overline{\mathbf{v}}\triangleq(q-v_{1})\cdots(q-v_{m})$. It is easy to verify that if $\mathbf{v}\in S_{m}^{U}$, then $\overline{\mathbf{v}}\in S_{m}^{L}$. Then it follows that $\left|S_{m}^{U}\right|\leq\left|S_{m}^{L}\right|<(0.87q)^{m}$.
Let $m=\left\lfloor\frac{d}{2}\log n\right\rfloor$ and
$Q=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\exists i,j\text{ such that }v_{i},v_{i+1}<(q+1)/2\text{ and }v_{j}.v_{j+1}\geq(q+1)/2\right\}.$
Then we have $\left|Q\right|\geq q^{m}-\left|S_{m}^{L}\right|-\left|S_{m}^{U}\right|>q^{m}-2(0.87q)^{m}$. Now let $k=\left\lfloor n/m\right\rfloor$ and define
=$\mathcal{R}^{\prime}=\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(k)}\mathbf{x}^{(k+1)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 1\leq i\leq k\\ \mathbf{x}^{(k+1)}\in\Sigma_{q}^{n-k\left\lfloor n/m\right\rfloor}.\end{array}\right\}$
It is easy to see that $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$ and thus
$\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}^{\prime}\right|$ $=\left|Q\right|^{k}q^{n-km}$
$>\left(q^{m}-2(0.87q)^{m}\right)^{k}q^{n-km}$
$=q^{n}\left(1-2(0.87)^{m}\right)^{k}$
$\overset{(a)}{\geq}q^{n}\left(1-2k(0.87)^{m}\right)$
$=q^{n}\left(1-2\left\lfloor\frac{n}{m}\right\rfloor 2^{\left\lfloor\frac{d}{2}\log n\right\rfloor\log(0.87)}\right)$
$\geq q^{n}\left(1-2\frac{n}{m}2^{\left\lfloor\frac{d}{2}\log n\right\rfloor\log(0.87)}\right)$
$\overset{(b)}{\geq}q^{n}\left(1-2\frac{n}{m}2^{\frac{d}{2}\log n\cdot\log(0.87)-\log(0.87)}\right)$
$=q^{n}\left(1-\frac{2}{0.87\left\lfloor\frac{d}{2}\log n\right\rfloor}n^{1+\frac{d}{2}\log(0.87)}\right)$
$\overset{(c)}{\geq}q^{n-1}.$
Here, (a) follows from the fact that $(1+x)^{r}\geq 1+rx$ for any integer $r\geq 1$ and any real number $x\geq-1$; (b) follows from the fact that $\log(0.87)<0$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor\geq\frac{d}{2}\log n-1$; (c) follow from the fact that $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. ∎
In the rest of this section, for any $q$-ary sequence $\mathbf{x}$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,b)$ and $U(\mathbf{x})=A(u(\mathbf{x}),\left|\mathbf{x}\right|,b)$. It is clear that $U(\mathbf{x})_{i}=u\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let
$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\}.$
Then $A(\mathbf{x})_{1}=\mathbf{x}_{I_{n,b}}$ and $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. Like in Section IV-A, let $U(\mathbf{x})_{1,I_{j}}$ ($1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$) be all runs in $U(\mathbf{x})_{1}$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$.
Recall that $I_{j}$ is an interval in $[n/b]$. For all $j\geq 1$ We associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$. In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$. In other words, $I_{j}^{L}$ is the interval of length at most $b-1$ in $[n]$ to the left of $I_{j}^{\prime}$. See Figure 1 for an illustration. Note that $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$ are dependent on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
[…]
(a) $U(\mathbf{x})$: $I_{j}$ in $\left[n/b\right]$
[…]
(b) $\mathbf{x}$: $I_{j}^{\prime}$ and $I_{j}^{L}$ in $\left[n\right]$
Figure 1: Illustration of relationship between $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$.
Lemma V.2
Let $q\geq 2$, $b>1$ and $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular. Let $\eta(\cdot)$ be the function defined in Lemma IV.3. Then given $\eta\left(U(\mathbf{x})_{1}\right)$, we can
  1. $(1)$
    either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$;
  2. $(2)$
    or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$.
Proof:
Since $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, we have $u(\mathbf{x}^{\prime})=u(\mathbf{x})_{[n]\setminus(D_{1}\cup D_{2})}$. It follows that $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$. Suppose that $U(\mathbf{x}^{\prime})_{1}$ is obtained from $U(\mathbf{x})_{1}$ by deleting $U(\mathbf{x})_{1,k_{1}}$ and $U(\mathbf{x})_{1,k_{2}}$. Recall that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. So $U(\mathbf{x}^{\prime})_{1}$ is obtained from $u(\mathbf{x})_{I_{n,b}}$ by deleting $u(\mathbf{x})_{1+(k_{1}-1)b}$ and $u(\mathbf{x})_{1+(k_{2}-1)b}$.
Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.3, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$. In addition, by Lemma IV.4, we can
  • either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $k_{1}\in I_{j_{1}}$ and $k_{2}\in I_{j_{2}}$;
  • or find an interval $[c,d]\subseteq[n/b]$ of length at most $3d\log(n/b)$ such that $k_{1},k_{2}\in[c,d]$.
By II.2, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Note that for the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$. Now the proof is completed. ∎
In the following, we always assume $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Furthermore, suppose that $A(\mathbf{x})_{1}$ is $d$-regular. We need to find conditions on $\mathbf{x}$ which enable us to recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$. As in Section IV, we split our discussion into Sections V-A and V-B. Our main result in this section is presented in Theorem V.1.
V-A When Case (1) in Lemma V.2 happens
In this case, we already found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Note that $I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}=\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]$ and $I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}=\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]$. Then it follows that
$x_{k}=\begin{cases}x^{\prime}_{k},\text{ if }k<p_{j_{1}-1}b-b+2,\\ x^{\prime}_{k-b},\text{ if }p_{j_{1}}b<k<p_{j_{2}-1}b-b+2,\\ x^{\prime}_{k-2b},\text{ if }p_{j_{2}}b<k\leq n.\end{cases}$ (14)
We can assume $j_{2}-j_{1}\geq 2$, since otherwise, we have $\left|I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\cup I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\right|\leq 2bd\log(n/b)+b-1$, and therefore, this situation is covered by case (2) in Lemma V.2.
Our first goal is to recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$. Recall that $\left|I_{j}^{L}\right|\leq b-1$. So we can view $\mathbf{x}_{I_{j}^{L}}$ as an integer in $[0,N-1]$, where $N=q^{b-1}$. Note that $I_{1}^{L}=\emptyset$. For uniform discussion, we also let $\mathbf{x}_{I_{1}^{L}}=0$ and $\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}=\mathbf{x}_{\left[n-b+2,n\right]}$. Now define $L(\mathbf{x})=\mathbf{x}_{I_{1}^{L}}\cdots\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}}\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}\in[0,N-1]^{r\left(U(\mathbf{x})_{1}\right)+2}$. In other word, sequence $L(\mathbf{x})$ is the concatenation of $\mathbf{x}_{I_{1}^{L}},\ldots,\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}},\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}$. By Equation 14, the value of $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\notin\{j_{1},j_{1}+1,j_{2},j_{2}+1\}$.
Since $j_{2}\geq j_{1}+2$, it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures. To determine the values of $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, we apply Lemma II.7. Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7. Define
$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
Since $r(U(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+2)+10\log N+4}=2^{\log n+O_{q,b}(1)}$. Now the following lemma follows from above analysis and Lemma II.7.
Lemma V.3
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ from $\mathbf{x}^{\prime}$.
Once we know $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, it remains to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, where $I_{j_{1}}^{\prime}=\left[p_{j_{1}-1}b+1,j_{1}b\right]$ and $I_{j_{1}}^{\prime}=\left[p_{j_{2}-1}b+1,j_{2}b\right]$. Since $j_{2}-j_{1}\geq 2$, we have $(p_{j_{2}-1}-1)b+2>p_{j_{1}}b$. This implies that the $b$-burst-deletion in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ does not affect $\mathbf{x}_{I_{j_{1}}^{\prime}}$. Therefore, we conclude that $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. So to decode $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$ from $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}$, we need the function $\phi(\cdot)$ defined in Lemma II.2.
Next, we will apply $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ to decode $\mathbf{x}_{I_{j}^{\prime}}$. Recall that $U(\mathbf{x})_{1}$ is $d$-regular. So by definition, we have $\left|I_{j}^{\prime}\right|=(p_{j}-p_{j-1})b\leq db\log n$. Then by Lemma II.2, $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{\log(db\log n)+O_{q,b}(1)}$. Let $Q_{1}\geq\max\{n/b,N_{2}\}$ be the smallest prime. Define
$g^{(0)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{2N_{2}},$ (15)
$g^{(1)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{Q_{1}}.$
Lemma V.4
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Furthermore, suppose that $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known. Then given $g^{(0)}(\mathbf{x})$ and $g^{(1)}(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$ from $\mathbf{x}^{\prime}$.
Proof:
Since $j_{2}-j_{1}\geq 2$, we have $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. So by Lemma II.2, to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, it it sufficient to obtain $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known, it is easy to see that $\mathbf{x}_{I_{j}^{\prime}}$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{Q_{1}}$. Then it follows from Equation 15 that
$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{2N_{2}},$ (16)
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$ (17)
Since $0\leq\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right),\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 16 that $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)=\delta^{(0)}$. Combining this with Equation 17, we get
$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{Q_{1}},$ (18)
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$
Now similar to Lemma IV.6, we can uniquely solve $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$ from (18). ∎
V-B When Case (2) in Lemma V.2 happens
In this case, we can find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1 such that such that $D_{1},D_{2}\subseteq J$. This implies that $\mathbf{x}_{k}$ is known to us for all $k\notin J$. Suppose $J=[c_{1},c_{2}]$. Then $\mathbf{x}^{\prime}_{[c_{1},c_{2}-2b]}\in\mathcal{B}_{b}^{2}\left(\mathbf{x}_{J}\right)$. To recover $\mathbf{x}_{J}$, we apply a trivial $2$-$b$-burst-deletion correcting code of length $2\left|J\right|$, which is given in Lemma II.5.
Let $P=3bd\log(n/b)+b-1$ and
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/P\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/P\right\rceil-1.\end{cases}$
For $s\in\{0,1\}$, let $L_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$. Let the function $\psi(\cdot)$ be defined in Lemma II.5. Since $\left|J_{j}\right|\leq 2P$, we can view each $\psi\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{3}-1\right]$, where $N_{3}=2^{7b\left\lceil\log q\right\rceil\log(2P/b)+O_{q,b}(1)}$ and define
$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{0}}\psi\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{3}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{1}}\psi\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{3}}.\end{array}$ (19)
Lemma V.5
Suppose that we found an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1 such that such that $D_{1},D_{2}\subseteq J$. Then given $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$, we can uniquely recover $\mathbf{x}_{J}$ (and thus $\mathbf{x}$) from $\mathbf{x}^{\prime}$.
Proof:
We can find a $j_{0}$ such that $J\subseteq J_{j_{0}}$. Without loss of generality, we assume $j_{0}\in L_{0}$. Then $\psi\left(\mathbf{x}_{J_{j}}\right)$ is known to us for each $j\in L_{0}\setminus\{j_{0}\}$. In addition, if $J_{j_{0}}=[c_{1},c_{2}]$, it is clear that $\mathbf{x}^{\prime}_{[c_{1},c_{2}-2b]}=\mathcal{B}_{2}^{b}\left(\mathbf{x}_{J_{0}}\right)$. It follows from Equation 19, that
$h_{j_{0}}(\mathbf{x})\equiv\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}\psi\left(\mathbf{x}_{J_{j}}\right)\right)\pmod{N_{3}}.$
Since $0\leq h_{j_{0}}(\mathbf{x})<N_{2}$, we have
$h_{j_{0}}(\mathbf{x})=\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}\psi\left(\mathbf{x}_{J_{j}}\right)\right)\pmod{N_{3}}.$
Now by Lemma II.5, one can uniquely recover $\mathbf{x}_{J_{0}}$ and thus $\mathbf{x}_{J}$. ∎
Theorem V.1
Suppose $q\geq 2$, $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{2}$ as
$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{2}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O_{q,b}(1)$:
  • $q$ is even, $d=7$ and $n\geq 9$;
  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
Proof:
Suppose $\mathbf{x}\in\mathcal{C}_{2}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma V.2, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$. For the former case, Lemmas V.3 and V.4 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. For the latter case, Lemma V.5 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. Therefore, $\mathcal{C}_{2}$ can correct two $b$-burst-deletions.
By Lemma V.1, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $U(\mathbf{x})_{1}$ is $d$-regular. Now the claim about redundancy follows from the pigeonhole principle. ∎
Remark V.2
Let $q=2$. Then Theorem V.1 gives a binary code correcting two $b$-burst-deletions with redundancy at most $6\log n+(14b+11)\log\log n+O_{b}(1)$. Comparing this with Theorem IV.1, we can see that when $2\leq b\leq 16$, the code in Theorem IV.1 has lower redundancy, while when $b\geq 17$, the code in Theorem V.1 has lower redundancy.
VI Non-binary Two-deletion Correcting Codes
In this section, we always assume $q>2$. We aim to give a new construction of $q$-ary two-deletion correcting codes. Recall the definition of $u(\mathbf{x})$ for any $\mathbf{x}\in\Sigma_{q}^{n}$. If $\mathbf{x}^{\prime}\in\mathcal{B}_{2}(\mathbf{x})$, then $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}(u(\mathbf{x}))$. So if $u(\mathbf{x})$ belongs to a binary two-deletion correcting code, we can recover $u(\mathbf{x})$ from $u(\mathbf{x}^{\prime})$. We further assume that $\mathbf{x}$ (and thus $u(\mathbf{x})$) is $d$-regular. Then Lemma IV.4 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions. This enables us to approximately determine error positions in $\mathbf{x}$. For the latter case, we need a $q$-ary two-deletion correcting code, which is given in Lemma II.4. For the former case, we need a $q$-ary single-deletion correcting code. We use the code given in [23]
Lemma VI.1
[23] Suppose $q>2$ and $n>2$. There is a function $\mathsf{DVT}:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+\log q}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathsf{DVT}(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$.
Let $u(\mathbf{x})_{I_{j}}$ ($1\leq j\leq r\left(u(\mathbf{x})\right)$) be all the runs in $u(\mathbf{x})$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(u(\mathbf{x})\right)}=n$ and $p_{j-1}<p_{j}$ for all $j$. Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$. Therefore, by Lemma VI.1, $\mathsf{DVT}\left(I_{j}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$. Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime. Define
$f_{0}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{2N_{1}},$ (20)
$f_{1}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{Q}.$
The two functions $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$ will help to deal with case (1) in Lemma IV.4. For case (2) in Lemma IV.4, let $P=3d\log n$ and
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/P\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/P\right\rceil-1.\end{cases}$
Since $\left|J_{1}\right|\leq 2P=6d\log n$, we can view each $\xi_{1}\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{7\left\lceil\log q\right\rceil\log(6d\log n)+O_{q}(1)}$ (see Lemma II.4). For $s\in\{0,1\}$, let $K_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$ and define
$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{0}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{1}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}}.\end{array}$ (21)
Now we are ready to give our code.
Theorem VI.1
Suppose $q>2$. Let $N_{1},N_{2},Q$, $f_{0}(\mathbf{x})$, $f_{1}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{0}<2N_{1}$, $0\leq b_{1}<Q$, and $0\leq c_{1},c_{2}<N_{2}$, define the code $\mathcal{C}_{3}$ as
$\mathcal{C}_{3}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}\text{ is }d\text{-regular},\\ \eta(u(\mathbf{x}))=a,\\ f_{s}(\mathbf{x})=b_{s}\text{ for }s\in\{0,1\},\\ h_{e}(\mathbf{x})=c_{1},h_{o}(\mathbf{x})=c_{2}\end{array}\right\}$
Then $\mathcal{C}_{3}$ is a two-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{0}$, $b_{1}$, $c_{1}$ and $c_{2}$, such that the redundancy of $\mathcal{C}_{3}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
  • $q$ is even, $d=7$ and $n\geq 9$;
  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
The proof of Theorem VI.1 is similar to those of Theorems IV.1 and V.1, and thus omitted.
VII Conclusion
In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $7\log n+O(\log\log n)$ for all $b\geq 2$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique. Furthermore, when $b=2$, we show that there is a binary code with redundancy at most $6\log n+O\left(\log\log n\right)$. Inspired by these ideas, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
A more complex problem is to construct codes capable of correcting at most two bursts of deletions, where each burst has a length at most $b$. Our technique fails in this setup since we can not even know the number of bursts or the length of each burst occurring in a received sequence. We leave this problem for future research.
The following Lemma .1 is derived by slightly modifying the proof of [41, Proposition 3.3.7]. Before showing its proof, we list some facts.
Let $\ln(\cdot)$ be the natural logarithm. Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}-\cdots$.
Fact .1
For all $10\leq x<$ 1, we have $\ln(1-x)\leq-x-\frac{x^{2}}{2}$ and $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$. When $0\leq x\leq\frac{1}{2}$, we have $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
Proof:
When $10\leq x<$ 1, we have $\ln(1-x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{(-x)^{i}}{i}=-\sum_{i=1}^{\infty}\frac{x^{i}}{i}\leq-x-\frac{x^{2}}{2}$. Since $-\ln(1+x)=-x+\frac{x^{2}}{2}-\left(\ln(1+x)-x+\frac{x^{2}}{2}\right)$ and $\ln(1+x)-x+\frac{x^{2}}{2}\geq 0$ for all $x\geq 0$, we have $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$. At last, it is easy to see that $-\ln(1-x)=x+\frac{x^{2}}{2}+x^{3}\left(\frac{1}{3}+\frac{x}{4}+\frac{x^{2}}{5}+\cdots\right)\leq x+\frac{x^{2}}{2}+x^{3}(1+x+x^{2}+\cdots)=x+\frac{x^{2}}{2}+\frac{x^{3}}{1-x}$. When $x\leq\frac{1}{2}$, we have $1-x\geq\frac{1}{2}$ and thus $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$. ∎
Lemma .1
Suppose that $q\geq 2$ is fixed and $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. Then it holds that
$H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
Proof:
It is easy to verify that
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $=-\left(1-\frac{1}{q}-\epsilon\right)\log_{q}\left(\frac{1-1/q-\epsilon}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1}{q}+\epsilon\right)$ (22)
$=-\log_{q}\left(\frac{1}{q}\left(1-\frac{\epsilon q}{q-1}\right)\right)+\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)$
$=1-\frac{1}{\ln q}\left[\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)\right]$
$=1+\frac{1}{\ln q}\left[-\ln\left(1-\frac{\epsilon q}{q-1}\right)+\left(\frac{1}{q}+\epsilon\right)\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(1+\epsilon q\right)\right].$
Since $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q}\right\}$, we have $0<\frac{\epsilon q}{q-1}<\frac{1}{2}$ and $10<\epsilon q<$ 1. Then it follows from .1 that $-\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}}$, $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}$ and $-\ln(1+\epsilon q)\leq-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}$. Now by Equation 22, we have
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $\leq 1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q}{q-1}\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\;\;\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}\right)\right]$
$=1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q^{2}}{q-1}+\frac{\epsilon^{2}q^{3}(q-2)}{2(q-1)^{2}}\right)\right]$
$=1+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\frac{\epsilon^{2}q^{2}}{q-1}+\frac{\epsilon^{2}q^{2}(q-2)}{2(q-1)^{2}}\right]$
$=1-\frac{\epsilon^{2}q^{2}}{2(q-1)\ln q}+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}$
$\overset{(a)}{\leq}1-\frac{\epsilon^{2}}{4},$
where (a) follows from the fact that $\epsilon<\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)$. ∎