### v4

# Codes Correcting Two Bursts of Exactly $b$ Deletions

Abstract
In this paper, we investigate codes designed to correct two bursts of deletions, where each burst has a length of exactly $b$, where $b>1$. The previous best construction, achieved through the syndrome compression technique, had a redundancy of at most $7\log n+O\left(\log n/\log\log n\right)$ bits. In contrast, our work introduces a novel approach for constructing $q$-ary codes that attain a redundancy of at most $5\log n+O(\log\log n)$ bits for all $b>1$ and $q\geq 2$. Additionally, for the case where $b=1$, we present a new construction of $q$-ary two-deletion correcting codes with a redundancy of $5\log n+O(\log\log n)$ bits, for all $q>2$.
Index Terms:
deletion, burst-deletion, error-correcting codes, DNA-based storage
I Introduction
A subset $\mathcal{C}\subseteq\{0,1,\ldots,q-1\}^{n}$ (where $q\geq 2$) is called a $t$-deletion correcting code, if it has the property that if a codeword $\mathbf{x}\in\mathcal{C}$ is corrupted by deleting $t$ symbols to obtain a subsequence $\mathbf{y}\in\{0,1,\ldots,q-1\}^{n-t}$, then one can recover $\mathbf{x}$ from $\mathbf{y}$. The study of deletion correcting codes has a long history, dating back to at least the 1960s [1]. The seminal work in this field is [2], whereby proposing a linear-time decoding algorithm, Levenshtein proved that the binary code (Varshamov-Tenengolts code, or VT code for short) constructed in [3] can combat a single deletion error. In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4]. For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|C\right|\right)$ [^1] All logarithms in this paper are to the base $2$, is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lower bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6]. This implies that the VT code in [3] and the Tenengolts code [4] have redundancy optimal up to a constant.
Due to applications in DNA-based data storage [7, 8, 9], document exchange [10, 11], multiple-deletion correcting codes with low redundancy have attracted a lot of interest in recent years [10, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19, 24]. For general $t\geq 3$, the smallest redundancy, which is $(4t-1)\log n+o(\log n)$, was achieved by a construction given in [20]. For non-binary alphabets, Sima et al [15] presented a family of $q$-ary $t$-deletion correcting codes with redundancy $4t\log n+o(\log n)$ by using the syndrome compression technique. The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding. A straightforward application of this method will give a $q$-ary $t$-deletion correcting codes with redundancy $(4t-1)\log n+o(\log n)$, which is the best-known result in redundancy. When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary $2$-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ [21]. And in a following work [22] the authors presented a $q$-ary $2$-deletion correcting code with redundancy $5\log n+O(\log\log n)$ for all $q>2$. When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23]. In addition, they showed that there is a linear time encoder with near-optimal redundancy for their code.
If deletions occur at consecutive positions, we call them a burst of deletions. Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 25], wireless sensor networks, and satellite communication devices [26]. A code is called a $b$-burst-deletion correcting code, if it can correct any single burst of exactly $b$ deletions. In 1970, Levenshtein presented a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [27]. For $b\geq 3$, Cheng et al in 2014 constructed a class of binary codes with redundancy $b\log(n/b+1)$ [28], which was later improved to $\log n+(b-1)\log\log n+O(1)$ by Schoeny et al in 2017 [29]. Schoeny’s result was generalized to non-binary alphabets in [30, 31]. The best known redundancy for all $q\geq 2$ is $\log n+O(1)$, which was contributed recently by Sun et al[32]. It was proved in [29, 31] that the redundancy of a $b$-burst-deletion correcting code is at least $\log n+\Omega(1)$. Therefore, the codes in [27] and [32, Theorem 9] have redundancy optimal up to a constant. There are also a lot of works on codes correcting single burst of at most $b$ deletions [27, 29, 33, 34, 21, 35, 36, 32].
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. We call such codes $2$-$b$-burst correcting codes. To the best of our knowledge, there are no explicit results about such codes. A related result can be found in a work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively. Let $t=2$ and $t_{L}=b$. Then their result gives a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes. On the other hand, [17, 19, 21, 22, 24] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique. Specifically, [17] and [19, 24] presented binary codes with redundancy at most $7\log n+O(1)$ and $4\log n+O(\log\log n)$, respectively. For non-binary codes, the best-known redundancy is $5\log n+O(\log\log n)$ [21, 22]. Note that in a burst of deletions, all deletions occur consecutively. Therefore, it is reasonable to deem by intuition that there is no big difference between two-deletion correcting codes and $2$-$b$-burst-deletion correcting codes (where $b>1$). This raises a natural question: for $b>1$, is there a construction that leads to codes that are as good as, or even better than, the one given by the syndrome compression technique? Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$. Our contributions include:
  • We establish lower and upper bounds on the size (or equivalently, the redundancy) of $2$-$b$-burst-deletion correcting codes;
  • A binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+14b\log\log n+O(1)$, for all $b>1$;
  • A $q$-ary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O(1)$, for all $q>2$ and $b>1$;
  • A new construction of $q$-ary two-deletion correcting code of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q>2$.
Here, it is assumed that $q$ and $b$ are constants with respect to $n$. Therefore, our results show that for $2$-$b$-burst-deletion correcting codes, we can do almost as well as two-deletion correcting codes.
The rest of this paper is organized as follows. In Section II, we introduce some necessary definitions and related results. In Section III, we bound above and below the size (or equivalently, the redundancy) of codes correcting two bursts of exactly $b$ deletions. Section IV deals with codes for correcting two $b$-burst-deletions. In Section V, we give a new construction of non-binary two-deletion correcting code. Finally Section VI concludes this paper.
II Preliminaries
In this section, we first introduce some necessary definitions, auxiliary conclusions and related results.
For an integer $q\geq 2$ and a positive integer $n$, denote $\Sigma_{q}=\left\{0,1,\ldots,q-1\right\}$ and $\Sigma_{q}^{n}$ the set of all $q$-ary sequences with $n$ symbols. Let $\mathbf{x}\in\Sigma_{q}^{n}$ be a sequence. Unless otherwise stated, the $i$th coordinate of $\mathbf{x}$ is denoted by $x_{i}$, i.e., $\mathbf{x}=x_{1}\cdots x_{n}$. We call $n$ the length of $\mathbf{x}$ and denote $\left|\mathbf{x}\right|=n$. For a finite set $A$, we denote by $\left|A\right|$ the cardinality of $A$.
For two integers $m$ and $n$ such that $m\leq n$, let $[m,n]$ denote the set $\left\{m,m+1,\ldots,n\right\}$. If $m=1$, denote $[n]=[1,n]$. For a sequence $\mathbf{x}\in\Sigma_{q}^{n}$ and a subset $I=\left\{i_{1},i_{2},\ldots,i_{t}\right\}\subseteq[n]$ where $i_{1}<i_{2}<\cdots<i_{t}$, we define $\mathbf{x}_{I}\triangleq x_{i_{1}}x_{i_{2}}\cdots x_{i_{t}}$. For each $I\subseteq[n]$, we say $\mathbf{x}_{I}$ is a subsequence of $\mathbf{x}$. In particular, if $I$ is an interval of $[n]$ (i.e., $I=[i,j]$ for some $1\leq i\leq j\leq n$), we say $\mathbf{x}_{I}$ is a substring of $\mathbf{x}$. A run in $\mathbf{x}$ is a maximal substring consisting of the same symbols. The number of runs of $\mathbf{x}$ is denoted by $r(\mathbf{x})$. For example, if $\mathbf{x}=100101$, then there are five runs in $\mathbf{x}$: $1$, $00$, $1$, 00 and $1$. So $r\left(\mathbf{x}\right)=5$.
The concatenation of two sequences $\mathbf{x}$ and $\mathbf{y}$ is denoted by $\mathbf{x}\mathbf{y}$. For example, let $\mathbf{x}=102$ and $\mathbf{y}=121$ be two sequences in $\Sigma_{3}^{3}$, then $\mathbf{x}\mathbf{y}=102121\in\Sigma_{3}^{6}$. Let $b$ and $n$ be two positive integers satisfying $b<n$. When a substring of length $b$ is deleted, we refer to it as a deletion-burst of size $b$ or a $b$-burst-deletion; that is to say, from $\mathbf{x}\in\Sigma_{q}^{n}$, we obtain a subsequence $\mathbf{x}_{[n]\setminus[i,i+b-1]}$ for some $11\leq i\leq n-b+$ 1.
In this paper, we focus on codes correcting two $b$-burst-deletions. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$, where $n>2b$. There are two ways to define two $b$-burst-deletions:
  1. (D1)
    the two bursts are caused by two channels: $\mathbf{x}$ passes the first channel, resulting in $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$ and then $\mathbf{z}$ passes the second channel, resulting in $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$;
  2. (D2)
    the two bursts are caused by a single channel: symbols in $\mathbf{x}$ pass a channel sequentially and we receive $\mathbf{y}=\mathbf{x}_{[n]\setminus I_{1}\cup I_{2}}$, where $I_{1}$ and $I_{2}$ are two disjoint intervals of length $b$ in [n].
Remark II.1
There is another possibility: the two bursts might overlap and result in a single burst that is shorter than $2b$. We do not take this situation into account, since it is covered by a more comprehensive problem: correcting two bursts of deletions, where each burst has length at most $b$. Our idea in this paper fails in this situation. We left this problem for future research.
In fact, (D1) and (D2) are equivalent. Firstly, it is clear that (D1) covers (D2). Next, we show that (D2) also covers (D1).
Observation II.1
Let $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}$ is obtained from $\mathbf{x}$ by process (D1). Then there exist two intervals $I_{1}=[j_{1},j_{1}+b-1]$, $I_{2}=[j_{2},j_{2}+b-1]\subseteq[n]$, where $j_{2}-j_{1}\geq b$, such that $\mathbf{y}=\mathbf{x}_{[n]\setminus(I_{1}\cup I_{2})}$. In particular, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by two $b$-burst-deletions, we can always assume that $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two non-overlap substrings of length $b$ from $\mathbf{x}$.
Proof:
By assumption, there is $11\leq i_{1}\leq n-b+$ 1 and $11\leq i_{2}\leq n-2b+$ 1 such that $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$, where $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$. If $i_{2}\geq i_{1}$, let $j_{1}=i_{1}$ and $j_{2}=i_{2}+b$. Then the conclusion follows.
Now suppose $11\leq i_{2}\leq i_{1}-$ 1. If $i_{2}\leq i_{1}-b$, let $j_{1}=i_{2}$ and $j_{2}=i_{1}$. Then the conclusion follows. If $i_{1}-b<i_{2}<i_{1}$, it is clear that $\mathbf{y}=\mathbf{x}_{[n]\setminus[i_{2},i_{2}+2b-1]}$. Let $j_{1}=i_{2}$ and $j_{2}=i_{2}+b$. Then the conclusion follows. ∎
For $t\in\{1,2\}$ and $n>tb$, define
[…]
When $b=1$, we use notation $\mathcal{B}_{t}(\mathbf{x})$ instead of $\mathcal{B}_{t}^{b}\left(\mathbf{x}\right)$.
Definition II.1
Let $\mathcal{C}$ be a subset of $\Sigma_{q}^{n}$ with $\left|\mathcal{C}\right|\geq 2$. Suppose $t\in\{1,2\}$. We call $\mathcal{C}$ a $t$-$b$-burst-deletion correcting code if $\mathcal{B}_{t}^{b}(\mathbf{x})\cap\mathcal{B}_{t}^{b}(\mathbf{y})=\emptyset$ for any two distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}$. In particular, if $b=1$, we call $\mathcal{C}$ a $t$-deletion correcting code.
Clearly, if any $\mathbf{x}\in\mathcal{C}$ can be uniquely and efficiently recovered from any given $\mathbf{x}^{\prime}\in\mathcal{B}_{t}^{b}(\mathbf{x})$, then $\mathcal{C}$ is a $t$-$b$-burst-deletion correcting code. Here, “efficiently” means that the time complexity of decoding $\mathbf{x}$ from $\mathbf{y}$ is polynomial in $n$. In this paper, we construct $t$-$b$-burst-deletion correcting codes and show that any codeword can be uniquely and efficiently decoded.
The redundancy of a code $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is defined to be $\rho(\mathcal{C})=\log\left(q^{n}/\left|\mathcal{C}\right|\right)$. All logarithms in this paper are to the base $2$. In addition, we always assume that $q$ and $b$ are fixed with respect to the code-length $n$.
Let $n^{\prime}$ and $n$ be two positive integers satisfying $n^{\prime}<n$. For each sequence $\mathbf{x}\in\Sigma_{q}^{n}$, let $\widetilde{\mathbf{x}}$ be the zero padding of $\mathbf{x}$ to the shortest length that is greater than $n$ and is divisible by $n^{\prime}$, that is, $\widetilde{\mathbf{x}}=\mathbf{x}0^{\left\lceil n/n^{\prime}\right\rceil n^{\prime}-n}$, and then $n^{\prime}\mid\left|\widetilde{\mathbf{x}}\right|$. We can represent $\mathbf{x}$ as an $n^{\prime}\times\left\lceil n/n^{\prime}\right\rceil$ array $A\left(\mathbf{x},n,n^{\prime}\right)=\left[a_{i,j}\right]$, where $a_{i,j}=\tilde{x}_{i+n^{\prime}j}$ for all $1\leq i\leq n^{\prime}$ and $10\leq j\leq\left\lceil n/n^{\prime}\right\rceil-$ 1. In other words, the $i$-th row of $A\left(\mathbf{x},n,n^{\prime}\right)$ is
$A\left(\mathbf{x},n,n^{\prime}\right)_{i}\triangleq\left(\tilde{x}_{i},\tilde{x}_{i+n^{\prime}},\tilde{x}_{i+2n^{\prime}},\ldots,\tilde{x}_{i+n^{\prime}\left(\left\lceil\frac{n}{n^{\prime}}\right\rceil-1\right)}\right).$
We call $A\left(\mathbf{x},n,n^{\prime}\right)$ a matrix (or array) representation of $\mathbf{x}$. If $n$ and $n^{\prime}$ are clear from the context, we will denote $A(\mathbf{x},n,n^{\prime})$ by $A(\mathbf{x})$.
For example, let $n=7$, $\mathbf{x}=1011010\in\Sigma_{2}^{7}$ and $n^{\prime}=2$. Then $\widetilde{\mathbf{x}}=10110100\in\Sigma_{2}^{8}$ and
$A\left(\mathbf{x},7,2\right)=\begin{pmatrix}1&1&0&0\\ 0&1&1&0\end{pmatrix}.$
If $n^{\prime}=3$, then $\widetilde{\mathbf{x}}=\mathbf{x}00$ and so
$A\left(\mathbf{x},8,3\right)=\begin{pmatrix}1&1&0\\ 0&0&0\\ 1&1&0\end{pmatrix}.$
In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as matrix $A(\mathbf{x},n,b)$. To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$. All results in this paper still hold even if $b\nmid n$, as long as we replace $n/b$ by $\left\lceil n/b\right\rceil$. Throughout this paper, for a matrix $A$, denote by $A_{i}$ and $A_{i,j}$ the $i$-th row of $A$ and the entry in the $i$-th row and $j$-th column, respectively.
The next observation is a straightforward result of II.1.
Observation II.2
For each $b<n$ and $b\mid n$, we can represent a sequence $\mathbf{x}\in\Sigma_{2}^{n}$ as a $b\times n/b$ array $A(\mathbf{x})$. Any two $b$-burst-deletions in $\mathbf{x}$ will induce two deletions in each row of $A(\mathbf{x})$. Furthermore, if the positions of the two deletions in $A(\mathbf{x})_{1}$ are $j_{1}$ and $j_{2}$, then for each $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred at coordinate $j_{1}-1$ or $j_{1}$, and the other deletion occurred at coordinate $j_{2}-1$ or $j_{2}$.
II-A Related Results
Recently, there have been two notable developments in non-binary two-deletion correcting codes with low redundancy. In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for all even $q>2$. In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for all $q>2$. In Section V, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for all $q>2$.
Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available. A related result can be found in the work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively. For $t=2$ and $t_{L}=b$, their result provides a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. Their result was derived using the syndrome compression technique, which was later extended to syndrome compression with pre-coding in [20]. We will apply this extended technique to provide a $q$-ary code with redundancy at most $7\log n+O\left(\log n/\log\log n\right)$, for all $q\geq 2$.
For a subset $\mathcal{E}\subseteq\Sigma_{q}^{n}$ and a sequence $\mathbf{x}\in\mathcal{E}$, define
$\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)=\left\{\mathbf{y}\in\mathcal{E}~{}:~{}\mathbf{y}\neq\mathbf{x}\text{ and }\mathcal{B}_{2}^{b}\left(\mathbf{y}\right)\cap\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)\neq\emptyset\right\}.$
Lemma II.1
[13, 20, 37] Let $\mathcal{E}\subseteq\Sigma_{q}^{n}$ be a code and $N>\max\left\{\left|\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)\right|~{}:~{}\mathbf{x}\in\mathcal{E}\right\}$. Suppose that the function $f:\Sigma_{q}^{n}\rightarrow\{0,1\}^{R(n)}$ (where $R(n)$ is a function of $n$ and $R(n)\geq 2$) satisfies the following property:
  1. $(\textup{{P}}1)$
    if $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}\in\mathcal{N}_{\Sigma_{q}^{n}}\left(\mathbf{x}\right)$, then $f\left(\mathbf{x}\right)\neq f\left(\mathbf{y}\right)$.
Then there exists a function $\bar{f}:\mathcal{E}\rightarrow\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, computable in polynomial time, such that $\bar{f}\left(\mathbf{x}\right)\neq\bar{f}\left(\mathbf{y}\right)$ for any $\mathbf{x}\in\mathcal{E}$ and $\mathbf{y}\in\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$.
Let $\mathcal{E}$ be a $1$-$b$-burst correcting code and $\bar{f}$ be given in Lemma II.1. Then Lemma II.1 asserts that if $\mathbf{x},\mathbf{y}\in\mathcal{E}$ are distinct codewords and $\bar{f}(\mathbf{x})=\bar{f}(\mathbf{y})$, we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$. Therefore, for any $\mathbf{a}\in\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, the code =$\mathcal{E}^{\prime}=\left\{\mathbf{x}\in\mathcal{E}:\bar{f}(\mathbf{x})=a\right\}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, by the pigeonhole principle, there exists an $\mathbf{a}$ such that the redundancy of $\mathcal{E}^{\prime}$ is at most $\rho\left(\mathcal{E}^{\prime}\right)=\rho(\mathcal{E})+2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$ [^2] We select $\mathcal{E}$ to be a $1$-$b$-burst correcting code to obtain a better redundancy. If we take $\mathcal{E}=\Sigma_{q}^{n}$ we get a redundancy of $8\log n+o(\log n)$..
For the choice of $\mathcal{E}$, we have the following result.
Lemma II.2
[32, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n>b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
This lemma gives a $1$-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$. Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|\leq q^{b}n^{3}$. In fact, each codeword in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$ can be obtained in the following three steps:
  1. 1)
    Delete two substrings of length $b$ from $\mathbf{x}$, resulting in $\mathbf{z}^{(1)}$. There are at most $n^{2}$ possibilities for $\mathbf{z}^{(1)}$.
  2. 2)
    For each $\mathbf{z}^{(1)}$, insert a sequence of length $b$ into $\mathbf{z}^{(1)}$ and get a sequence $\mathbf{z}^{(2)}\in\Sigma_{q}^{n-b}$. For each $\mathbf{z}^{(1)}$, there are at most $q^{b}n$ possibilities for $\mathbf{z}^{(2)}$.
  3. 3)
    Insert a sequence of length $b$ into $\mathbf{z}^{(2)}$ to get a sequence $\mathbf{y}\in\Sigma_{q}^{n}$. Since $\mathcal{E}$ is a $1$-$b$-burst-deletion correcting code, for each $\mathbf{z}^{(2)}$, there is at most one $\mathbf{y}$ which is in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$.
Therefore, we can let $N=q^{b}n^{3}$. Now the redundancy of $\mathcal{E}^{\prime}$ is at most $7\log n+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$. To conclude our discussion, it remains to find an $f$ satisfying (P1) in Lemma II.1 such that $R(n)=O(\log n)$, which is given in Lemma II.5. The proof of Lemma II.5 is based on Lemmas II.3 and II.4.
The following lemma is a corollary of [17, Theorem 2].
Lemma II.3
[17, Theorem 2] For any integer $n\geq 3$, there exists a function $\xi:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{7\log n+O\left(1\right)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\xi\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}^{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
This result can be extended to arbitrary finite alphabets in the following way.
Lemma II.4
Suppose $q\geq 2$. There is a function $\xi_{1}:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{7\left\lceil\log q\right\rceil\log n+O_{q}(1)}$, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ and $\xi_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q}(1)$ is a constant dependent only on $q$.
Proof:
Any $\mathbf{x}\in\Sigma_{q}^{n}$ can be uniquely represented as an array
$M(\mathbf{x})\triangleq\begin{pmatrix}x_{1,1}&\cdots&x_{1,n}\\ \vdots&\cdots&\vdots\\ x_{\left\lceil\log q\right\rceil,1}&\cdots&x_{\left\lceil\log q\right\rceil,n}\end{pmatrix},$
where $x_{k,i}\in\{0,1\}$ be such that $x_{i}=\sum_{k=1}^{\left\lceil\log q\right\rceil}x_{k,i}2^{k-1}$ for all $1\leq i\leq n$.
Denote the $k$-th row of $M(\mathbf{x})$ by $M(\mathbf{x})_{k}$. Suppose $\mathbf{y}\in\mathcal{B}_{2}(\mathbf{x})$. It is clear that $M(\mathbf{y})_{k}\in\mathcal{B}_{2}\left(M(\mathbf{x})_{k}\right)$ for all $1\leq k\leq\left\lceil\log q\right\rceil$. Let $\xi(\cdot)$ be the function defined in Lemma II.3. For $\mathbf{x}\in\Sigma_{q}^{n}$, define $\xi_{1}(\mathbf{x})\triangleq\left(\xi\left(M(\mathbf{x})_{1}\right),\ldots,\xi\left(M(\mathbf{x})_{\left\lceil\log q\right\rceil}\right)\right)$. Then by Lemma II.3, given $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ and $\xi_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi\left(M(\mathbf{x})_{k}\right)$ is a binary vector of length $7\log n+O(1)$, we can see that $\xi_{1}(\mathbf{x})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log n+O_{q}(1)$. ∎
Lemma II.5
Suppose $b>1$, $n>2b$ and $q\geq 2$. There is a function $\psi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)}$, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$.
Proof:
Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{y})=A(\mathbf{y},n-b,b)$. Since $\mathbf{y}\in\mathcal{B}_{2}^{b}(\mathbf{x})$, we have $A(\mathbf{y})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. Define $\psi(\mathbf{x})\triangleq\left(\xi_{1}\left(A(\mathbf{x})_{1}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b}\right)\right)$. Then by Lemma II.4, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi_{1}(A(\mathbf{x})_{i})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log(n/b)+O_{q}(1)$, we can see that $\psi(\mathbf{x})$ is a binary vector of length $7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)$. ∎
Taking $f=\psi$ gives a function satisfying (P1) as needed.
Before proceeding to subsequent sections, we introduce two useful lemmas. The first one is a well-known result from number theory. The second one will be used in Section IV-B.
Lemma II.6 (Bertrand’s Postulate/Bertrand–Chebyshev theorem)
For every integer $n\geq 2$, there is always at least one prime $p$ such that $n\leq p<2n$.
Let $N\geq 2$ be an integer. Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$. Let ???? denote an unknown symbol (not in $[0,N-1]$). If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ be such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for all $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two). For our purpose in Section IV-B, assume $j\geq i+2$. For a sequence $\mathbf{x}$ over the alphabet $[0,N-1]$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,2)$ and $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
Lemma II.7
For any $0\leq a_{1},a_{2}<2N$ and $0\leq b<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
  • (C1)
    $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\equiv a_{1}\pmod{2N}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\equiv a_{2}\pmod{2N}$;
  • (C2)
    $W(\mathbf{x})\equiv b\pmod{nN^{2}}$, where $W(\mathbf{x})=\textup{Syn}\left(A(\mathbf{x})_{1}\right)+(2N-1)\cdot\textup{Syn}\left(A(\mathbf{x})_{2}\right)$.
Then the code $\mathcal{C}$ can correct two bursts (of length two) of erasures. In particular, there is a function $\varphi:[0,N-1]^{n}\rightarrow\Sigma_{2}^{\log n+4\log N+2}$, efficiently computable, such that for any $\mathbf{x}\in[0,N-1]^{n}$, given $\varphi(\mathbf{x})$, we can efficiently and uniquely recover $\mathbf{x}$ from $\mathbf{y}$, where $\mathbf{y}$ is any given sequence obtained from $\mathbf{x}$ by two bursts of erasures.
Proof:
Firstly, if the correctness of the code $\mathcal{C}$ is proved, we can define
$\varphi(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\pmod{2N},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\pmod{2N},W(\mathbf{x})\pmod{nN^{2}}\right).$
Here, we view $\varphi(\mathbf{x})$ as a binary vector. There are at most $(2N)^{2}\cdot(nN^{2})$ values of $\varphi(\mathbf{x})$. As a result, the length of $\varphi(\mathbf{x})$, when viewed as a binary vector, is at most $\log n+4\log N+2$. It remains to prove the correctness of $\mathcal{C}$.
Suppose $\mathbf{y}$ is obtained from a codeword $\mathbf{x}\in\mathcal{C}$ by two bursts of erasures. Then $A(\mathbf{y})_{i}$ is obtained from $A(\mathbf{x})_{i}$ by two erasures for each $i=1,2$. Denote the error positions in $A(\mathbf{y})_{1}$ are $i_{1},i_{2}$, and the error positions in $A(\mathbf{y})_{2}$ are $i_{3},i_{4}$, where $i_{1}<i_{2}$ and $i_{3}\leq i_{4}$. Note that $i_{1},i_{2},i_{3}$ and $i_{4}$ are known to us. Clearly, we have $i_{3}\in\{i_{1},i_{1}-1\}$ and $i_{4}\in\{i_{2},i_{2}-1\}$. This implies that $i_{4}-i_{3}\in\{i_{2}-i_{1},i_{2}-i_{1}-1,i_{2}-i_{1}+1\}$.
Next, we describe how to decode $\mathbf{x}$ from $\mathbf{y}$. Clearly, it is sufficient to recover the values of $A(\mathbf{x})_{1,i_{1}}$, $A(\mathbf{x})_{1,i_{2}}$, $A(\mathbf{x})_{2,i_{3}}$ and $A(\mathbf{x})_{2,i_{4}}$. Let
$\delta_{1}=\left(a_{1}-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\right)\pmod{2N},$ (1)
$\delta_{2}=\left(a_{2}-\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\right)\pmod{2N}.$
Since $0\leq A(\mathbf{x})_{1,i_{1}}+A(\mathbf{x})_{1,i_{2}},A(\mathbf{x})_{2,i_{3}}+A(\mathbf{x})_{2,i_{4}}<2N$ and $0\leq\delta_{1},\delta_{2}<2N$, it follows from Condition (C1) and Equation 1 that $\delta_{1}=A(\mathbf{x})_{1,i_{1}}+A(\mathbf{x})_{1,i_{2}}$ and $\delta_{2}=A(\mathbf{x})_{2,i_{3}}+A(\mathbf{x})_{2,i_{4}}$.
For simpler notations, denote $\alpha_{1}=A(\mathbf{x})_{1,i_{1}}$ and $\alpha_{2}=A(\mathbf{x})_{2,i_{3}}$. Then we have $A(\mathbf{x})_{1,i_{2}}=\delta_{1}-\alpha_{1}$ and $A(\mathbf{x})_{2,i_{4}}=\delta_{2}-\alpha_{2}$. Therefore, it remains to obtain the values of $\alpha_{1}$ and $\alpha_{2}$. To that end, let
$\Delta=\left(b-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{2,j}\right)\pmod{nN^{2}}.$ (2)
Since each term in the right-hand side of (2) is known, we can obtain the value of $\Delta$. Furthermore, by (2), Condition (C2), and the relationship between $A(\mathbf{x})$ and $A(\mathbf{y})$, it holds that
$\Delta$ $=\left(W(\mathbf{x})-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{2,j}\right)\pmod{nN^{2}}$ (3)
$=\left(i_{1}\alpha_{1}+i_{2}(\delta_{1}-\alpha_{1})+(2N-1)(i_{3}\alpha_{2}+i_{4}(\delta_{2}-\alpha_{2}))\right)\pmod{nN^{2}}.$
Let $\Delta^{\prime}=\left(i_{2}\delta_{1}+(2N-1)i_{4}\delta_{2}-\Delta\right)\pmod{nN^{2}}$, which can be calculated since values of $\delta_{1}$, $\delta_{2}$ and $\Delta$ are known to us. Then it follows from Equation 3 that
$(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}\equiv\Delta^{\prime}\pmod{nN^{2}}.$ (4)
Since $i_{1}<i_{2}\leq\left\lceil n/2\right\rceil$ and $i_{4}-i_{3}\leq\left\lceil n/2\right\rceil$, we have $21\leq i_{2}-i_{1},i_{4}-i_{3}\leq n/$ 2. Combining this with the fact $10\leq\alpha_{1},\alpha_{2}\leq N-$ 1, we conclude that $0\leq(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}<nN^{2}$. Now Equation 4 implies
$(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}=\Delta^{\prime}.$ (5)
Recall that $i_{4}-i_{3}\in\{i_{2}-i_{1}-1,i_{2}-i_{1},i_{2}-i_{1}+1\}$. By definition, the two bursts of erasures in $\mathbf{x}$ do not overlap. It follows that $i_{2}\geq i_{1}+2$ if $i_{4}-i_{3}=i_{2}-i_{1}-1$. Therefore, we always have $i_{4}-i_{3}\geq 1$. This implies $(2N-1)(i_{4}-i_{3})>1$. Then it follows from Equation 5 that
$(i_{2}-i_{1})\alpha_{1}\equiv\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}.$ (6)
Note that $10\leq\alpha_{1}\leq N-$ 1. When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it it easy to see that $0\leq(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. When $i_{4}-i_{3}=i_{2}-i_{1}-1$, since $i_{2}-i_{1}\geq 2$, we have $(2N-1)(i_{4}-i_{3})-(i_{2}-i_{1})\alpha_{1}=(i_{2}-i_{1})(2N-1-\alpha_{1})-(2N-1)\geq 2N-1-2\alpha_{1}>0$ and hence $(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. Now it follows from Equation 6 that $(i_{2}-i_{1})\alpha_{1}=\Delta^{\prime\prime}$, where $\Delta^{\prime\prime}=\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}$. From this, we get $\alpha_{1}=\frac{\Delta^{\prime\prime}}{(i_{2}-i_{1})}$. Then by Equation 5, we have $\alpha_{2}=\frac{\Delta^{\prime}-(i_{2}-i_{1})\alpha_{1}}{(2N-1)(i_{4}-i_{3})}$. Now the proof is completed. ∎
III Bounds
We could not find any existing upper or lower bounds on the maximum size of a $2$-$b$-burst-deletion code. In this section, we will derive these bounds.
Let $M_{q,n,b}$ be the maximum size of a $2$-$b$-burst-deletion correcting code in $\Sigma_{q}^{n}$, where $n>2b$.
Theorem III.1
$M_{n,q,b}\geq\frac{q^{n-2b}}{\binom{n}{2}^{2}}$.
Proof:
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. A subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [38, page 100, Theorem 1] that
$M_{q,n,b}\geq\sum_{\mathbf{x}\in\Sigma_{q}^{n}}\frac{1}{d(\mathbf{x})+1}.$ (7)
For $\mathbf{x}\in\Sigma_{q}^{n}$, by II.1, we conclude that each $\mathbf{y}$ (including $\mathbf{x}$ itself) with $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$ can be obtained as follows: 1) deleting two non-overlap substrings of length $b$ from $\mathbf{x}$, resulting in a sequence $\mathbf{z}\in\Sigma_{q}^{n-2b}$; 2) inserting two sequences of length $b$ into $\mathbf{z}$. Therefore, we have $d(\mathbf{x})+1\leq\binom{n}{2}^{2}q^{2b}$. ∎
In [39, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
Theorem III.2
For $q\geq 2$, let $f(q)=\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. When $n\geq 30$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
$M_{q,n,b}\leq\left(\frac{3b^{2}}{q^{2b-2}(q-1)^{2}}+\frac{(1.121)^{3b}}{n}\right)\frac{q^{n}}{n^{2}}.$
Proof:
Define $m=n/b-1$. Let $\mathcal{C}\subseteq\Sigma_{q}^{n}$ be a $2$-$b$-busrt-deletion correcting code. Set $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and $t=\left(1-\frac{1}{q}-\epsilon\right)m$. We partition $\mathcal{C}$ into two disjoint subsets: $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$, where $\mathcal{C}_{1}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\geq t+2\text{ for some }1\leq i\leq b\right\}$ and $\mathcal{C}_{2}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\leq t+1\text{ for all }1\leq i\leq b\right\}$ (recall that $r(\cdot)$ denotes the number of runs). To derive an upper bound of $\left|\mathcal{C}\right|$, it is sufficient to upper bound $\left|\mathcal{C}_{1}\right|$ and $\left|\mathcal{C}_{2}\right|$.
For any $\mathbf{x}\in\mathcal{C}$, define $\mathcal{A}_{2}^{b}(\mathbf{x})=\left\{A(\mathbf{x}^{\prime}):\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})\right\}$. Let $A(\ell,k)$ be the array obtained by deleting the $\ell$-th and the $k$-th columns of $A(\mathbf{x})$ for $1\leq\ell\neq k\leq n/b$. It is clear that $A(\ell,k)\in\mathcal{A}_{2}^{b}(\mathbf{x})$ and
$\cup_{1\leq\ell\neq k\leq\frac{n}{b}}\left\{A(\ell,k)_{i}\right\}=\mathcal{B}_{2}(A(\mathbf{x})_{i}),$
for all $1\leq i\leq b$. By [40, eq. (11)], we know that
$\binom{r-1}{2}\leq\left|\mathcal{B}_{2}(\mathbf{v})\right|\leq\binom{r+1}{2}$
for any sequence $\mathbf{v}\in\Sigma_{q}^{n}$ with exactly $r$ runs. Since $\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|$ and $\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|=\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|$ for all $\mathbf{x}$, it follows that
$\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|\geq\binom{\max_{1\leq i\leq b}\left\{r\left(A(\mathbf{x})_{i}\right)\right\}-1}{2}\geq\binom{t+1}{2},$ (8)
for all $\mathbf{x}\in\mathcal{C}_{1}$. Since $\mathcal{C}$ is a $2$-$b$-burst-deletion correcting code, we know that $\mathcal{C}_{1}$ is also a $2$-$b$-burst-deletion correcting code. So we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for all distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}_{1}$. Then it follows from Equation 8 that $\left|\mathcal{C}_{1}\right|\binom{t+1}{2}\leq\sum_{\mathbf{x}\in\mathcal{C}_{1}}\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\leq q^{n-2b}$ and hence
$\left|\mathcal{C}_{1}\right|\leq\frac{q^{n-2b}}{\binom{t+1}{2}}\leq\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{t^{2}}$ $=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{\left(1-1/q-\epsilon\right)^{2}\left(n/b-1\right)^{2}}$ (9)
$=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2b^{2}}{(1-1/q-\epsilon)^{2}(1-b/n)^{2}}$
$=\frac{2b^{2}q^{n+2}}{n^{2}q^{2b}(q-1)^{2}}\cdot\frac{1}{\left(1-\frac{q}{q-1}\epsilon\right)^{2}\left(1-b/n\right)^{2}}$
$\leq\frac{3b^{2}q^{n+2}}{n^{2}q^{2b}(q-1)^{2}}$
as long as $\left(1-\frac{q}{q-1}\epsilon\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, which is possible when $n$ is sufficiently large.
Next, we proceed to upper bound $\left|\mathcal{C}_{2}\right|$. To that end, define a set
$\mathcal{C}^{\prime}\triangleq\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(b)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in\Sigma_{q}^{n/b},\;\forall 1\leq i\leq b\\ r\left(\mathbf{x}^{(i)}\right)\leq t+1,\;\forall 1\leq i\leq b\end{array}\right\}.$
Since $r\left(A(\mathbf{x})_{i}\right)\leq t+1$ for any $\mathbf{x}\in\mathcal{C}_{2}$ and $1\leq i\leq b$, we conclude that $\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|$. So it suffices to estimate an upper bound of $\left|\mathcal{C}^{\prime}\right|$.
Since $m=n/b-1$, by definition of $\mathcal{C}^{\prime}$, we have
$\left|\mathcal{C}^{\prime}\right|$ =|$=\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})\leq t+1\right\}\right|^{b}$ (10)
$=\left|\cup_{j=1}^{t+1}\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|^{b}$
$=\left(\sum_{j=1}^{m+1}\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|\right)^{b}$
=$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
where (a) follows from the well-known result (see the proof of [41, Theorem 3.1]):
|$\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|=\binom{m}{j-1}q(q-1)^{j-1}.$
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [42, Proposition 3.3.3] that
$\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\leq q^{mH_{q}(t/m)}=q^{mH_{q}(1-1/q-\epsilon)},$ (11)
where $H_{q}(x)\triangleq x\log_{q}(q-1)-x\log_{q}x-(1-x)\log_{q}(1-x)$ is the $q$-ary entropy function. Here, $\log_{q}x=\frac{\log x}{\log q}$ for any positive real number $x$.
If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. Now by Lemma .1, we have
$H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
Now it follows from Equations 10 and 11 that
$\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}\overset{(b)}{<}(1.121)^{3b}\frac{q^{n}}{n^{3}},$ (12)
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 9 and 12. ∎
Then next corollary is a direct consequence of Theorems III.1 and III.2.
Corollary III.1
Suppose that $q$ and $b$ are fixed. Let $f(q)$ be defined in Theorem III.2. When $n\geq\max\left\{30,\frac{(1.21)^{3b}q^{2b-2}(q-1)^{2}}{b^{2}}\right\}$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
$2\log n+(2b-2)\log q+2\log(q-1)-2\log b-2\leq\rho(\mathcal{C})\leq 4\log n+2b\log q,$
where $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code with maximum size.
IV Codes For Correcting Two $b$-Burst-Deletions
In this section, we construct $q$-ary $2$-$b$-burst-deletion correcting codes with redundancy at most $5\log n+O(\log\log n)$, for any $q\geq 2$ and $b>1$. Our idea is to first locate positions of deletions in short intervals (which is accomplished in Section IV-A, Lemma IV.5) and then correct errors in these intervals (which is accomplished in Sections IV-B and IV-C).
IV-A Approximately determine positions of deletions
Definition IV.1 (regularity)
A sequence $\mathbf{x}\in\Sigma_{2}^{n}$ is said to be $d$-regular if each substring of $\mathbf{x}$ of length at least $d\log n$ contains both $00$ and $11$.
For the number of regular sequences, we have the following lemma.
Lemma IV.1
[19, Lemma 11] The number of $d$-regular sequences of length $n$ is at least $2^{n-1}$, as long as $d\geq 7$ and $n$ be such that $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{0.15d-1}\geq 12$. In particular, when $d=7$, it suffices to require $n\geq 9$.
Lemma IV.2
[19, Theorem 7] Suppose $d\geq 7$. There exists a function $\eta:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{4\log n+10\log\log n+O_{d}(1)}$, computable in linear time, such that for any $d$-regular sequence $\mathbf{x}\in\Sigma_{2}^{n}$, given $\eta\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{d}(1)$ denotes a constant depending only on $d$.
If the first row $A(\mathbf{x})_{1}$ is regular, then by correcting the two deletions in $A(\mathbf{x})_{1}$, we can approximate the location of the two deletions, as ensured by the following lemma.
Lemma IV.3
[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma_{2}^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$. Then we can
  1. $(1)$
    either find distinct runs $\mathbf{x}_{J_{1}}$ and $\mathbf{x}_{J_{2}}$ of $\mathbf{x}$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, such that $i_{1}\in J_{1}$ and $i_{2}\in J_{2}$.
  2. $(2)$
    or find an interval $J\subseteq[n]$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, of length at most $3d\log n$ such that $i_{1},i_{2}\in J$.
We briefly explain why Lemma IV.3 is helpful. Suppose $\mathbf{x}\in\Sigma_{2}^{n}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$. Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},n-2b,b)$. Then $A(\mathbf{x}^{\prime})_{i}\in\mathcal{B}_{2}(A(\mathbf{x})_{i})$ for all $1\leq i\leq 2$ If $A(\mathbf{x})_{1}$ is $d$-regular (where $d\geq 7$) and $\eta(A(\mathbf{x})_{1})$ is given, Lemma IV.2 ensures that we can decode $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ and Lemma IV.3 ensures that error positions in $A(\mathbf{x})_{1}$ can be approximately determined in one or two short intervals. Then by II.2, we can approximate error positions in $A(\mathbf{x})_{i}$ within one or two short intervals, for all $2\leq i\leq b$.
However, the above analysis is valid only when $\mathbf{x}$ is a binary sequence. Since we deal with binary and non-binary cases in a unified way, it is convenient to generalize the notion of regularity to general alphabets. To that end, we associate with any $q$-ary sequence a binary sequence.
Let $q\geq 2$. For any $x\in\Sigma_{q}$, we can uniquely write it as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<\left\lceil q/2\right\rceil$. Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$. Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$ and when $q=2$, we have $u(\mathbf{x})=\mathbf{x}$. It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
Remark IV.1
Although any $x\in\Sigma_{q}$ can be decomposed as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, it might exist some $u\in\{0,1\}$ and $0\leq v<\left\lceil q/2\right\rceil$ such that $\left\lceil q/2\right\rceil u+v\notin\Sigma_{q}$. In fact, if $q$ is even, then for any $u\in\{0,1\}$ and any $0\leq v<\left\lceil q/2\right\rceil$, we have $\left\lceil q/2\right\rceil u+v\in\Sigma_{q}$. But when $q$ is odd, we have $\left\lceil q/2\right\rceil+\left\lceil q/2\right\rceil-1=q\notin\Sigma_{q}$.
Definition IV.2
A sequence $\mathbf{x}\in\Sigma_{q}^{n}$ is said to be $d$-regular if $u(\mathbf{x})$ is $d$-regular. In other words, $\mathbf{x}$ is $d$-regular if and only if any substring of $\mathbf{x}$ of length at least $d\log n$ contains two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$.
[…]
Lemma IV.4
Let $q\geq 2$.
  • If $q$ is even, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n}/2\geq q^{n-1}$ for all $d\geq 7$ and $n$ such that $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{0.15d-1}\geq 12$. In particular, when $d=7$, it suffices to require $n\geq 9$.
  • If $q$ is odd, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n-1}$ for all $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. Note that when $d\geq 10$, we have $-1-\frac{d}{2}\log(0.87)>0$ and hence this condition could be satisfied when $n$ is sufficiently large. In particular, when $d=10$, it suffices to require $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
Proof:
Suppose first that $q$ is even. Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. Then it follows from Lemma IV.1 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
Now suppose that $q\geq 3$ is odd. In this case, the previous argument does not hold. This is because when $q$ is odd, $u_{x}=1$ and $v_{x}=(q-1)/2$, we have $\left\lceil q/2\right\rceil u_{x}+(q-1)/2=q\notin\Sigma_{q}$. Fortunately, we can apply similar ideas in proofs of [19, Lemma 11] and [22, Lemma 5] to derive our lower bound on $\left|\mathcal{R}_{q,n,d}\right|$. For $m\geq 2$, define
$S_{m}^{L}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }0\leq v_{i},v_{i+1}\leq(q-1)/2\right\},$
$S_{m}^{U}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }(q+1)/2\leq v_{i},v_{i+1}<q\right\}.$
We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$. By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}\right)^{3}=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$. Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$. For $\mathbf{v}\in S_{m}^{L}$, it must holds that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$. So it follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
For $\mathbf{v}\in\Sigma_{q}^{m}$, let $\overline{\mathbf{v}}\triangleq(q-v_{1})\cdots(q-v_{m})$. It is easy to verify that if $\mathbf{v}\in S_{m}^{U}$, then $\overline{\mathbf{v}}\in S_{m}^{L}$. Then it follows that $\left|S_{m}^{U}\right|\leq\left|S_{m}^{L}\right|<(0.87q)^{m}$.
Let $m=\left\lfloor\frac{d}{2}\log n\right\rfloor$ and
$Q=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\exists i,j\text{ such that }v_{i},v_{i+1}<(q+1)/2\text{ and }v_{j}.v_{j+1}\geq(q+1)/2\right\}.$
Then we have $\left|Q\right|\geq q^{m}-\left|S_{m}^{L}\right|-\left|S_{m}^{U}\right|>q^{m}-2(0.87q)^{m}$. Now let $k=\left\lfloor n/m\right\rfloor$ and define
=$\mathcal{R}^{\prime}=\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(k)}\mathbf{x}^{(k+1)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 1\leq i\leq k\\ \mathbf{x}^{(k+1)}\in\Sigma_{q}^{n-k\left\lfloor n/m\right\rfloor}.\end{array}\right\}$
It is straightforward to see that $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$ and thus
$\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}^{\prime}\right|$ $=\left|Q\right|^{k}q^{n-km}$
$>\left(q^{m}-2(0.87q)^{m}\right)^{k}q^{n-km}$
$=q^{n}\left(1-2(0.87)^{m}\right)^{k}$
$\overset{(a)}{\geq}q^{n}\left(1-2k(0.87)^{m}\right)$
$=q^{n}\left(1-2\left\lfloor\frac{n}{m}\right\rfloor 2^{\left\lfloor\frac{d}{2}\log n\right\rfloor\log(0.87)}\right)$
$\geq q^{n}\left(1-2\frac{n}{m}2^{\left\lfloor\frac{d}{2}\log n\right\rfloor\log(0.87)}\right)$
$\overset{(b)}{\geq}q^{n}\left(1-2\frac{n}{m}2^{\frac{d}{2}\log n\cdot\log(0.87)-\log(0.87)}\right)$
$=q^{n}\left(1-\frac{2}{0.87\left\lfloor\frac{d}{2}\log n\right\rfloor}n^{1+\frac{d}{2}\log(0.87)}\right)$
$\overset{(c)}{\geq}q^{n-1}.$
Here, (a) follows from the fact that $(1+x)^{r}\geq 1+rx$ for any integer $r\geq 1$ and any real number $x\geq-1$; (b) follows from the fact that $\log(0.87)<0$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor\geq\frac{d}{2}\log n-1$; (c) follow from the fact that $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. ∎
Notations: In the rest of this section, for any $q$-ary sequence $\mathbf{x}$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,b)$ and $U(\mathbf{x})=A(u(\mathbf{x}),\left|\mathbf{x}\right|,b)$. Denote the $i$-th row of $A(\mathbf{x})$ and $U(\mathbf{x})$ by $A(\mathbf{x})_{i}$ and $U(\mathbf{x})_{i}$, respectively. It is clear that $U(\mathbf{x})_{i}=u\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let
$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\},$
which is a subset of $[n]$. Then we have $A(\mathbf{x})_{1}=\mathbf{x}_{I_{n,b}}$ and $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
Recall that $r\left(U(\mathbf{x})_{1}\right)$ is the number of runs in $U(\mathbf{x})_{1}$. For $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$, let $U(\mathbf{x})_{1,I_{j}}$ be the $j$-th run. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$. By definition, $I_{j}$ is an interval in $[n/b]$. For all $j\geq 1$, we associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$. In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$. In other words, $I_{j}^{L}$ is the interval of length at most $b-1$ in $[n]$ to the left of $I_{j}^{\prime}$. See Figure 1 for an illustration. Note that $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$ are dependent on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
[…]
(a) $U(\mathbf{x})$: $I_{j}$ in $\left[n/b\right]$
[…]
(b) $\mathbf{x}$: $I_{j}^{\prime}$ and $I_{j}^{L}$ in $\left[n\right]$
Figure 1: Illustration of relationship between $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$.
Lemma IV.5
Let $q\geq 2$, $b>1$ and $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular. Let $\eta(\cdot)$ be the function defined in Lemma IV.2. Then given $\eta\left(U(\mathbf{x})_{1}\right)$, we can
  1. $(1)$
    either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$;
  2. $(2)$
    or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$.
Proof:
Since $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, we have $u(\mathbf{x}^{\prime})=u(\mathbf{x})_{[n]\setminus(D_{1}\cup D_{2})}$. It follows that $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$. Suppose that $U(\mathbf{x}^{\prime})_{1}$ is obtained from $U(\mathbf{x})_{1}$ by deleting $U(\mathbf{x})_{1,k_{1}}$ and $U(\mathbf{x})_{1,k_{2}}$. Recall that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. So $U(\mathbf{x}^{\prime})_{1}$ is obtained from $u(\mathbf{x})_{I_{n,b}}$ by deleting $u(\mathbf{x})_{1+(k_{1}-1)b}$ and $u(\mathbf{x})_{1+(k_{2}-1)b}$.
Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.2, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$. In addition, by Lemma IV.3, we can
  • either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $k_{1}\in I_{j_{1}}$ and $k_{2}\in I_{j_{2}}$;
  • or find an interval $[c,d]\subseteq[n/b]$ of length at most $3d\log(n/b)$ such that $k_{1},k_{2}\in[c,d]$.
By II.2, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. For the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$. Now the proof is completed. ∎
From now on, we assume that $\mathbf{x}\in\Sigma_{q}^{n}$ and $A(\mathbf{x})_{1}$ is $d$-regular, where $d$ will be specified later. Furthermore, suppose $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$. Then by II.2, we have $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$. Since $U(\mathbf{x})_{1}$ is regular, Lemma IV.2 ensures that when given $\eta\left(U(\mathbf{x})_{1}\right)$, one can uniquely recover $U(\mathbf{x})_{1}$ from $U\left(\mathbf{x}^{\prime}\right)_{1}$. Then by Lemma IV.3, we can know the approximate positions of the two deletions in $U(\mathbf{x})_{1}$. According to Lemma IV.5, we can approximately determine the positions of the two bursts in $\mathbf{x}$. Based on the two cases in Lemma IV.5, we split our discussion into Section IV-B and Section IV-C. Our codes for correcting two $b$-burst-deletions are presented in Theorem IV.1, whose proof is divided into Lemmas IV.6 and IV.7.
Before proceeding, we need the following observation.
Observation IV.1
Let $m>b$. Suppose $\mathbf{u}\in\Sigma_{q}^{m}$ and $\mathbf{v}\in\mathcal{B}_{1}^{b}(\mathbf{u})$. Then for any $b\leq i\leq m-b$, we have $\mathbf{v}_{[1,i-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[1,i]}\right)$ and $\mathbf{v}_{[i+1,m-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[i+1,m]}\right)$.
IV-B When Case (1) in Lemma IV.5 happens
In this case, we already found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Note that $I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}=\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]$ and $I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}=\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]$. Then it follows that
$x_{k}=\begin{cases}x^{\prime}_{k},\text{ if }k<p_{j_{1}-1}b-b+2,\\ x^{\prime}_{k-b},\text{ if }p_{j_{1}}b<k<p_{j_{2}-1}b-b+2,\\ x^{\prime}_{k-2b},\text{ if }p_{j_{2}}b<k\leq n.\end{cases}$ (13)
In particular, $\mathbf{x}_{I_{j}^{\prime}}$ is know to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, to decode $\mathbf{x}$ from $\mathbf{x}^{\prime}$, it is sufficient to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
In the rest of this subsection, it is always assumed that $j_{2}-j_{1}\geq 2$, since otherwise, we have $\left|I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\cup I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\right|\leq 2bd\log(n/b)+b-1$, and therefore, this situation is covered by case (2) in Lemma IV.5 (or equivalently, case (2) in Lemma IV.3).
Since $j_{2}-j_{1}\geq 2$, deletions in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ do not affect $\mathbf{x}_{I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}}$. As a result, (13) implies $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}}\right)$. Applying IV.1 to $\mathbf{u}=\mathbf{x}_{I_{j_{1}-1}\cup I_{j_{1}}}$, $\mathbf{v}=\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$, $m=\left(p_{j_{1}}-p_{j_{1}-2}\right)b$ and $i=\left(p_{j_{1}-1}-p_{j_{1}-2}\right)b$, we conclude that
$\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}-1}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right),$ (14)
$\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right).$
Similarly, we have
$\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}-1}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right),$ (15)
$\mathbf{x}^{\prime}_{\left[(p_{j_{2}-1}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right).$
Recall that the range of $j_{1}$ is $\left[1,r(U(\mathbf{x})_{1})\right]$. To make (14) valid when $j_{1}=1$, we let $\mathbf{x}_{I_{0}^{\prime}}=0^{b}$ (the sequence consisting of $b$ symbols 00) and $\mathbf{x}^{\prime}_{\left[p_{-2}b+1,-b\right]}$ be an empty sequence.
By (14) and (15), to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$, we need to apply to each $\mathbf{x}_{I_{j}^{\prime}}$ a code for correcting single $b$-burst-deletion. Let $\phi(\cdot)$ be the function defined in Lemma II.2. Then by Lemma II.2, our goal boils down to recovering the values of $\phi\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $A(\mathbf{x})_{1}$ is $d$-regular, we have $\left|I_{j}^{\prime}\right|\leq db\log(n/b)$ for all $0\leq j\leq r(U(\mathbf{x})_{1})$. Therefore, each $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $[0,N-1]$, where $N=2^{\log\log(n)+O_{q,d,b}(1)}$. Here, $O_{q,d,b}(1)$ denotes a constant dependent only on $q,d$ and $b$.
Define $\overline{\phi}(\mathbf{x})=\phi\left(\mathbf{x}_{I_{0}^{\prime}}\right)\phi\left(\mathbf{x}_{I_{1}^{\prime}}\right)\cdots\phi\left(\mathbf{x}_{I_{r(U(\mathbf{x})_{1})}^{\prime}}\right)\in[0,N-1]^{r(U(\mathbf{x})_{1})+1}$. According to (13), when given $\mathbf{x}^{\prime}$, the value of $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, sequence $\overline{\phi}(\mathbf{x})$ suffered two bursts (of length two) of erasures. To correct errors in $\overline{\phi}(\mathbf{x})$, let $\varphi(\cdot)$ be the function defined in Lemma II.7 and define
$f(\mathbf{x})=\varphi\left(\overline{\phi}(\mathbf{x})\right).$
Recall that we have $r(U(\mathbf{x})_{1})\leq n/b$. Then By Lemma II.7, $f(\mathbf{x})$ can be viewed as an integer in $[0,N_{1}-1]$, where $N_{1}=2^{\log n+4\log\log n+O_{q,d,b}(1)}$.
The following lemma follows from the above analysis, Lemmas II.2 and II.7.
Lemma IV.6
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f(\mathbf{x})$, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
IV-C When Case (2) in Lemma IV.5 happens
Recall that case (2) in Lemma IV.5 corresponds to case (2) in Lemma IV.3. Therefore, for case (2), instead of looking at $\mathbf{x}$ and $\mathbf{x}^{\prime}$, we look at $A(\mathbf{x})$ and $A(\mathbf{x}^{\prime})$. In this case, we have already decoded $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ by using $\eta\left(U(\mathbf{x})_{1}\right)$. Furthermore, by case (2) in Lemma IV.3, we found an interval $[c_{1},c_{2}]\subseteq[n/b]$ of length at most $3db\log(n/b)$, which contains the positions of the two deletions in $U(\mathbf{x})_{1}$ (and hence in $A(\mathbf{x})_{1}$). By II.2, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $[c_{1}-1,c_{2}]$, for all $2\leq i\leq b$. Therefore, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin[c_{1}-1,c_{2}]$.
Note that $\left|[c_{1}-1,c_{2}]\right|$=$\left|J\right|+1\leq 3db\log(n/b)+1$. Let $P=3db\log(n/b)+1$ and define
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/bP\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/bP\right\rceil-1.\end{cases}$
Since $c_{1}$ and $c_{2}$ are known, we can find some $j_{0}$ such that $[c_{1}-1,c_{2}]\subseteq J_{j_{0}}$. Suppose $J_{j_{0}}=[d_{1},d_{2}]$. According to above discussion, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin J_{j_{0}}$. To decode $A(\mathbf{x})_{i}$ from $A(\mathbf{x}^{\prime})_{i}$, it remains to recover $A(\mathbf{x})_{i,J_{j_{0}}}$.It is easy to see that $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i,J_{j_{0}}}\right)$. Therefore, we can apply Lemma II.4 to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$.
Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. For each $11\leq j\leq\left\lceil n/bP\right\rceil-$ 1, define
$g_{j}\left(\mathbf{x}\right)=\begin{cases}\left(\xi_{1}\left(A(\mathbf{x})_{1,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q>2;\\ \left(\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q=2.\end{cases}$
Note that in the above definition of $g_{j}(\mathbf{x})$, we distinguish the two cases where $q>2$ and $q=2$. This is because when $q=2$, we have $A(\mathbf{x})=U(\mathbf{x})$, and hence the first row of $A(\mathbf{x})$ was already recovered.
Since $\left|J_{j}\right|\leq 2P=6db\log\left(n/b\right)+2$ for all $j$, by Lemma II.4, each $\xi_{1}\left(A(\mathbf{x})_{i,J_{j}}\right)$ is a binary vector of length $7\left\lceil\log q\right\rceil\log\log n$ $+O_{q,d,b}(1)$. Therefore, each $g_{j}(\mathbf{x})$ is a binary vector of length at most $7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)$ (when $q>2$) or $7(b-1)\log\log n+O_{d,b}(1)$ (when $q=2$). Therefore, we can view $g_{j}(\mathbf{x})$ as an integer in $\left[0,N_{2}-1\right]$, where
$N_{2}=\begin{cases}2^{7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)},\mbox{ if }q>2,\\ 2^{7(b-1)\log\log n+O_{d,b}(1)},\mbox{ if }q=2.\\ \end{cases}$ (16)
For $s\in\{0,1\}$, let $L_{s}=\left\{1\leq j\leq\left\lceil n/bP\right\rceil-1:j\equiv s\pmod{2}\right\}$. For $s\in\{0,1\}$, define
$h^{(s)}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{s}}g_{j}(\mathbf{x})\pmod{N_{2}}.$ (17)
Lemma IV.7
For case (2) in Lemma IV.5 (and hence case (2) in Lemma IV.3), given $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$, we can uniquely recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $\mathbf{x}^{\prime}$, for each $2\leq i\leq b$. In particular, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
Proof:
Without loss of generality, assume $j_{0}\in L_{0}$. Then $g_{j}(\mathbf{x})$ is known to us for each $j\in L_{0}\setminus\{j_{0}\}$. By above discussion and Lemma II.4, to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ for all $2\leq i\leq b$, it is sufficient to obtain the value of $g_{j_{0}}(\mathbf{x})$. It follows from Equation 17, that
$g_{j_{0}}(\mathbf{x})\equiv\left(h^{(0)}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}g_{j}(\mathbf{x})\right)\pmod{N_{2}}.$
Since $0\leq g_{j_{0}}(\mathbf{x})<N_{2}$, we have
$g_{j_{0}}(\mathbf{x})=\left(h^{(0)}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}g_{j}(\mathbf{x})\right)\pmod{N_{2}}.$
Now the proof is completed. ∎
We are ready to present our main result in this section.
Theorem IV.1
Suppose $q\geq 2$, $n>2b$ and $b>1$. Let $N_{1}$, $N_{2}$, $f(\mathbf{x})$, $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$ be defined as above. Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$. For any $0\leq a<N_{0}$, $0\leq b<N_{1}$ and $0\leq c_{0},c_{1}<N_{2}$, define the code $\mathcal{C}_{1}$ as
$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ h^{(0)}(\mathbf{x})=c_{0},h^{(1)}(\mathbf{x})=c_{1}\end{array}\right\}.$
Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O_{q,b}(1)$:
  • $q>2$ is even, $d=7$ and $n\geq 9$;
  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
If $q=2$, $d=7$ and $n\geq 9$, the redundancy is upper bounded by $5\log n+14b\log\log n+O_{b}(1)$.
Proof:
Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma IV.5, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$. For the former case, Lemma IV.6 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. For the latter case, Lemma IV.7 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. Therefore, $\mathcal{C}_{1}$ can correct two $b$-burst-deletions.
By Lemma IV.4, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $A(\mathbf{x})_{1}$ is $d$-regular. It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, and $c_{1}$, such that
$\left|\mathcal{C}_{1}\right|\geq\frac{q^{n-1}}{N_{0}\cdot N_{1}\cdot(N_{2})^{2}},$
which implies by definition that
$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+\log N_{1}+2\log N_{2}+1$
$=5\log n+14\log\log n+2\log N_{2}+O_{q,b}(1).$
Now the claim for redundancy follows from (16). ∎
V Non-binary Two-deletion Correcting Codes
In this section, it is always assumed that $q>2$. We aim to give a new construction of $q$-ary two-deletion correcting codes. Recall the definition of $u(\mathbf{x})$ for any $\mathbf{x}\in\Sigma_{q}^{n}$. If $\mathbf{x}^{\prime}\in\mathcal{B}_{2}(\mathbf{x})$, then $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}(u(\mathbf{x}))$. So if $u(\mathbf{x})$ belongs to a binary two-deletion correcting code, we can recover $u(\mathbf{x})$ from $u(\mathbf{x}^{\prime})$. We further assume that $\mathbf{x}$ (and thus $u(\mathbf{x})$) is $d$-regular. Then Lemma IV.3 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions. This enables us to approximately determine error positions in $\mathbf{x}$. For the latter case, we need a $q$-ary two-deletion correcting code, which is given in Lemma II.4. For the former case, we need a $q$-ary single-deletion correcting code. We use the code given in [23]
Lemma V.1
[23] Suppose $q>2$ and $n>2$. There is a function $\mathsf{DVT}:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+\log q}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathsf{DVT}(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$.
Let $u(\mathbf{x})_{I_{j}}$ ($1\leq j\leq r\left(u(\mathbf{x})\right)$) be all the runs in $u(\mathbf{x})$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(u(\mathbf{x})\right)}=n$ and $p_{j-1}<p_{j}$ for all $j$. Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$. Therefore, by Lemma V.1, $\mathsf{DVT}\left(I_{j}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$. Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime. By Lemma II.6, we have $n\leq Q<2n$. Define
$f_{0}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{2N_{1}},$ (18)
$f_{1}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{Q}.$
The two functions $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$ can help to deal with case (1) in Lemma IV.3, as shown in the next lemma.
Lemma V.2
Suppose that the two deletions in $\mathbf{x}$ occurred in two known intervals $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$, respectively. Then given $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$, we can uniquely recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
Proof:
Notice that $x_{k}$ is known to us for all $k\notin I_{j_{1}}\cup I_{j_{2}}$. It remains to recover $\mathbf{x}_{I_{j_{1}}}$ and $\mathbf{x}_{I_{j_{2}}}$. It is easy to see that $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}+1,p_{j_{1}}-1\right]}\in\mathcal{B}_{1}\left(\mathbf{x}_{I_{j_{1}}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2},p_{j_{2}}-2\right]}\in\mathcal{B}_{1}\left(\mathbf{x}_{I_{j_{2}}}\right)$. By Lemma V.1, to recover $\mathbf{x}_{I_{j_{1}}}$ and $\mathbf{x}_{I_{j_{2}}}$, it is sufficient to know the values of $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)$ and $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)$. Note that $\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)$ is known to us for all $j\neq j_{1},j_{2}$. Let $\delta_{0}=\left(f_{0}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\right)\pmod{2N_{1}}$ and $\delta_{1}=\left(f_{1}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\right)\pmod{Q}$. Then it follows from Equation 18 that
$\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{0}\pmod{2N_{1}},$ (19)
$j_{1}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+j_{2}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{1}\pmod{Q}.$ (20)
Since $0\leq\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)<2N_{1}$ and $0\leq\delta_{0}<2N_{1}$, we can conclude from Equation 19 that $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)=\delta_{0}$. Combining this with Equation 20, we get
$\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{0}\pmod{Q},$ (21)
$j_{1}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+j_{2}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{1}\pmod{Q}.$
Since $1\leq j_{1}<j_{2}\leq n\leq Q$, we have $j_{1}\not\equiv j_{2}\pmod{Q}$. Therefore, system (21) has a unique solution in the field $\mathbb{F}_{Q}$: $\left(\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)\pmod{Q},\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\pmod{Q}\right)$. Since $0\leq\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right),\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)<N_{1}\leq Q$, we have $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)=\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)\pmod{Q}$ and $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)=\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\pmod{Q}$. Now the proof is completed. ∎
For case (2) in Lemma IV.3, we follow similar idea in Section IV-C. Let $P=3d\log n$ and
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/P\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/P\right\rceil-1.\end{cases}$
Since $\left|J_{1}\right|\leq 2P=6d\log n$, we can view each $\xi_{1}\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{7\left\lceil\log q\right\rceil\log(6d\log n)+O_{q}(1)}$ (see Lemma II.4). For $s\in\{0,1\}$, let $K_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$ and define
$h^{(s)}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{s}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}}.\$ (22)
Lemma V.3
Suppose we are in case (2) in Lemma IV.3. Given $h^{(0)}$ and $h^{(1)}$, we can uniquely recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
The proof is similar to that of Lemma IV.7, and thus omitted.
Our $q$-ary two-deletion correcting code is given in the next theorem.
Theorem V.1
Suppose $q>2$. Let $N_{1},N_{2},Q$, $f_{0}(\mathbf{x})$, $f_{1}(\mathbf{x})$, $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{0}<2N_{1}$, $0\leq b_{1}<Q$, and $0\leq c_{0},c_{1}<N_{2}$, define the code $\mathcal{C}_{2}$ as
$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}\text{ is }d\text{-regular},\\ \eta(u(\mathbf{x}))=a,\\ f_{s}(\mathbf{x})=b_{s}\text{ for }s\in\{0,1\},\\ h^{(0)}(\mathbf{x})=c_{0},h^{(1)}(\mathbf{x})=c_{1}\end{array}\right\}.$
Then $\mathcal{C}_{2}$ is a two-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{0}$, $b_{1}$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
  • $q$ is even, $d=7$ and $n\geq 9$;
  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
The proof of Theorem V.1 follows directly from Lemmas V.2 and V.3. The claim for redundancy follows from Lemma IV.4 and Lemma II.6.
VI Conclusion
In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $5\log n+O(\log\log n)$ for all $b\geq 2$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique. Inspired by these ideas, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
A more complex problem is to construct codes capable of correcting at most two bursts of deletions, where each burst has a length at most $b$. Our technique fails in this setup since we can not even know the number of bursts or the length of each burst occurring in a received sequence. We leave this problem for future research.
The following Lemma .1 is derived by slightly modifying the proof of [42, Proposition 3.3.7]. Before showing its proof, we list some facts.
Let $\ln(\cdot)$ be the natural logarithm. Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}-\cdots$.
Fact .1
For all $10\leq x<$ 1, we have $\ln(1-x)\leq-x-\frac{x^{2}}{2}$ and $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$. When $0\leq x\leq\frac{1}{2}$, we have $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
Proof:
When $10\leq x<$ 1, we have $\ln(1-x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{(-x)^{i}}{i}=-\sum_{i=1}^{\infty}\frac{x^{i}}{i}\leq-x-\frac{x^{2}}{2}$. Since $-\ln(1+x)=-x+\frac{x^{2}}{2}-\left(\ln(1+x)-x+\frac{x^{2}}{2}\right)$ and $\ln(1+x)-x+\frac{x^{2}}{2}\geq 0$ for all $x\geq 0$, we have $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$. At last, it is easy to see that $-\ln(1-x)=x+\frac{x^{2}}{2}+x^{3}\left(\frac{1}{3}+\frac{x}{4}+\frac{x^{2}}{5}+\cdots\right)\leq x+\frac{x^{2}}{2}+x^{3}(1+x+x^{2}+\cdots)=x+\frac{x^{2}}{2}+\frac{x^{3}}{1-x}$. When $x\leq\frac{1}{2}$, we have $1-x\geq\frac{1}{2}$ and thus $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$. ∎
Lemma .1
Suppose that $q\geq 2$ is fixed and $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. Then it holds that
$H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
Proof:
It is easy to verify that
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $=-\left(1-\frac{1}{q}-\epsilon\right)\log_{q}\left(\frac{1-1/q-\epsilon}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1}{q}+\epsilon\right)$ (23)
$=-\log_{q}\left(\frac{1}{q}\left(1-\frac{\epsilon q}{q-1}\right)\right)+\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)$
$=1-\frac{1}{\ln q}\left[\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)\right]$
$=1+\frac{1}{\ln q}\left[-\ln\left(1-\frac{\epsilon q}{q-1}\right)+\left(\frac{1}{q}+\epsilon\right)\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(1+\epsilon q\right)\right].$
Since $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q}\right\}$, we have $0<\frac{\epsilon q}{q-1}<\frac{1}{2}$ and $10<\epsilon q<$ 1. Then it follows from .1 that $-\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}}$, $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}$ and $-\ln(1+\epsilon q)\leq-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}$. Now by Equation 23, we have
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $\leq 1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q}{q-1}\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\;\;\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}\right)\right]$
$=1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q^{2}}{q-1}+\frac{\epsilon^{2}q^{3}(q-2)}{2(q-1)^{2}}\right)\right]$
$=1+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\frac{\epsilon^{2}q^{2}}{q-1}+\frac{\epsilon^{2}q^{2}(q-2)}{2(q-1)^{2}}\right]$
$=1-\frac{\epsilon^{2}q^{2}}{2(q-1)\ln q}+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}$
$\overset{(a)}{\leq}1-\frac{\epsilon^{2}}{4},$
where (a) follows from the fact that $\epsilon<\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)$. ∎