### v5

# Codes Correcting Two Bursts of Exactly $b$ Deletions

Abstract
In this paper, we investigate codes designed to correct two bursts of deletions, where each burst has a length of exactly $b$, where $b>1$. The previous best construction, achieved through the syndrome compression technique, had a redundancy of at most $7\log n+O\left(\log n/\log\log n\right)$ bits. In contrast, our work introduces a novel approach for constructing $q$-ary codes that attain a redundancy of at most $5\log n+O(\log\log n)$ bits for all $b>1$ and $q\geq 2$. Additionally, for the case where $b=1$, we present a new construction of $q$-ary two-deletion correcting codes with a redundancy of $5\log n+O(\log\log n)$ bits, for all $q>2$.
Index Terms:
deletion, burst-deletion, error-correcting codes, DNA-based storage
I Introduction
Asubset $\mathcal{C}\subseteq\{0,1,\ldots,q-1\}^{n}$ (where $q\geq 2$) is called a $t$-deletion correcting code, if it has the property that if a codeword $\mathbf{x}\in\mathcal{C}$ is corrupted by deleting $t$ symbols to obtain a subsequence $\mathbf{y}\in\{0,1,\ldots,q-1\}^{n-t}$, then one can recover $\mathbf{x}$ from $\mathbf{y}$. The study of deletion correcting codes has a long history, dating back to at least the 1960s [1]. The seminal work in this field is [2], whereby proposing a linear-time decoding algorithm, Levenshtein proved that the binary code (Varshamov-Tenengolts code, or VT code for short) constructed in [3] can combat a single deletion error. In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4]. For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|\mathcal{C}\right|\right)$ [^1] All logarithms in this paper are to the base $2$., is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lower bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6]. This implies that the VT code in [3] and the Tenengolts code [4] have redundancy optimal up to a constant.
Due to applications in DNA-based data storage [7, 8, 9], document exchange [10, 11], multiple-deletion correcting codes with low redundancy have attracted a lot of interest in recent years [10, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19, 24]. For general $t\geq 3$, the smallest redundancy, which is $(4t-1)\log n+o(\log n)$, was achieved by a construction given in [20]. For non-binary alphabets, Sima et al [15] presented a family of $q$-ary $t$-deletion correcting codes with redundancy $4t\log n+o(\log n)$ by using the syndrome compression technique. The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding. A straightforward application of this method will give a $q$-ary $t$-deletion correcting code with redundancy $(4t-1)\log n+o(\log n)$, which is the best-known result in redundancy. When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary 22-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ [21]. And in a following work [22] the authors presented a $q$-ary 22-deletion correcting code with redundancy $5\log n+O(\log\log n)$ for all $q>2$. When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23]. In addition, they showed that there is a linear time encoder with near-optimal redundancy for their code.
If deletions occur at consecutive positions, we call them a burst of deletions. Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 25], wireless sensor networks, and satellite communication devices [26]. A code is called a $b$-burst-deletion correcting code, if it can correct any single burst of exactly $b$ deletions. In 1970, Levenshtein presented a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [27]. For $b\geq 3$, Cheng et al in 2014 constructed a class of binary codes with redundancy $b\log(n/b+1)$ [28], which was later improved to $\log n+(b-1)\log\log n+O(1)$ by Schoeny et al in 2017 [29]. Schoeny’s result was generalized to non-binary alphabets in [30, 31]. The best known redundancy for all $q\geq 2$ is $\log n+O(1)$, which was contributed recently by Sun et al[32]. It was proved in [29, 31] that the redundancy of a $b$-burst-deletion correcting code is at least $\log n+\Omega(1)$. Therefore, the codes in [27] and [32, Theorem 9] have redundancy optimal up to a constant. There are also a lot of works on codes correcting single burst of at most $b$ deletions [27, 29, 33, 34, 35, 21, 36, 37, 32]. For readers’ convenience, we summarize previous results on codes correcting bursts of deletions in Table I.
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. We call such codes 22-$b$-burst correcting codes. To the best of our knowledge, there are no explicit results about such codes. A related result can be found in a work of Sima et al [34, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively ($(t,t_{L})$ burst deletions, for short). Let $t=2$ and $t_{L}=b$. Then their result gives a binary 22-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes. On the other hand, [17, 19, 21, 22, 24] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique. Specifically, [17] and [19, 24] presented binary codes with redundancy at most $7\log n+O(1)$ and $4\log n+O(\log\log n)$, respectively. For non-binary codes, the best-known redundancy is $5\log n+O(\log\log n)$ [21, 22]. Note that in a burst of deletions, all deletions occur consecutively. Therefore, it is reasonable to deem by intuition that there is no big difference between two-deletion correcting codes and 22-$b$-burst-deletion correcting codes (where $b>1$). This raises a natural question: for $b>1$, is there a construction that leads to codes that are as good as, or even better than, the one given by the syndrome compression technique? Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$. Our contributions include:
  • We establish lower and upper bounds on the size (or equivalently, the redundancy) of 22-$b$-burst-deletion correcting codes;
  • A binary 22-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+14b\log\log n+O(1)$, for any $b>1$;
  • A $q$-ary 22-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O(1)$, for any $q>2$ and $b>1$;
  • A new construction of $q$-ary two-deletion correcting codes of length $n$ with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for any $q>2$.
Here, it is assumed that $q$ and $b$ are constants with respect to $n$. Therefore, our results show that for 22-$b$-burst-deletion correcting codes, we can do almost as well as two-deletion correcting codes.
TABLE I: Previous codes correcting bursts of deletions and corresponding methods
Redundancies of $q$-ary Codes References      Core Methods
single burst
of size
exactly $b$
$b\log(n/b+1)$
($q=2$)
[28]
1. representing each codeword as an array with $b$ rows
2. imposing a VT constraint on each row
$\log n+(b-1)\log\log n+O(1)$
($q=2$, $q>2$)
[29]
[30]
1. representing each codeword as an array with $b$ rows
2. encoding the first row with a VT code with run-length limited constraint
3. encoding the rest rows with shifted VT codes
$\log n+O(1)$ ($q\geq 2$) [32]
1. representing each codeword as an array with $b$ rows
2. representing this array as a $q^{b}$-ary vector
3. imposing two types of sum constraints on this vector
4. consider the signature of this vector
5. imposing a VT-type constraint together with three types of sum constraints
   on the signature
single burst
of size
at most $b$
$\log n+1$ ($b=2,q=2$)
[27]
[36]
1. imposing a sum constraint on rank sequences of codewords
       or
2. imposing a VT-type constraint on differential sequences of codewords
$(b-1)\log n+\left(\binom{b}{2}-1\right)\log\log n$
            +O(1)+O(1)
($q=2$)
[29]
   extension of the construction of $b$-burst-deletion correcting codes
   in the same work
$\left\lceil\log b\right\rceil\left(\log n+\binom{b+1}{2}\log\log n\right)$
           +O(1)+O(1)
($q=2$)
[33]
1. denoting each $t\in[1,b]$ as $t=2^{i}\cdot j$, where $i\geq 0$ and $j$ is odd
2. representing each code word as an array with $2^{i}$ rows
3. imposing a VT-type constraint alongside a “balanced” constraint on the
   first row to approximate the error positions
4. representing each codeword as an array with $t$ rows and encoding
   each row with a shifted VT code
$\log n+\binom{b+1}{2}\log\log n+O(1)$
($q=2$)
[35]
1. each codeword $\mathbf{x}$ is required to be $(\bm{p},\delta)$-dense
2. associating with $\mathbf{x}$ a vector of integers $a_{\bm{p}}(\mathbf{x})$
3. imposing a constraint on the number of $\bm{p}$ in $\mathbf{x}$ and a VT-type constraint
   on $a_{\bm{p}}(\mathbf{x})$ to approximate positions of errors
5. for each $1\leq t\leq b$, representing $\mathbf{x}$ as an array with $t$ rows and
   encoding each row with shifted VT codes
$4\log n+o\left(\log n\right)$ [34] syndrome compression technique
$\log n+8\log\log n+o\left(\log\log n\right)$
($q=2$)
[21]
1. applying the same method in [35] to approximate positions of errors
2. applying a code with $4\log n+o(\log n)$ redundant bits to correct burst
   deletions in short intervals
$\log n+(8\log q+8)\log\log n$
      $o\left(\log q\log\log n\right)$
($q>2$ is even)
1. representing each codeword as a binary array with $\left\lceil\log q\right\rceil$ rows
2. encoding the first row with the binary code in the same work
3. applying a code with $4\log n+o(\log n)$ redundant bits to correct
   errors in remaining rows
$\log n+8\log\log n+o\left(\log\log n\right)$
($q\geq 2$)
[37]    generalization of binary codes in [21]
$\log n+\log q\log\log n+O(1)$
($b=2$, $q>2$ is even)
[36]
1. representing each codeword as a binary array with $\left\lceil\log q\right\rceil$ rows
2. encoding the first row with the binary code in the same work
   with additional pattern-limited constraint
3. applying a $P$-bounded version of the binary code to remaining rows
$\log n+b\log\log n+O(1)$
($q\geq 2$)
[32]
1. associating each codeword with a binary sequence
2. applying the same method in [35] to the binary sequence to approximate
   error positions
3. for each $1\leq b^{\prime}\leq b$, applying a bounded $b^{\prime}$-burst-deletion correcting code
   to correct errors in short intervals
$(t,t_{L})$
burst
deletions
$4t\log n+o(\log\log n)$
($q=2$)
[34]
1. $t$-mixed sequences
2. syndrome compression technique
The rest of this paper is organized as follows. In Section II, we introduce some necessary definitions and related results. In Section III, we bound above and below the size (or equivalently, the redundancy) of codes correcting two bursts of exactly $b$ deletions. Section IV deals with codes for correcting two $b$-burst-deletions. In Section V, we give a new construction of non-binary two-deletion correcting codes. Finally Section VI concludes this paper.
II Preliminaries
In this section, we introduce some necessary definitions, auxiliary conclusions and related results.
For an integer $q\geq 2$ and a positive integer $n$, denote $\Sigma_{q}=\left\{0,1,\ldots,q-1\right\}$ and $\Sigma_{q}^{n}$ the set of all $q$-ary sequences with $n$ symbols. Let $\mathbf{x}\in\Sigma_{q}^{n}$ be a sequence. Unless otherwise stated, the $i$th coordinate of $\mathbf{x}$ is denoted by $x_{i}$, i.e., $\mathbf{x}=x_{1}\cdots x_{n}$. We call $n$ the length of $\mathbf{x}$ and denote $\left|\mathbf{x}\right|=n$. For a finite set $A$, we denote by $\left|A\right|$ the cardinality of $A$.
For two integers $m$ and $n$ such that $m\leq n$, let $[m,n]$ denote the set $\left\{m,m+1,\ldots,n\right\}$. If $m=1$, denote $[n]=[1,n]$. For a sequence $\mathbf{x}\in\Sigma_{q}^{n}$ and a subset $I=\left\{i_{1},i_{2},\ldots,i_{t}\right\}\subseteq[n]$ where $i_{1}<i_{2}<\cdots<i_{t}$, we define $\mathbf{x}_{I}\triangleq x_{i_{1}}x_{i_{2}}\cdots x_{i_{t}}$. For each $I\subseteq[n]$, we say $\mathbf{x}_{I}$ is a subsequence of $\mathbf{x}$. In particular, if $I$ is an interval of $[n]$ (i.e., $I=[i,j]$ for some $1\leq i\leq j\leq n$), we say $\mathbf{x}_{I}$ is a substring of $\mathbf{x}$. A run in $\mathbf{x}$ is a maximal substring consisting of the same symbols. The number of runs of $\mathbf{x}$ is denoted by $r(\mathbf{x})$. For example, if $\mathbf{x}=100101$, then there are five runs in $\mathbf{x}$: $1$, $00$, 11, 0 and $1$. So $r\left(\mathbf{x}\right)=5$.
The concatenation of two sequences $\mathbf{x}$ and $\mathbf{y}$ is denoted by $\mathbf{x}\mathbf{y}$. For example, let $\mathbf{x}=102$ and $\mathbf{y}=121$ be two sequences in $\Sigma_{3}^{3}$, then $\mathbf{x}\mathbf{y}=102121\in\Sigma_{3}^{6}$. Let $b$ and $n$ be two positive integers satisfying $b<n$. When a substring of length $b$ is deleted, we refer to it as a deletion-burst of size $b$ or a $b$-burst-deletion; that is to say, from $\mathbf{x}\in\Sigma_{q}^{n}$, we obtain a subsequence $\mathbf{x}_{[n]\setminus[i,i+b-1]}$ for some $1\leq i\leq n-b+1$.
In this paper, we focus on codes correcting two $b$-burst-deletions. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$, where $n>2b$. There are two ways to define two $b$-burst-deletions:
  1. (D1)
    the two bursts are caused by two channels: $\mathbf{x}$ passes the first channel, resulting in $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$ and then $\mathbf{z}$ passes the second channel, resulting in $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$;
  2. (D2)
    the two bursts are caused by a single channel: symbols in $\mathbf{x}$ pass a channel sequentially and we receive $\mathbf{y}=\mathbf{x}_{[n]\setminus I_{1}\cup I_{2}}$, where $I_{1}$ and $I_{2}$ are two disjoint intervals of length $b$ in [n].
Remark II.1
There is another possibility: the two bursts might overlap and result in a single burst that is shorter than $2b$. We do not take this situation into account, since it is covered by a more comprehensive problem: correcting two bursts of deletions, where each burst has length at most $b$. Our idea in this paper fails in this situation. We left this problem for future research.
In fact, (D1) and (D2) are equivalent. Firstly, it is clear that (D1) covers (D2). Next, we show that (D2) also covers (D1).
Observation II.1
Let $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}$ is obtained from $\mathbf{x}$ by process (D1). Then there exist two intervals $I_{1}=[j_{1},j_{1}+b-1]$, $I_{2}=[j_{2},j_{2}+b-1]\subseteq[n]$, where $j_{2}-j_{1}\geq b$, such that $\mathbf{y}=\mathbf{x}_{[n]\setminus(I_{1}\cup I_{2})}$. In particular, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by two $b$-burst-deletions, we can always assume that $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two non-overlapping substrings of length $b$ from $\mathbf{x}$.
Proof:
By assumption, there is $1\leq i_{1}\leq n-b+1$ and $1\leq i_{2}\leq n-2b+1$ such that $\mathbf{y}=\mathbf{z}_{[n-b]\setminus[i_{2},i_{2}+b-1]}$, where $\mathbf{z}=\mathbf{x}_{[n]\setminus[i_{1},i_{1}+b-1]}$. If $i_{2}\geq i_{1}$, let $j_{1}=i_{1}$ and $j_{2}=i_{2}+b$. Then the conclusion follows.
Now suppose $1\leq i_{2}\leq i_{1}-1$. If $i_{2}\leq i_{1}-b$, let $j_{1}=i_{2}$ and $j_{2}=i_{1}$. Then the conclusion follows. If $i_{1}-b<i_{2}<i_{1}$, it is clear that $\mathbf{y}=\mathbf{x}_{[n]\setminus[i_{2},i_{2}+2b-1]}$. Let $j_{1}=i_{2}$ and $j_{2}=i_{2}+b$. Then the conclusion follows. ∎
For $t\in\{1,2\}$ and $n>tb$, define
$\mathcal{B}_{t}^{b}\left(\mathbf{x}\right)=\left\{\mathbf{y}\in\Sigma_{q}^{n-tb}:\begin{array}[]{c}\mathbf{y}\text{ is obtained from }\mathbf{x}\\ \text{by }t\;b\text{-burst-deletion(s)}\end{array}\right\}.$
When $b=1$, we use notation $\mathcal{B}_{t}(\mathbf{x})$ instead of $\mathcal{B}_{t}^{b}\left(\mathbf{x}\right)$.
Definition II.1
Let $\mathcal{C}$ be a subset of $\Sigma_{q}^{n}$ with $\left|\mathcal{C}\right|\geq 2$. Suppose $t\in\{1,2\}$. We call $\mathcal{C}$ a $t$-$b$-burst-deletion correcting code if $\mathcal{B}_{t}^{b}(\mathbf{x})\cap\mathcal{B}_{t}^{b}(\mathbf{y})=\emptyset$ for any two distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}$. In particular, if $b=1$, we call $\mathcal{C}$ a $t$-deletion correcting code.
Clearly, if any $\mathbf{x}\in\mathcal{C}$ can be uniquely and efficiently recovered from any given $\mathbf{x}^{\prime}\in\mathcal{B}_{t}^{b}(\mathbf{x})$, then $\mathcal{C}$ is a $t$-$b$-burst-deletion correcting code. Here, “efficiently” means that the time complexity of decoding $\mathbf{x}$ from $\mathbf{y}$ is polynomial in $n$. In this paper, we construct 22-$b$-burst-deletion correcting codes and show that any codeword can be uniquely and efficiently decoded.
The redundancy of a code $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is defined to be $\rho(\mathcal{C})=\log\left(q^{n}/\left|\mathcal{C}\right|\right)$. All logarithms in this paper are to the base $2$. In addition, we always assume that $q$ and $b$ are fixed with respect to the code-length $n$.
[…]
$A\left(\mathbf{x},n^{\prime}\right)_{i}\triangleq\left(\tilde{x}_{i},\tilde{x}_{i+n^{\prime}},\tilde{x}_{i+2n^{\prime}},\ldots,\tilde{x}_{i+n^{\prime}\left(\left\lceil\frac{n}{n^{\prime}}\right\rceil-1\right)}\right).$
We call $A\left(\mathbf{x},n^{\prime}\right)$ a matrix (or array) representation of $\mathbf{x}$. If $n^{\prime}$ is clear from the context, we will denote $A\left(\mathbf{x},n^{\prime}\right)$ by $A(\mathbf{x})$.
For example, let $n=7$, $\mathbf{x}=1011010\in\Sigma_{2}^{7}$ and $n^{\prime}=2$. Then $\widetilde{\mathbf{x}}=10110100\in\Sigma_{2}^{8}$ and
$A\left(\mathbf{x},2\right)=\begin{pmatrix}1&1&0&0\\ 0&1&1&0\end{pmatrix}$.
If $n^{\prime}=3$, then $\widetilde{\mathbf{x}}=\mathbf{x}00$ and so
$A\left(\mathbf{x},3\right)=\begin{pmatrix}1&1&0\\ 0&0&0\\ 1&1&0\end{pmatrix}$.
In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as matrix $A(\mathbf{x},b)$. To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$. All results in this paper still hold even if $b\nmid n$, as long as we replace $n/b$ by $\left\lceil n/b\right\rceil$. Throughout this paper, for a matrix $A$, denote by $A_{i}$ and $A_{i,j}$ the $i$-th row of $A$ and the entry in the $i$-th row and $j$-th column, respectively.
The next observation is a straightforward result of II.1.
Observation II.2
For each $b<n$ and $b\mid n$, we can represent a sequence $\mathbf{x}\in\Sigma_{q}^{n}$ as a $b\times n/b$ array $A(\mathbf{x})$. Any two $b$-burst-deletions in $\mathbf{x}$ will induce two deletions in each row of $A(\mathbf{x})$. Furthermore, if the positions of the two deletions in $A(\mathbf{x})_{1}$ are $j_{1}$ and $j_{2}$, then for each $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred at coordinate $j_{1}-1$ or $j_{1}$, and the other deletion occurred at coordinate $j_{2}-1$ or $j_{2}$.
II-A Related Results
Recently, there have been two notable developments in non-binary two-deletion correcting codes with low redundancy. In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for any even $q>2$. In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for any $q>2$. In Section V, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for any $q>2$.
Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available. A related result can be found in the work of Sima et al [34, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively. For $t=2$ and $t_{L}=b$, their result provides a binary 22-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. Their result was derived using the syndrome compression technique, which was later extended to syndrome compression with pre-coding in [20]. We will apply this extended technique to provide a $q$-ary code with redundancy at most $7\log n+O\left(\log n/\log\log n\right)$, for all $q\geq 2$.
For a subset $\mathcal{E}\subseteq\Sigma_{q}^{n}$ and a sequence $\mathbf{x}\in\mathcal{E}$, define
$\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)=\left\{\mathbf{y}\in\mathcal{E}~:~\mathbf{y}\neq\mathbf{x}\text{ and }\mathcal{B}_{2}^{b}\left(\mathbf{y}\right)\cap\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)\neq\emptyset\right\}.$
In other words, $\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$ is the set of all sequences (except $\mathbf{x}$) in $\mathcal{E}$ whose error-ball intersects with that of $\mathbf{x}$.
Lemma II.1
[13, 20, 38] Let $\mathcal{E}\subseteq\Sigma_{q}^{n}$ be a code and $N>\max\left\{\left|\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)\right|~:~\mathbf{x}\in\mathcal{E}\right\}$. Suppose that the function $f:\Sigma_{q}^{n}\rightarrow\{0,1\}^{R(n)}$ (where $R(n)$ is a function of $n$ and $R(n)\geq 2$) satisfies the following property:
  1. $(\textup{{P}}1)$
    if $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{y}\in\mathcal{N}_{\Sigma_{q}^{n}}\left(\mathbf{x}\right)$, then $f\left(\mathbf{x}\right)\neq f\left(\mathbf{y}\right)$.
Then there exists a function $\bar{f}:\mathcal{E}\rightarrow\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, computable in polynomial time [^2] In this paper, when saying that a function is computable in polynomial/linear time, we mean that this function is computable in time polynomial/linear in the code-length $n$. such that $\bar{f}\left(\mathbf{x}\right)\neq\bar{f}\left(\mathbf{y}\right)$ for any $\mathbf{x}\in\mathcal{E}$ and $\mathbf{y}\in\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$.
Let $\mathcal{E}$ be a 11-$b$-burst correcting code and $\bar{f}$ be given in Lemma II.1. Then Lemma II.1 asserts that if $\mathbf{x},\mathbf{y}\in\mathcal{E}$ are distinct codewords and $\bar{f}(\mathbf{x})=\bar{f}(\mathbf{y})$, we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$. Therefore, for any $\mathbf{a}\in\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, the code $\mathcal{E}^{\prime}=\left\{\mathbf{x}\in\mathcal{E}:\bar{f}(\mathbf{x})=a\right\}$ is a 22-$b$-burst-deletion correcting code. Furthermore, by the pigeonhole principle, there exists an $\mathbf{a}$ such that the redundancy of $\mathcal{E}^{\prime}$ is at most $\rho\left(\mathcal{E}^{\prime}\right)=\rho(\mathcal{E})+2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$. [^3] We select $\mathcal{E}$ to be a 11-$b$-burst correcting code to obtain a better redundancy. If we take $\mathcal{E}=\Sigma_{q}^{n}$ we get a redundancy of $8\log n+o(\log n)$.
For the choice of $\mathcal{E}$, we have the following result.
Lemma II.2
[32, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n\geq b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
This lemma gives a 11-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$. Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|<q^{b}n^{3}$. In fact, each codeword in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$ can be obtained in the following three steps:
  1. 1)
    Delete two substrings of length $b$ from $\mathbf{x}$, resulting in $\mathbf{z}^{(1)}$. There are less than $n^{2}$ possibilities for $\mathbf{z}^{(1)}$.
  2. 2)
    For each $\mathbf{z}^{(1)}$, insert a sequence of length $b$ into $\mathbf{z}^{(1)}$ and get a sequence $\mathbf{z}^{(2)}\in\Sigma_{q}^{n-b}$. For each $\mathbf{z}^{(1)}$, there are at most $q^{b}n$ possibilities for $\mathbf{z}^{(2)}$.
  3. 3)
    Insert a sequence of length $b$ into $\mathbf{z}^{(2)}$ to get a sequence $\mathbf{y}\in\Sigma_{q}^{n}$. Since $\mathcal{E}$ is a 11-$b$-burst-deletion correcting code, for each $\mathbf{z}^{(2)}$, there is at most one $\mathbf{y}$ which is in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$.
Therefore, we can let $N=q^{b}n^{3}$. Now the redundancy of $\mathcal{E}^{\prime}$ is at most $7\log n+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$. To conclude our discussion, it remains to find an $f$ satisfying (P1) in Lemma II.1 such that $R(n)=O(\log n)$, which is given in Lemma II.5. The proof of Lemma II.5 is based on Lemmas II.3 and II.4.
The following lemma is a corollary of [17, Theorem 2].
Lemma II.3
[17, Theorem 2] For any integer $n\geq 3$, there exists a function $\xi:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{7\log n+O\left(1\right)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\xi\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ (i.e., $\mathbf{y}$ is obtained from $\mathbf{x}$ by two deletions), one can uniquely and efficiently recover $\mathbf{x}$.
This result can be extended to arbitrary finite alphabets in the following way.
Lemma II.4
Suppose $q\geq 2$. There is a function $\xi_{1}:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{7\left\lceil\log q\right\rceil\log n+O_{q}(1)}$, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ and $\xi_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q}(1)$ is a constant dependent only on $q$.
Proof:
Any $\mathbf{x}\in\Sigma_{q}^{n}$ can be uniquely represented as an array
$M(\mathbf{x})\triangleq\begin{pmatrix}x_{1,1}&\cdots&x_{1,n}\\ \vdots&\cdots&\vdots\\ x_{\left\lceil\log q\right\rceil,1}&\cdots&x_{\left\lceil\log q\right\rceil,n}\end{pmatrix}$,
where $x_{k,i}\in\{0,1\}$ such that $x_{i}=\sum_{k=1}^{\left\lceil\log q\right\rceil}x_{k,i}2^{k-1}$ for all $1\leq i\leq n$.
Denote the $k$-th row of $M(\mathbf{x})$ by $M(\mathbf{x})_{k}$. Suppose $\mathbf{y}\in\mathcal{B}_{2}(\mathbf{x})$. It is clear that $M(\mathbf{y})_{k}\in\mathcal{B}_{2}\left(M(\mathbf{x})_{k}\right)$ for all $1\leq k\leq\left\lceil\log q\right\rceil$. Let $\xi(\cdot)$ be the function defined in Lemma II.3. For $\mathbf{x}\in\Sigma_{q}^{n}$, define $\xi_{1}(\mathbf{x})\triangleq\left(\xi\left(M(\mathbf{x})_{1}\right),\ldots,\xi\left(M(\mathbf{x})_{\left\lceil\log q\right\rceil}\right)\right)$. Then by Lemma II.3, given $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ and $\xi_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi\left(M(\mathbf{x})_{k}\right)$ is a binary vector of length $7\log n+O(1)$, we can see that $\xi_{1}(\mathbf{x})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log n+O_{q}(1)$. ∎
Lemma II.5
Suppose $b>1$, $n>2b$ and $q\geq 2$. There is a function $\psi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)}$, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$.
Proof:
Let $A(\mathbf{x})=A(\mathbf{x},b)$ and $A(\mathbf{y})=A(\mathbf{y},b)$. Since $\mathbf{y}\in\mathcal{B}_{2}^{b}(\mathbf{x})$, we have $A(\mathbf{y})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. Define $\psi(\mathbf{x})\triangleq\left(\xi_{1}\left(A(\mathbf{x})_{1}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b}\right)\right)$. Then by Lemma II.4, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi_{1}(A(\mathbf{x})_{i})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log(n/b)+O_{q}(1)$, we can see that $\psi(\mathbf{x})$ is a binary vector of length $7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)$. ∎
Taking $f=\psi$ gives a function satisfying (P1) as needed.
Before proceeding to subsequent sections, we introduce a useful lemma, which will be used in Section IV-B.
Let $N\geq 2$ be an integer. Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$. Let ?? denote an unknown symbol (not in [0,N1][0,N-1]). If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for any $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two). For our purpose in Section IV-B, assume $j\geq i+2$. For a sequence $\mathbf{x}$ over the alphabet [0,N1][0,N-1], denote $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$. In the following lemma, denote $A(\mathbf{x})=A(\mathbf{x},2)$, i.e., the matrix representation of $\mathbf{x}$ with two rows.
Lemma II.6
For any $0\leq a_{1},a_{2}<2N$ and $0\leq b<nN^{2}$, define $\mathcal{C}$ to be the set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
  • (C1)
    $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\equiv a_{1}\pmod{2N}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\equiv a_{2}\pmod{2N}$;
  • (C2)
    $W(\mathbf{x})\equiv b\pmod{nN^{2}}$, where $W(\mathbf{x})=\textup{Syn}\left(A(\mathbf{x})_{1}\right)+(2N-1)\cdot\textup{Syn}\left(A(\mathbf{x})_{2}\right)$.
Then the code $\mathcal{C}$ can correct two bursts (of length two) of erasures. In particular, there is a function $\varphi:[0,N-1]^{n}\rightarrow\Sigma_{2}^{\log n+4\log N+2}$, efficiently computable, such that for any $\mathbf{x}\in[0,N-1]^{n}$, given $\varphi(\mathbf{x})$, we can efficiently and uniquely recover $\mathbf{x}$ from $\mathbf{y}$, where $\mathbf{y}$ is any given sequence obtained from $\mathbf{x}$ by two bursts of erasures.
Proof:
Firstly, if the correctness of the code $\mathcal{C}$ is proved, we can define
[…]
$\quad\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\pmod{2N}$,
$\quad W(\mathbf{x})\pmod{nN^{2}}\Bigg{)}$.
Here, we view $\varphi(\mathbf{x})$ as a binary vector. There are at most $(2N)^{2}\cdot(nN^{2})$ values of $\varphi(\mathbf{x})$. As a result, the length of $\varphi(\mathbf{x})$, when viewed as a binary vector, is at most $\log n+4\log N+2$. It remains to prove the correctness of $\mathcal{C}$.
Suppose $\mathbf{y}$ is obtained from a codeword $\mathbf{x}\in\mathcal{C}$ by two bursts of erasures. Then $A(\mathbf{y})_{i}$ is obtained from $A(\mathbf{x})_{i}$ by two erasures for each $i=1,2$. Denote the error positions in $A(\mathbf{y})_{1}$ are $i_{1},i_{2}$, and the error positions in $A(\mathbf{y})_{2}$ are $i_{3},i_{4}$, where $i_{1}<i_{2}$ and $i_{3}\leq i_{4}$. Note that $i_{1},i_{2},i_{3}$ and $i_{4}$ are known to us. Clearly, we have $i_{3}\in\{i_{1},i_{1}-1\}$ and $i_{4}\in\{i_{2},i_{2}-1\}$. This implies that $i_{4}-i_{3}\in\{i_{2}-i_{1},i_{2}-i_{1}-1,i_{2}-i_{1}+1\}$.
Next, we describe how to decode $\mathbf{x}$ from $\mathbf{y}$. Clearly, it is sufficient to recover the values of $A(\mathbf{x})_{1,i_{1}}$, $A(\mathbf{x})_{1,i_{2}}$, $A(\mathbf{x})_{2,i_{3}}$ and $A(\mathbf{x})_{2,i_{4}}$. Let
$\delta_{1}=\left(a_{1}-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\right)\pmod{2N}$, (1)
$\delta_{2}=\left(a_{2}-\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\right)\pmod{2N}$.
Since $0\leq A(\mathbf{x})_{1,i_{1}}+A(\mathbf{x})_{1,i_{2}},A(\mathbf{x})_{2,i_{3}}+A(\mathbf{x})_{2,i_{4}}<2N$ and $0\leq\delta_{1},\delta_{2}<2N$, it follows from Condition (C1) and Equation 1 that $\delta_{1}=A(\mathbf{x})_{1,i_{1}}+A(\mathbf{x})_{1,i_{2}}$ and $\delta_{2}=A(\mathbf{x})_{2,i_{3}}+A(\mathbf{x})_{2,i_{4}}$.
For simpler notations, denote $\alpha_{1}=A(\mathbf{x})_{1,i_{1}}$ and $\alpha_{2}=A(\mathbf{x})_{2,i_{3}}$. Then we have $A(\mathbf{x})_{1,i_{2}}=\delta_{1}-\alpha_{1}$ and $A(\mathbf{x})_{2,i_{4}}=\delta_{2}-\alpha_{2}$. Therefore, it remains to obtain the values of $\alpha_{1}$ and $\alpha_{2}$. To that end, let
$\Delta=$ $\left(b-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{1,j}\right$. (2)
$\quad\quad\left.-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{2,j}\right)\pmod{nN^{2}}$.
$\Delta$ $=\left(W(\mathbf{x})-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{x})_{2,j}\right)\pmod{nN^{2}}$ (3)
$=\left(i_{1}\alpha_{1}+i_{2}(\delta_{1}-\alpha_{1})+(2N-1)(i_{3}\alpha_{2}+i_{4}(\delta_{2}-\alpha_{2}))\right)\pmod{nN^{2}}$.
Since each term in the right-hand side of (2) is known, we can obtain the value of $\Delta$. Furthermore, by (2), Condition (C2), and the relationship between $A(\mathbf{x})$ and $A(\mathbf{y})$, we have (3).
Let $\Delta^{\prime}=\left(i_{2}\delta_{1}+(2N-1)i_{4}\delta_{2}-\Delta\right)\pmod{nN^{2}}$, which can be calculated since values of $\delta_{1}$, $\delta_{2}$ and $\Delta$ are known to us. Then it follows from Equation 3 that
$(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}\equiv\Delta^{\prime}\pmod{nN^{2}}$. (4)
Since $i_{1}<i_{2}\leq\left\lceil n/2\right\rceil$ and $i_{4}-i_{3}\leq\left\lceil n/2\right\rceil$, we have $1\leq i_{2}-i_{1},i_{4}-i_{3}\leq n/2$. Combining this with the fact $0\leq\alpha_{1},\alpha_{2}\leq N-1$, we conclude that $0\leq(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}<nN^{2}$. Now Equation 4 implies
$(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}=\Delta^{\prime}.$ (5)
Recall that $i_{4}-i_{3}\in\{i_{2}-i_{1}-1,i_{2}-i_{1},i_{2}-i_{1}+1\}$. By definition, the two bursts of erasures in $\mathbf{x}$ do not overlap. It follows that $i_{2}\geq i_{1}+2$ if $i_{4}-i_{3}=i_{2}-i_{1}-1$. Therefore, we always have $i_{4}-i_{3}\geq 1$. This implies $(2N-1)(i_{4}-i_{3})>1$. Then it follows from Equation 5 that
$(i_{2}-i_{1})\alpha_{1}\equiv\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}$. (6)
Note that $0\leq\alpha_{1}\leq N-1$. When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it is easy to see that $0\leq(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. When $i_{4}-i_{3}=i_{2}-i_{1}-1$, since $i_{2}-i_{1}\geq 2$, we have $(2N-1)(i_{4}-i_{3})-(i_{2}-i_{1})\alpha_{1}=(i_{2}-i_{1})(2N-1-\alpha_{1})-(2N-1)\geq 2N-1-2\alpha_{1}>0$ and hence $(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. Now it follows from Equation 6 that $(i_{2}-i_{1})\alpha_{1}=\Delta^{\prime\prime}$, where $\Delta^{\prime\prime}=\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}$. From this, we get $\alpha_{1}=\frac{\Delta^{\prime\prime}}{(i_{2}-i_{1})}$. Then by Equation 5, we have $\alpha_{2}=\frac{\Delta^{\prime}-(i_{2}-i_{1})\alpha_{1}}{(2N-1)(i_{4}-i_{3})}$. Now the proof is completed. ∎
III Bounds
We could not find any existing upper or lower bounds on the maximum size of a 22-$b$-burst-deletion code. In this section, we will derive these bounds.
Let $M_{q,n,b}$ be the maximum size of a 22-$b$-burst-deletion correcting code in $\Sigma_{q}^{n}$, where $n>2b$.
Theorem III.1
The maximum size of a 22-$b$-burst-deletion correcting code satisfies $M_{n,q,b}\geq\frac{q^{n-2b}}{\binom{n}{2}^{2}}$.
Proof:
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. By definition, a subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a 22-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [39, page 100, Theorem 1] that
$M_{q,n,b}\geq\sum_{\mathbf{x}\in\Sigma_{q}^{n}}\frac{1}{d(\mathbf{x})+1}.$ (7)
For $\mathbf{x}\in\Sigma_{q}^{n}$, by II.1, we conclude that each $\mathbf{y}$ (including $\mathbf{x}$ itself) with $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$ can be obtained as follows: 1) deleting two non-overlap substrings of length $b$ from $\mathbf{x}$, resulting in a sequence $\mathbf{z}\in\Sigma_{q}^{n-2b}$; 2) inserting two sequences of length $b$ into $\mathbf{z}$. Therefore, we have $d(\mathbf{x})+1\leq\binom{n}{2}^{2}q^{2b}$. ∎
In [40, Section IV-B], the authors proved an upper bound of 11-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
Theorem III.2
For $q\geq 2$, let $f(q)=\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. If $n\geq 30$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, the maximum size of a 22-$b$-burst-deletion correcting code satisfies
$M_{q,n,b}\leq\left(\frac{3b^{2}}{q^{2b-2}(q-1)^{2}}+\frac{(1.121)^{3b}}{n}\right)\frac{q^{n}}{n^{2}}.$
Proof:
Define m=n/b1m=n/b-1. Let $\mathcal{C}\subseteq\Sigma_{q}^{n}$ be a 22-$b$-burst-deletion correcting code. Set $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and $t=\left(1-\frac{1}{q}-\epsilon\right)m$. We partition $\mathcal{C}$ into two disjoint subsets: $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$, where $\mathcal{C}_{1}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\geq t+2\text{ for some }1\leq i\leq b\right\}$ and $\mathcal{C}_{2}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\leq t+1\text{ for all }1\leq i\leq b\right\}$ (recall that $r(\cdot)$ denotes the number of runs). To derive an upper bound of $\left|\mathcal{C}\right|$, it is sufficient to upper bound $\left|\mathcal{C}_{1}\right|$ and $\left|\mathcal{C}_{2}\right|$.
For any $\mathbf{x}\in\mathcal{C}$, define $\mathcal{A}_{2}^{b}(\mathbf{x})=\left\{A(\mathbf{x}^{\prime}):\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})\right\}$. Let $A(\ell,k)$ be the array obtained by deleting the $\ell$-th and the $k$-th columns of $A(\mathbf{x})$ for $1\leq\ell\neq k\leq n/b$. It is clear that $A(\ell,k)\in\mathcal{A}_{2}^{b}(\mathbf{x})$ and
$\cup_{1\leq\ell\neq k\leq\frac{n}{b}}\left\{A(\ell,k)_{i}\right\}=\mathcal{B}_{2}(A(\mathbf{x})_{i}),$
for all $1\leq i\leq b$. By [41, eq. (11)], we know that
$\binom{r-1}{2}\leq\left|\mathcal{B}_{2}(\mathbf{v})\right|\leq\binom{r+1}{2}$
for any sequence $\mathbf{v}\in\Sigma_{q}^{n}$ with exactly $r$ runs. Since $\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|$ and $\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|=\left|\mathcal{A}_{2}^{b}(\mathbf{x})\right|$ for all $\mathbf{x}$, it follows that
$\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|$ $\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|$ (8)
$\geq\binom{\max_{1\leq i\leq b}\left\{r\left(A(\mathbf{x})_{i}\right)\right\}-1}{2}\geq\binom{t+1}{2}$,
for all $\mathbf{x}\in\mathcal{C}_{1}$. Since $\mathcal{C}$ is a 22-$b$-burst-deletion correcting code, we know that $\mathcal{C}_{1}$ is also a 22-$b$-burst-deletion correcting code. So we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for all distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}_{1}$. Then it follows from Equation 8 that $\left|\mathcal{C}_{1}\right|\binom{t+1}{2}\leq\sum_{\mathbf{x}\in\mathcal{C}_{1}}\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\leq q^{n-2b}$ and hence
$\left|\mathcal{C}_{1}\right|\leq\frac{q^{n-2b}}{\binom{t+1}{2}}$ $\leq\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{t^{2}}$ (9)
$=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{\left(1-1/q-\epsilon\right)^{2}\left(n/b-1\right)^{2}}$
$=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2b^{2}}{(1-1/q-\epsilon)^{2}(1-b/n)^{2}}$
$=\frac{2b^{2}q^{n+2}}{n^{2}q^{2b}(q-1)^{2}}\cdot\frac{1}{\left(1-\frac{q}{q-1}\epsilon\right)^{2}\left(1-b/n\right)^{2}}$
$\leq\frac{3b^{2}q^{n+2}}{n^{2}q^{2b}(q-1)^{2}}$
as long as $\left(1-\frac{q}{q-1}\epsilon\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, which is possible when $n$ is sufficiently large.
Next, we proceed to upper bound $\left|\mathcal{C}_{2}\right|$. To that end, define $\mathcal{C}^{\prime}$ to be the following set
$\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(b)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in\Sigma_{q}^{n/b},\;\forall 1\leq i\leq b\\ r\left(\mathbf{x}^{(i)}\right)\leq t+1,\;\forall 1\leq i\leq b\end{array}\right\}.$
Since $r\left(A(\mathbf{x})_{i}\right)\leq t+1$ for any $\mathbf{x}\in\mathcal{C}_{2}$ and $1\leq i\leq b$, we conclude that $\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|$. So it suffices to estimate an upper bound of $\left|\mathcal{C}^{\prime}\right|$.
Since m=n/b1m=n/b-1, by definition of $\mathcal{C}^{\prime}$, we have
$\left|\mathcal{C}^{\prime}\right|$ $=\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})\leq t+1\right\}\right|^{b}$ (10)
$=\left|\cup_{j=1}^{t+1}\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|^{b}$
$=\left(\sum_{j=1}^{t+1}\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|\right)^{b}$
$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}$
$=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
where (a) follows from the well-known result (see the proof of [42, Theorem 3.1]):
$\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|=\binom{m}{j-1}q(q-1)^{j-1}.$
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [43, Proposition 3.3.3] that
$\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\leq q^{mH_{q}(t/m)}=q^{mH_{q}(1-1/q-\epsilon)},$ (11)
where $H_{q}(x)\triangleq x\log_{q}(q-1)-x\log_{q}x-(1-x)\log_{q}(1-x)$ is the $q$-ary entropy function. Here, $\log_{q}x=\frac{\log x}{\log q}$ for any positive real number $x$.
If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. Now by Lemma A.1, we have
$H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
Now it follows from Equations 10 and 11 that
$\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}$ $\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}$ (12)
$=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}$
[…]
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $\sqrt[n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 9 and 12. ∎
Then next corollary is a direct consequence of Theorems III.1 and III.2.
Corollary III.1
Suppose that $q$ and $b$ are fixed. Let $f(q)$ be defined in Theorem III.2. When $n\geq\max\left\{30,\frac{(1.21)^{3b}q^{2b-2}(q-1)^{2}}{b^{2}}\right\}$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
$2\log n+\log\left(\frac{q^{2b-2}(q-1)^{2}}{4b^{2}}\right)\leq\rho(\mathcal{C})\leq 4\log n+2b\log q,$
where $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a 22-$b$-burst-deletion correcting code with maximum size.
Proof:
Firstly, the upper bound follows straightforward from Theorem III.1 just by noticing that $\binom{n}{2}^{2}\leq n^{4}$. Since $n\geq\frac{(1.21)^{3b}q^{2b-2}(q-1)^{2}}{b^{2}}$, we have $\left(\frac{3b^{2}}{q^{2b-2}(q-1)^{2}}+\frac{(1.121)^{3b}}{n}\right)\frac{q^{n}}{n^{2}}\leq\frac{4b^{2}}{q^{2b-2}(q-1)^{2}}\frac{q^{n}}{n^{2}}$. Now the lower bound follows from Theorem III.2. ∎
IV Codes For Correcting Two $b$-Burst-Deletions
In this section, we construct $q$-ary 22-$b$-burst-deletion correcting codes with redundancy at most $5\log n+O(\log\log n)$, for any $q\geq 2$ and $b>1$. Our idea is to first locate positions of deletions in short intervals (which is accomplished in Section IV-A, Lemma IV.5) and then correct errors in these intervals (which is accomplished in Sections IV-B and IV-C).
IV-A Approximately determine positions of deletions
Definition IV.1 (regularity)
A sequence $\mathbf{x}\in\Sigma_{2}^{n}$ is said to be $d$-regular if each substring of $\mathbf{x}$ of length at least $d\log n$ contains both $00$ and $11$.
For the number of regular sequences, we have the following lemma.
Lemma IV.1
[19, Lemma 11] The number of $d$-regular sequences of length $n$ is at least $2^{n-1}$, as long as $d\geq 7$ and $n$ be such that $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{0.15d-1}\geq 12$. In particular, when $d=7$, it suffices to require $n\geq 9$.
The next lemma ensures that we can efficiently correct two deletions in $d$-regular sequences.
Lemma IV.2
[19, Theorem 7] Suppose $d\geq 7$. There exists a function $\eta:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{4\log n+10\log\log n+O_{d}(1)}$, computable in linear time, such that for any $d$-regular sequence $\mathbf{x}\in\Sigma_{2}^{n}$, given $\eta\left(\mathbf{x}\right)$ and any $\mathbf{x}^{\prime}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{d}(1)$ denotes a constant depending only on $d$.
Furthermore, once $\mathbf{x}$ is recovered, locations of the two deletions in $\mathbf{x}$ can be approximated, as ensured by the following lemma.
Lemma IV.3
[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma_{2}^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$. When given $\mathbf{x}$ and $\mathbf{x}^{\prime}$, we can
  1. $(1)$
    either find distinct runs $\mathbf{x}_{J_{1}}$ and $\mathbf{x}_{J_{2}}$ of $\mathbf{x}$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, such that $i_{1}\in J_{1}$ and $i_{2}\in J_{2}$.
  2. $(2)$
    or find an interval $J\subseteq[n]$, uniquely determined by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, of length at most $3d\log n$ such that $i_{1},i_{2}\in J$.
A proof of Lemma IV.3 is given in [21]. Here we show the intuition behind this lemma.
Example IV.1
  1. $(1)$
    Let $\mathbf{x}=000111$ and $\mathbf{x}^{\prime}=0011$. Then $\mathbf{x}^{\prime}$ is obtained from $\mathbf{x}$ by deleting one bit in the run $000$ and one bit in the run $111$. This corresponds to case (1) in Lemma IV.3.
  2. $(2)$
    $\mathbf{x}=10010101110$ and $\mathbf{x}^{\prime}=100101110$. Then $\mathbf{x}^{\prime}$ can be obtained from $\mathbf{x}$ by deleting $x_{2}$ and $x_{4}$, or by deleting $x_{3}$ and $x_{4}$, or by deleting $x_{4}$ and $x_{5}$, or by deleting $x_{7}$ and $x_{8}$, or by deleting $x_{7}$ and $x_{9}$. Therefore, we can only locate error positions in the substring $001010111$, which is the concatenation of a run $00$, an alternating substring $1010$ and a run $111$. If $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular, then each run has length at most $d\log n$ and each alternating substring has length at most $d\log n$. This is why in case (2) of Lemma IV.3 we can locate error positions in an interval of length at most $3d\log n$.
Now we briefly explain why Lemma IV.3 is helpful. Suppose $\mathbf{x}\in\Sigma_{2}^{n}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$. Let $A(\mathbf{x})=A(\mathbf{x},b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},b)$ be the matrix representations of $\mathbf{x}$ and $\mathbf{x}^{\prime}$, respectively. If $A(\mathbf{x})_{1}$ is $d$-regular (where $d\geq 7$) and $\eta(A(\mathbf{x})_{1})$ is given, Lemma IV.2 ensures that we can decode $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ and Lemma IV.3 ensures that error positions in $A(\mathbf{x})_{1}$ can be approximately determined in one or two short intervals. By II.2, this further reveals information of error positions in remaining rows of $A(\mathbf{x})$, or equivalently, error positions in $\mathbf{x}$. Detailed analysis will be given Lemma IV.5. Before that, it is convenient to generalize the notion of regularity to general alphabets. To that end, we associate with any $q$-ary sequence a binary sequence.
Let $q\geq 2$. For any $x\in\Sigma_{q}$, we can uniquely write it as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<\left\lceil q/2\right\rceil$. Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$. We call $u(\mathbf{x})$ the binary sequence associated with $\mathbf{x}$. Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$ and when $q=2$, we have $u(\mathbf{x})=\mathbf{x}$. It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
Example IV.2
Let $q=3$. We have that $\left\lceil q/2\right\rceil=2$. If $x=1$, then $u_{x}=0$ and $v_{x}=1$. If $x=2$, then $u_{1}=1$ and $v_{x}=0$.
Remark IV.1
Although any $x\in\Sigma_{q}$ can be decomposed as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, it might exist some $u\in\{0,1\}$ and $0\leq v<\left\lceil q/2\right\rceil$ such that $\left\lceil q/2\right\rceil u+v\notin\Sigma_{q}$. In fact, if $q$ is even, then for any $u\in\{0,1\}$ and any $0\leq v<\left\lceil q/2\right\rceil$, we have $\left\lceil q/2\right\rceil u+v\in\Sigma_{q}$. But when $q$ is odd, we have $\left\lceil q/2\right\rceil+\left\lceil q/2\right\rceil-1=q\notin\Sigma_{q}$.
Definition IV.2
A sequence $\mathbf{x}\in\Sigma_{q}^{n}$ is said to be $d$-regular if $u(\mathbf{x})$ is $d$-regular. In other words, $\mathbf{x}$ is $d$-regular if and only if any substring of $\mathbf{x}$ of length at least $d\log n$ contains two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$.
Let $\mathcal{R}_{q,n,d}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\mathbf{x}\text{ is }d\text{-regular}\right\}$. To estimate a lower bound on our code size in Theorem IV.1, we need a lower bound on $\left|\mathcal{R}_{q,n,d}\right|$.
Lemma IV.4
Let $q\geq 2$.
  • If $q$ is even, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n}/2\geq q^{n-1}$ for all $d\geq 7$ and $n$ such that $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{0.15d-1}\geq 12$. In particular, when $d=7$, it suffices to require $n\geq 9$.
  • If $q$ is odd, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n-1}$ for all $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. Note that when $d\geq 10$, we have $1-\frac{d}{2}\log(0.87)>0$ and hence this condition could be satisfied when $n$ is sufficiently large. In particular, when $d=10$, it suffices to require $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
Proof:
Suppose first that $q$ is even. Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<q/2$. On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $0\leq v_{x}<q/2$. Then it follows from Lemma IV.1 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
Now suppose that $q\geq 3$ is odd. In this case, the previous argument does not hold. This is because when $q$ is odd, $u_{x}=1$ and $v_{x}=(q-1)/2$, we have $\left\lceil q/2\right\rceil u_{x}+(q-1)/2=q\notin\Sigma_{q}$. Fortunately, we can apply similar ideas in proofs of [19, Lemma 11] and [22, Lemma 5] to derive our lower bound on $\left|\mathcal{R}_{q,n,d}\right|$. For $m\geq 2$, define
$S_{m}^{L}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\begin{array}[]{c}\nexists 1\leq i<m\text{ such that}\\ 0\leq v_{i},v_{i+1}\leq(q-1)/2\end{array}\right\},$
$S_{m}^{U}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\begin{array}[]{c}\nexists 1\leq i<m\text{ such that}\\ (q+1)/2\leq v_{i},v_{i+1}<q\end{array}\right\}.$
We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$. By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}\right)^{3}=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$. Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$. For $\mathbf{v}\in S_{m}^{L}$, it must hold that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$. So it follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
For $\mathbf{v}\in\Sigma_{q}^{m}$, let $\overline{\mathbf{v}}\triangleq(q-v_{1})\cdots(q-v_{m})$. It is easy to verify that if $\mathbf{v}\in S_{m}^{U}$, then $\overline{\mathbf{v}}\in S_{m}^{L}$. Then it follows that $\left|S_{m}^{U}\right|\leq\left|S_{m}^{L}\right|<(0.87q)^{m}$.
Let $m=\left\lfloor\frac{d}{2}\log n\right\rfloor$ and
$Q=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\begin{array}[]{c}\exists i,j\text{ such that }v_{i},v_{i+1}<(q+1)/2\\ \text{and }v_{j}.v_{j+1}\geq(q+1)/2\end{array}\right\}.$
In other words, $Q$ is the set of length-$m$ sequences which contain two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$. Then we have $\left|Q\right|\geq q^{m}-\left|S_{m}^{L}\right|-\left|S_{m}^{U}\right|>q^{m}-2(0.87q)^{m}$. Now let $k=\left\lfloor n/m\right\rfloor$ and define $\mathcal{R}^{\prime}$ to be the following set
$\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(k)}\mathbf{x}^{(k+1)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 1\leq i\leq k\\ \mathbf{x}^{(k+1)}\in\Sigma_{q}^{n-km}\end{array}\right\}.$
It is easy to see that for any $\mathbf{x}\in\mathcal{R}^{\prime}$, any substring of $\mathbf{x}$ of length at least $d\log n$ must contain some $\mathbf{x}^{(i)}$, where $1\leq i\leq k$. Therefore, $\mathbf{x}$ is $d$-regular and hence $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$. This implies that
$\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}^{\prime}\right|$ $=\left|Q\right|^{k}q^{n-km}$
$>\left(q^{m}-2(0.87q)^{m}\right)^{k}q^{n-km}$
$=q^{n}\left(1-2(0.87)^{m}\right)^{k}$
$\overset{(a)}{\geq}q^{n}\left(1-2k(0.87)^{m}\right)$
$=q^{n}\left(1-2\left\lfloor\frac{n}{m}\right\rfloor 2^{\left\lfloor\frac{d}{2}\log n\right\rfloor\log(0.87)}\right)$
$\geq q^{n}\left(1-2\frac{n}{m}2^{\left\lfloor\frac{d}{2}\log n\right\rfloor\log(0.87)}\right)$
$\overset{(b)}{\geq}q^{n}\left(1-2\frac{n}{m}2^{\frac{d}{2}\log n\cdot\log(0.87)-\log(0.87)}\right)$
$=q^{n}\left(1-\frac{2}{0.87\left\lfloor\frac{d}{2}\log n\right\rfloor}n^{1+\frac{d}{2}\log(0.87)}\right)$
$\overset{(c)}{\geq}q^{n-1}.$
Here, (a) follows from the fact that $(1+x)^{r}\geq 1+rx$ for any integer $r\geq 1$ and any real number $x\geq-1$; (b) follows from the fact that $\log(0.87)<0$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor\geq\frac{d}{2}\log n-1$; (c) follow from the fact that $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. ∎
In Appendix B, we will discuss how to encode a sequence into a $d$-regular sequence.
Now we return to the aim of this subsection: approximately determine error positions. Let $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}$ be obtained from $\mathbf{x}$ by a 22-$b$-burst-deletion. Let $u(\mathbf{x})$ and $u(\mathbf{x}^{\prime})$ be the binary sequences associated with $\mathbf{x}$ and $\mathbf{x}^{\prime}$, respectively (see the paragraph prior to Example IV.2). Then we have $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}^{b}\left(u(\mathbf{x})\right)$. By the relationship between $\mathbf{x}$ and $u(\mathbf{x})$, it suffices to locate the 22-$b$-burst-deletion in $u(\mathbf{x})$. In the rest of this section, let $U(\mathbf{x})=A(u(\mathbf{x}),b)$ be the matrix representation of $u(\mathbf{x})$. Denote the $i$-th row of $U(\mathbf{x})$ by $U(\mathbf{x})_{i}$. Similarly, we can define $U(\mathbf{x}^{\prime})$ and $U(\mathbf{x}^{\prime})_{i}$.
Since $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}^{b}\left(u(\mathbf{x})\right)$, it follows from II.2 that $U(\mathbf{x}^{\prime})_{i}$ is obtained from $U(\mathbf{x})_{i}$ by two deletions. Suppose that $U(\mathbf{x})_{1}$ is $d$-regular and $\eta\left(U(\mathbf{x})_{1}\right)$ is given. According to Lemma IV.2, we can efficiently recover $U(\mathbf{x})_{1}$. Then by Lemma IV.3, the two deletions in $U(\mathbf{x})_{1}$ can be approximately located. In Lemma IV.5 below, we will show that this helps to locate errors in $u(\mathbf{x})$. To explain the idea, more notations are needed.
For $1\leq i\leq b$ and $1\leq j\leq n/b$, let $U(\mathbf{x})_{i,j}$ be the $j$-th coordinate of $U(\mathbf{x})_{i}$. By the relationship between $u(\mathbf{x})$ and $U(\mathbf{x})$, it is easy to see that $U(\mathbf{x})_{1,j}=u(\mathbf{x})_{1+(j-1)b}$. This justifies the following definition
$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\}.$
Note that $I_{n,b}$ is a subset of $[n]$. Clearly, we have $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
Example IV.3
Let $n=10$, $b=2$ and $\mathbf{x}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}2\in\Sigma_{3}^{10}$. Then $u(\mathbf{x})={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}1$. By definition, we have $I_{n,b}=\{1,3,5,7,9\}$ and $u(\mathbf{x})_{I_{n,b}}=01100$. On the other hand, the matrix representations of $u(\mathbf{x})$ is
$U(\mathbf{x})=\begin{pmatrix}0&1&1&0&0\\ 0&1&0&0&1\end{pmatrix}$.
It is easy to see that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
Recall that $r\left(U(\mathbf{x})_{1}\right)$ is the number of runs in $U(\mathbf{x})_{1}$. For $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$, let $U(\mathbf{x})_{1,I_{j}}$ be the $j$-th run of $U(\mathbf{x})_{1}$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$. By definition, $I_{j}$ is an interval in $[n/b]$. For all $j\geq 1$, we associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$. In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$. In other words, $I_{j}^{L}$ is the interval of length at most b1b-1 in $[n]$ to the left of $I_{j}^{\prime}$. Note that $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$ are dependent on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
Example IV.4
Let $\mathbf{x}$ be the sequence in Example IV.3. There are three runs in $U(\mathbf{x})_{1}=01100$: 0, $11$ and $00$. Recall that $p_{j}$ indicates where the $j$-th run ends. Then we have $p_{1}=1$, $p_{2}=3$, $p_{3}=5$, $I_{1}=[1,1]$, $I_{2}=[2,3]$ and $I_{3}=[4,5]$. All intervals $I_{j}^{\prime}$ and $I_{j}^{L}$ are listed in the following:
$\begin{array}[]{cc}I_{1}^{L}=\emptyset,I_{1}^{\prime}=[1,2],\\ I_{2}^{L}=[2,2],I_{2}^{\prime}=[3,6],\\ I_{3}^{L}=[6,6],I_{3}^{\prime}=[7,10].\end{array}$
We briefly explain why we define intervals $I_{j}^{\prime}$ and $I_{j}^{L}$. A $b$-burst-deletion in $u(\mathbf{x})$ induces one deletion in each row of $U(\mathbf{x})$. If it is known that the deletion in the first row $U(\mathbf{x})_{1}$ occurs in the interval $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, we can only locate the deletion in the $i$-th row in the interval $\left[p_{j-1},p_{j}\right]$ when $i\geq 2$ (see II.2). Note that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. When looking at $u(\mathbf{x})$, interval $I_{j}^{L}$ (if not empty) corresponds to the last b1b-1 coordinates in the $p_{j-1}$-th column of $U(\mathbf{x})$ and interval $I_{j}^{\prime}$ corresponds to columns of $U(\mathbf{x})$ indexed by $I_{j}$. As a result, the $b$-burst-deletion in $u(\mathbf{x})$ can be located in the interval $I_{j}^{L}\cup I_{j}^{\prime}$. See Figure 1 for an illustration.
[…]
(a) $U(\mathbf{x})$: $I_{j}$ in $\left[n/b\right]$
[…]
(b) $\mathbf{x}$: $I_{j}^{\prime}$ and $I_{j}^{L}$ in $\left[n\right]$
Figure 1: Illustration of relationship between $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$.
In the rest of this section, let $A(\mathbf{x})=A(\mathbf{x},b)$ be the matrix representation of $\mathbf{x}$ and $A(\mathbf{x})_{1}$ be the first row of $A(\mathbf{x})$.
Lemma IV.5
Let $q\geq 2$, $b>1$ and $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. That is to say, $\mathbf{x}^{\prime}$ is obtained from $\mathbf{x}$ by a 22-$b$-burst-deletion. Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular. Let $\eta(\cdot)$ be the function defined in Lemma IV.2. Then given $\eta\left(U(\mathbf{x})_{1}\right)$, we can
  1. $(1)$
    either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$;
  2. $(2)$
    or find an interval $J\subseteq[n]$ of length at most $3bd\log(n/b)+b-1$, such that $D_{1},D_{2}\subseteq J$.
Proof:
Since $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, we have $u(\mathbf{x}^{\prime})=u(\mathbf{x})_{[n]\setminus(D_{1}\cup D_{2})}$. It follows that $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$. Suppose that $U(\mathbf{x}^{\prime})_{1}$ is obtained from $U(\mathbf{x})_{1}$ by deleting $U(\mathbf{x})_{1,k_{1}}$ and $U(\mathbf{x})_{1,k_{2}}$. Recall that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. This implies that $U(\mathbf{x}^{\prime})_{1}$ is obtained from $u(\mathbf{x})_{I_{n,b}}$ by deleting $u(\mathbf{x})_{1+(k_{1}-1)b}$ and $u(\mathbf{x})_{1+(k_{2}-1)b}$.
Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.2, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$. Then we can know the number of runs in $U(\mathbf{x})_{1}$ and $I_{j}$ for all $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$. In addition, by Lemma IV.3, we can
  • either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $k_{1}\in I_{j_{1}}$ and $k_{2}\in I_{j_{2}}$;
  • or find an interval $c,d]\subseteq[n/b]$ of length at most $3d\log(n/b)$ such that $k_{1},k_{2}\in[c,d]$.
Similar to the discussion in the paragraph above Lemma IV.5, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. For the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$. Now the proof is completed. ∎
From now on, we assume that $\mathbf{x}\in\Sigma_{q}^{n}$ and $A(\mathbf{x})_{1}$ is $d$-regular, where $d$ will be specified later. Let $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$, where $D_{1},D_{2}\subseteq[n]$ are two unknown disjoint intervals of length $b$. Suppose that $\eta\left(U(\mathbf{x})_{1}\right)$ is given. The first step of decoding $\mathbf{x}$ from $\mathbf{x}^{\prime}$ is to decode $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$. After this step, we can know the number of runs in $U(\mathbf{x})_{1}$ and $I_{j}$ for all $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$. Then by Lemma IV.5, one can approximately locate the two bursts in $\mathbf{x}$ if $\eta\left(U(\mathbf{x})_{1}\right)$ is given. Based on the two cases in Lemma IV.5, we split our discussion into Section IV-B and Section IV-C. Our codes for correcting two $b$-burst-deletions are presented in Theorem IV.1, whose proof is divided into Lemmas IV.6 and IV.7. In Section IV-B and Section IV-C, we follow notations defined in this subsection.
Before proceeding, we need the following observation.
Observation IV.1
Let $m>b$. Suppose $\mathbf{u}\in\Sigma_{q}^{m}$ and $\mathbf{v}$ is obtained from $\mathbf{u}$ by single $b$-burst-deletion. Then for any $b\leq i\leq m-b$, we have $\mathbf{v}_{[1,i-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[1,i]}\right)$ and $\mathbf{v}_{[i+1,m-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[i+1,m]}\right)$.
IV-B When Case (1) in Lemma IV.5 happens
In this case, we already found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Note that $I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}=\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]$ and $I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}=\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]$. Then it follows that
$x_{k}=\begin{cases}x^{\prime}_{k},\text{ if }k<p_{j_{1}-1}b-b+2,\\ x^{\prime}_{k-b},\text{ if }p_{j_{1}}b<k<p_{j_{2}-1}b-b+2,\\ x^{\prime}_{k-2b},\text{ if }p_{j_{2}}b<k\leq n.\end{cases}$ (13)
In particular, $\mathbf{x}_{I_{j}^{\prime}}$ is known to us for any $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, to decode $\mathbf{x}$ from $\mathbf{x}^{\prime}$, it is sufficient to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
In the rest of this subsection, it is always assumed that $j_{2}-j_{1}\geq 2$, since otherwise, we have $\left|I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\cup I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\right|\leq 2bd\log(n/b)+b-1$, and therefore, this situation is covered by case (2) in Lemma IV.5 (or equivalently, case (2) in Lemma IV.3).
Since $j_{2}-j_{1}\geq 2$, deletions in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ do not affect $\mathbf{x}_{I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}}$. As a result, we have $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}\right)$ (i.e., $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$ is obtained from $\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}$ by single $b$-burst-deletion) and $\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}}\right)$ due to the fact that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\subseteq I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\subseteq I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}$. Applying IV.1 to $\mathbf{u}=\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}$ (i.e., the concatenation of $\mathbf{x}_{I_{j_{1}-1}}$ and $\mathbf{x}_{I_{j_{1}}}$), $\mathbf{v}=\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$, $m=\left(p_{j_{1}}-p_{j_{1}-2}\right)b$ and $i=\left(p_{j_{1}-1}-p_{j_{1}-2}\right)b$, we conclude that
$\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}-1}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right),$ (14)
$\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right).$
Similarly, we have
$\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}-1}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right),$ (15)
$\mathbf{x}^{\prime}_{\left[(p_{j_{2}-1}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right).$
Recall that the range of $j_{1}$ is $\left[1,r(U(\mathbf{x})_{1})\right]$. To make (14) valid when $j_{1}=1$, we let $\mathbf{x}_{I_{0}^{\prime}}=0^{b}$ (the sequence consisting of $b$ symbols 0) and $\mathbf{x}^{\prime}_{\left[p_{-1}b+1,-b\right]}$ be an empty sequence.
By (14) and (15), to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$, we need to apply to each $\mathbf{x}_{I_{j}^{\prime}}$ a code for correcting single $b$-burst-deletion. Let $\phi(\cdot)$ be the function defined in Lemma II.2. Then by Lemma II.2, our goal boils down to recovering the values of $\phi\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $U(\mathbf{x})_{1}$ is $d$-regular and each $U(\mathbf{x})_{1,I_{j}}$ is a run of $U(\mathbf{x})_{1}$, we have $\left|I_{j}\right|\leq d\log(n/b)$ and hence $\left|I_{j}^{\prime}\right|\leq db\log(n/b)$ for all $0\leq j\leq r(U(\mathbf{x})_{1})$. Therefore, each $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in [0,N1][0,N-1], where $N=2^{\log\log(n)+O_{q,d,b}(1)}$. Here, $O_{q,d,b}(1)$ denotes a constant dependent only on $q,d$ and $b$.
Define $\overline{\phi}(\mathbf{x})=\phi\left(\mathbf{x}_{I_{0}^{\prime}}\right)\phi\left(\mathbf{x}_{I_{1}^{\prime}}\right)\cdots\phi\left(\mathbf{x}_{I_{r(U(\mathbf{x})_{1})}^{\prime}}\right)\in[0,N-1]^{r(U(\mathbf{x})_{1})+1}$. According to (13), when given $\mathbf{x}^{\prime}$, the value of $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, sequence $\overline{\phi}(\mathbf{x})$ suffered two bursts (of length two) of erasures. To correct errors in $\overline{\phi}(\mathbf{x})$, let $\varphi(\cdot)$ be the function defined in Lemma II.6 and define
$f(\mathbf{x})=\varphi\left(\overline{\phi}(\mathbf{x})\right).$
Recall that we have $r(U(\mathbf{x})_{1})\leq n/b$. Then By Lemma II.6, $f(\mathbf{x})$ can be viewed as an integer in $0,N_{1}-1]$, where $N_{1}=2^{\log n+4\log\log n+O_{q,d,b}(1)}$.
The following lemma follows from the above analysis, Lemmas II.2 and II.6.
Lemma IV.6
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f(\mathbf{x})$, we can efficiently recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
IV-C When Case (2) in Lemma IV.5 happens
Recall that case (2) in Lemma IV.5 corresponds to case (2) in Lemma IV.3. Therefore, for case (2), instead of looking at $\mathbf{x}$ and $\mathbf{x}^{\prime}$, we look at $A(\mathbf{x})$ and $A(\mathbf{x}^{\prime})$. In this case, we have already decoded $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ by using $\eta\left(U(\mathbf{x})_{1}\right)$. Furthermore, by case (2) in Lemma IV.3, we found an interval $c_{1},c_{2}]\subseteq[n/b]$ of length at most $3d\log(n/b)$, which contains the positions of the two deletions in $U(\mathbf{x})_{1}$ (and hence in $A(\mathbf{x})_{1}$). By II.2, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $c_{1}-1,c_{2}]$, for all $2\leq i\leq b$. Let $A(\mathbf{x})_{i,k}$ be the $k$-th coordinate in the $A(\mathbf{x})_{i}$. Then $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin[c_{1}-1,c_{2}]$. In fact, it holds that $A(\mathbf{x})_{i,k}=A(\mathbf{x}^{\prime})_{i,k}$ when $1\leq k<c_{1}-1$ and $A(\mathbf{x})_{i,k}=A(\mathbf{x}^{\prime})_{i,k-2}$ when $c_{2}<k\leq n/b$.
Let $P=3d\log(n/b)+1$. Define
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/bP\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/bP\right\rceil-1.\end{cases}$
Since $c_{1}$ and $c_{2}$ are known and $\left|[c_{1}-1,c_{2}]\right|\leq 3d\log(n/b)+1$, we can find some $j_{0}$ such that $c_{1}-1,c_{2}]\subseteq J_{j_{0}}$. Suppose $J_{j_{0}}=[d_{1},d_{2}]$. According to above discussion, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin J_{j_{0}}$. To decode $A(\mathbf{x})_{i}$ from $A(\mathbf{x}^{\prime})_{i}$, it remains to recover $A(\mathbf{x})_{i,J_{j_{0}}}$. It is easy to see that $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$ is obtained from $A(\mathbf{x})_{i,J_{j_{0}}}$ by two deletions. Therefore, we can apply Lemma II.4 to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$.
Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. For each $1\leq j\leq\left\lceil n/bP\right\rceil-1$, let function $g_{j}(\mathbf{x})$ be defined in (16).
$g_{j}\left(\mathbf{x}\right)=\begin{cases}\left(\xi_{1}\left(A(\mathbf{x})_{1,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q>2;\\ \left(\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q=2.\end{cases}$ (16)
Note that in the above definition of $g_{j}(\mathbf{x})$, we distinguish the two cases where $q>2$ and $q=2$. This is because when $q=2$, we have $A(\mathbf{x})=U(\mathbf{x})$, and hence the first row of $A(\mathbf{x})$ was already recovered.
Since $\left|J_{j}\right|\leq 2P=6d\log\left(n/b\right)+2$ for all $j$, by Lemma II.4, each $\xi_{1}\left(A(\mathbf{x})_{i,J_{j}}\right)$ is a binary vector of length $7\left\lceil\log q\right\rceil\log\log n$ $O_{q,d,b}(1)$. Therefore, each $g_{j}(\mathbf{x})$ is a binary vector of length at most $7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)$ (when $q>2$) or $7(b-1)\log\log n+O_{d,b}(1)$ (when $q=2$). Therefore, we can view $g_{j}(\mathbf{x})$ as an integer in $\left[0,N_{2}-1\right]$, where
[…]
For $s\in\{0,1\}$, let $L_{s}=\left\{1\leq j\leq\left\lceil n/bP\right\rceil-1:j\equiv s\pmod{2}\right\}$. For $s\in\{0,1\}$, define
$h^{(s)}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{s}}g_{j}(\mathbf{x})\pmod{N_{2}}$. (18)
Lemma IV.7
For case (2) in Lemma IV.5 (and hence case (2) in Lemma IV.3), given $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$, we can uniquely recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $\mathbf{x}^{\prime}$, for all $i$. In particular, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
Proof:
Without loss of generality, assume $j_{0}\in L_{0}$. Then $g_{j}(\mathbf{x})$ is known to us for each $j\in L_{0}\setminus\{j_{0}\}$. By above discussion and Lemma II.4, to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ for all $2\leq i\leq b$, it is sufficient to obtain the value of $g_{j_{0}}(\mathbf{x})$. It follows from Equation 18, that
$g_{j_{0}}(\mathbf{x})\equiv\left(h^{(0)}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}g_{j}(\mathbf{x})\right)\pmod{N_{2}}$.
Since $0\leq g_{j_{0}}(\mathbf{x})<N_{2}$, we have
$g_{j_{0}}(\mathbf{x})=\left(h^{(0)}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}g_{j}(\mathbf{x})\right)\pmod{N_{2}}$.
Now the proof is completed. ∎
We are ready to present our main result in this section.
Theorem IV.1
Suppose $q\geq 2$, $n>2b$ and $b>1$. Let $N_{1}$, $N_{2}$, $f(\mathbf{x})$, $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$ be defined as above. Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$. For any $0\leq a<N_{0}$, $0\leq b<N_{1}$ and $0\leq c_{0},c_{1}<N_{2}$, define the code $\mathcal{C}_{1}$ as
$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ h^{(0)}(\mathbf{x})=c_{0},h^{(1)}(\mathbf{x})=c_{1}\end{array}\right\}.$
Then $\mathcal{C}_{1}$ is a 22-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O_{q,b}(1)$:
  • $q>2$ is even, $d=7$ and $n\geq 9$;
  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
If $q=2$, $d=7$ and $n\geq 9$, the redundancy is upper bounded by $5\log n+14b\log\log n+O_{b}(1)$.
Proof:
Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma IV.5, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $3bd\log(n/b)+b-1$, such that $D_{1},D_{2}\subseteq J$. For the former case, Lemma IV.6 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. For the latter case, Lemma IV.7 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. Therefore, $\mathcal{C}_{1}$ can correct two $b$-burst-deletions.
By Lemma IV.4, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $A(\mathbf{x})_{1}$ is $d$-regular. It then follows from the pigeonhole principle that there are some $a$, $b$, $c_{0}$, and $c_{1}$, such that
$\left|\mathcal{C}_{1}\right|\geq\frac{q^{n-1}}{N_{0}\cdot N_{1}\cdot(N_{2})^{2}},$
which implies by definition that
$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+\log N_{1}+2\log N_{2}+1$
$=5\log n+14\log\log n+2\log N_{2}+O_{q,b}(1).$
Now the claim for redundancy follows from (17). ∎
V Non-binary Two-deletion Correcting Codes
In this section, it is always assumed that $q>2$. Recall that $\mathcal{B}_{t}(\mathbf{x})$ is the set of sequences obtained from $\mathbf{x}$ by $t$ deletions. We aim to give a new construction of $q$-ary two-deletion correcting codes. Recall the definition of $u(\mathbf{x})$ for any $\mathbf{x}\in\Sigma_{q}^{n}$. If $\mathbf{x}^{\prime}\in\mathcal{B}_{2}(\mathbf{x})$, then $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}(u(\mathbf{x}))$. So if $u(\mathbf{x})$ belongs to a binary two-deletion correcting code, we can recover $u(\mathbf{x})$ from $u(\mathbf{x}^{\prime})$. We further assume that $\mathbf{x}$ (and thus $u(\mathbf{x})$) is $d$-regular. Then Lemma IV.3 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions. This enables us to approximately determine error positions in $\mathbf{x}$. For the latter case, we need a $q$-ary two-deletion correcting code, which is given in Lemma II.4. For the former case, we need a $q$-ary single-deletion correcting code. We use the code given in [23].
Lemma V.1
[23] Suppose $q>2$ and $n>2$. There is a function $\mathsf{DVT}:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+\log q}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\mathsf{DVT}(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$.
Let $u(\mathbf{x})_{I_{j}}$ ($1\leq j\leq r\left(u(\mathbf{x})\right)$) be all the runs in $u(\mathbf{x})$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(u(\mathbf{x})\right)}=n$ and $p_{j-1}<p_{j}$ for all $j$. Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$. Therefore, by Lemma V.1, $\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$. Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime. By the following lemma, we have $n\leq Q<2n$.
Lemma V.2 (Bertrand–Chebyshev theorem)
For every integer $n\geq 2$, there is always at least one prime $p$ such that $n\leq p<2n$.
Define
$f_{0}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{2N_{1}}$, (19)
$f_{1}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{Q}$.
The two functions $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$ can help to deal with case (1) in Lemma IV.3, as shown in the next lemma.
Lemma V.3
Suppose that the two deletions in $\mathbf{x}$ occurred in two known intervals $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$, respectively. Then given $f_{0}(\mathbf{x})$ and $f_{1}(\mathbf{x})$, we can uniquely recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
Proof:
Notice that $x_{k}$ is known to us for all $k\notin I_{j_{1}}\cup I_{j_{2}}$. It remains to recover $\mathbf{x}_{I_{j_{1}}}$ and $\mathbf{x}_{I_{j_{2}}}$. It is easy to see that $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}+1,p_{j_{1}}-1\right]}\in\mathcal{B}_{1}\left(\mathbf{x}_{I_{j_{1}}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2},p_{j_{2}}-2\right]}\in\mathcal{B}_{1}\left(\mathbf{x}_{I_{j_{2}}}\right)$. By Lemma V.1, to recover $\mathbf{x}_{I_{j_{1}}}$ and $\mathbf{x}_{I_{j_{2}}}$, it is sufficient to know the values of $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)$ and $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)$. Note that $\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)$ is known to us for all $j\neq j_{1},j_{2}$. Let $\delta_{0}=\left(f_{0}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\right)\pmod{2N_{1}}$ and $\delta_{1}=\left(f_{1}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\right)\pmod{Q}$. Then it follows from Equation 19 that
$\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{0}\pmod{2N_{1}}$, (20)
$j_{1}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+j_{2}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{1}\pmod{Q}$. (21)
Since $0\leq\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)<2N_{1}$ and $0\leq\delta_{0}<2N_{1}$, we can conclude from Equation 20 that $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)=\delta_{0}$. Combining this with Equation 21, we get
$\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{0}\pmod{Q}$, (22)
$j_{1}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)+j_{2}\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\equiv\delta_{1}\pmod{Q}$.
Since $1\leq j_{1}<j_{2}\leq n\leq Q$, we have $j_{1}\not\equiv j_{2}\pmod{Q}$. Therefore, system (22) has a unique solution in the field $\mathbb{F}_{Q}$: $\left(\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)\pmod{Q},\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\pmod{Q}\right)$. Since $0\leq\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right),\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)<N_{1}\leq Q$, we have $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)=\mathsf{DVT}\left(\mathbf{x}_{I_{j_{1}}}\right)\pmod{Q}$ and $\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)=\mathsf{DVT}\left(\mathbf{x}_{I_{j_{2}}}\right)\pmod{Q}$. Now the proof is completed. ∎
For case (2) in Lemma IV.3, we follow similar idea in Section IV-C. Let $P=3d\log n$ and
$J_{j}=\begin{cases}\left[(j-1)P+1,(j+1)P\right],\mbox{ if }1\leq j\leq\left\lceil n/P\right\rceil-2,\\ \left[(j-1)P+1,n\right],\mbox{ if }j=\left\lceil n/P\right\rceil-1.\end{cases}$
Let $\xi_{1}(\cdot)$ be the function in Lemma II.4, which can help to correct two deletions in $q$-ary sequences. Since $\left|J_{1}\right|\leq 2P=6d\log n$, we can view each $\xi_{1}\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{7\left\lceil\log q\right\rceil\log(6d\log n)+O_{q}(1)}$ (see Lemma II.4). For $s\in\{0,1\}$, let $K_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$ and define
$h^{(s)}(\mathbf{x})=\mathop{\sum}\limits_{j\in K_{s}}\xi_{1}\left(\mathbf{x}_{J_{j}}\right)\pmod{N_{2}}$.\\ (23)
Lemma V.4
Suppose we are in case (2) in Lemma IV.3. Given $h^{(0)}$ and $h^{(1)}$, we can uniquely recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
The proof is similar to that of Lemma IV.7, and thus omitted.
Our $q$-ary two-deletion correcting code is given in the next theorem.
Theorem V.1
Suppose $q>2$. Let $N_{1},N_{2},Q$, $f_{0}(\mathbf{x})$, $f_{1}(\mathbf{x})$, $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{0}<2N_{1}$, $0\leq b_{1}<Q$, and $0\leq c_{0},c_{1}<N_{2}$, define the code $\mathcal{C}_{2}$ as
$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}\text{ is }d\text{-regular},\\ \eta(u(\mathbf{x}))=a,\\ f_{s}(\mathbf{x})=b_{s}\text{ for }s\in\{0,1\},\\ h^{(0)}(\mathbf{x})=c_{0},h^{(1)}(\mathbf{x})=c_{1}\end{array}\right\}.$
Then $\mathcal{C}_{2}$ is a two-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there are some $a$, $b_{0}$, $b_{1}$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
  • $q$ is even, $d=7$ and $n\geq 9$;
  • $q$ is odd, $d=10$ and $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
The proof of Theorem V.1 follows directly from Lemmas V.3 and V.4. The claim for redundancy follows from Lemma IV.4 and Lemma V.2.
VI Conclusion
In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $5\log n+O(\log\log n)$ for all $b\geq 2$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique. Inspired by these ideas, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
In this work, it is required that the two bursts have the same length. Allowing the two deletion-bursts to have different sizes would make the error model more general and applicable to a wider range of practical scenarios. Indeed, such a generalization would be more realistic, as burst deletions in real-world applications (e.g., communication systems or storage systems) may not always have equal lengths. The case where the two bursts have different lengths seems more involved. Extending our work to handle bursts of different sizes is a promising direction for future research.
A more complex problem is to construct codes capable of correcting at most two bursts of deletions, where each burst has a length at most $b$. Our technique fails in these setup since we can not even know the number of bursts or the length of each burst occurring in a received sequence. This problem is also left for future research.
Acknowledgement
The authors express their gratitude to the anonymous reviewers for their detailed and constructive comments which are very helpful to the improvement of the presentation of this paper. The authors would also like to thank Prof. Alberto Ravagnani, the associate editor, for his excellent editorial job.
Appendix A A Lemma
The following Lemma A.1 is derived by slightly modifying the proof of [43, Proposition 3.3.7]. Before showing its proof, we list some facts. Let $\ln(\cdot)$ be the natural logarithm. Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}$-$\cdots$.
Fact A.1
For all $0\leq x<1$, we have $\ln(1-x)\leq-x-\frac{x^{2}}{2}$ and $\ln(1+x)\leq-x+\frac{x^{2}}{2}$. When $0\leq x\leq\frac{1}{2}$, we have $\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
Proof:
When $0\leq x<1$, we have $\ln(1-x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{(-x)^{i}}{i}=-\sum_{i=1}^{\infty}\frac{x^{i}}{i}\leq-x-\frac{x^{2}}{2}$. Since $\ln(1+x)=-x+\frac{x^{2}}{2}-\left(\ln(1+x)-x+\frac{x^{2}}{2}\right)$ and $\ln(1+x)-x+\frac{x^{2}}{2}\geq 0$ for all $x\geq 0$, we have $\ln(1+x)\leq-x+\frac{x^{2}}{2}$. At last, it is easy to see that $\ln(1-x)=x+\frac{x^{2}}{2}+x^{3}\left(\frac{1}{3}+\frac{x}{4}+\frac{x^{2}}{5}+\cdots\right)\leq x+\frac{x^{2}}{2}+x^{3}(1+x+x^{2}+\cdots)=x+\frac{x^{2}}{2}+\frac{x^{3}}{1-x}$. When $x\leq\frac{1}{2}$, we have $1-x\geq\frac{1}{2}$ and thus $\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$. ∎
Lemma A.1
Suppose that $q\geq 2$ is fixed and $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. Then it holds that
$H_{q}(1-\frac{1}{q}-\epsilon)\leq 1-\frac{\epsilon^{2}}{4}.$
Proof:
It is easy to verify that
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ (24)
$=$ $-\left(1-\frac{1}{q}-\epsilon\right)\log_{q}\left(\frac{1-1/q-\epsilon}{q-1}\right)$
$\quad-\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1}{q}+\epsilon\right)$
$=$ $-\log_{q}\left(\frac{1}{q}\left(1-\frac{\epsilon q}{q-1}\right)\right)$
$\quad+\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)$
$=$ $1-\frac{1}{\ln q}\left[\ln\left(1-\frac{\epsilon q}{q-1}\right)\right$.
$\quad\left.-\left(\frac{1}{q}+\epsilon\right)\ln\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)\right]$
$=$ $1+\frac{1}{\ln q}\left[-\ln\left(1-\frac{\epsilon q}{q-1}\right)+\left(\frac{1}{q}+\epsilon\right)\ln\left(1-\frac{\epsilon q}{q-1}\right)\right$.
$\quad\quad\quad\quad\quad\left.-\left(\frac{1}{q}+\epsilon\right)\ln\left(1+\epsilon q\right)\right].$
Since $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q}\right\}$, we have $0<\frac{\epsilon q}{q-1}<\frac{1}{2}$ and $0<\epsilon q<1$. Then it follows from A.1 that $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}}$, $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}$ and $\ln(1+\epsilon q)\leq-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}$. Now by Equation 24, we have
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$
$\leq$ $1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}\right$.
$\quad\quad\left.+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}\right)\right]$
$=$ $1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}\right$.
$\quad\quad\quad\quad\quad\quad\quad\left.+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q^{2}}{q-1}+\frac{\epsilon^{2}q^{3}(q-2)}{2(q-1)^{2}}\right)\right]$
$=$ $1+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}$
$\quad\quad\quad\quad\quad+\frac{1}{\ln q}\left[\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\frac{\epsilon^{2}q^{2}}{q-1}+\frac{\epsilon^{2}q^{2}(q-2)}{2(q-1)^{2}}\right]$
$=$ $1-\frac{\epsilon^{2}q^{2}}{2(q-1)\ln q}+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}$
$\overset{(a)}{\leq}$ $1-\frac{\epsilon^{2}}{4},$
where (a) follows from the fact that $\epsilon<\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)$. ∎
Appendix B Encoding sequences into regular sequences
In this section, we give two methods to encode a sequence into a $d$-regular sequence using only one redundant symbol. The first method is an instantiation of the encoding idea in the proof of [19, Lemma 11]. The second method relies on the sequence replacement technique. The first method works for more flexible parameters $n$ and $d$. The time complexity of the second method is lower.
For any $q\geq 2$, let $Q\subseteq\Sigma_{q}^{m}$ be the set of sequences containing two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$. According to [19, Lemma 10], it holds that $\left|Q\right|\geq 2^{m}-2\times(1.62)^{m+2}$ when $q=2$. By the definition of $q$-ary $d$-regular sequences, it is clear that $\left|Q\right|\geq q^{m}-2\times(1.62)^{2}\times(0.81q)^{m}$ for any even $q$. In Lemma IV.4, it is shown that $\left|Q\right|\geq q^{m}-2\times(0.87q)^{m}$.
B-A The first method
Suppose that $d$ and $n$ satisfy conditions in Lemma IV.4. In this subsection, let $m=\left\lfloor\frac{d}{2}\log n\right\rfloor$, $k=\left\lfloor n/m\right\rfloor$ and
$\mathcal{R}^{\prime}=\bigg{\{}\mathbf{x}=\mathbf{x}^{(0)}\cdots\mathbf{x}^{(k-1)}$ $\mathbf{x}^{(k)}\in\Sigma_{q}^{n}:$
$\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 0\leq i\leq k-1\\ \mathbf{x}^{(k)}\in\Sigma_{q}^{n-km}\end{array}\bigg{\}}.$
Then in the proof of [19, Lemma 11] and Lemma IV.4, it is shown that each sequence in $d$-regular and $\left|\mathcal{R}^{\prime}\right|\geq q^{n-1}$. We aim to encode a sequence of length n1n-1 into $\mathcal{R}^{\prime}$. Before showing the algorithm, we need to define some functions.
For $\mathbf{x}\in\Sigma_{q}^{n-1}$, define $f_{\textup{dec}}(\mathbf{x})=\sum_{i=1}^{n-1}x_{i}q^{i-1}$. It is clear that $f_{\textup{dec}}(\mathbf{x})\in\left[0,q^{n-1}-1\right]$ and $f_{\textup{dec}}(\mathbf{x})$ can be computed in $O(n)$ time.
We further assume that $n$ is sufficiently large such that $m\geq 8$ when $q$ is odd, or such that $m\geq 12$ when $q$ is even. Under this assumption, it is easy to verify that $\left|Q\right|\geq q^{m-1}\geq q^{n-km}$. Since $\left|Q\right|^{k}-1+\left|Q\right|^{k}\left(q^{n-km}-1\right)=\left|\mathcal{R}^{\prime}\right|-1\geq q^{n-1}-1$, each integer $a$ in $\left[0,q^{n-1}-1\right]$ can be uniquely written as $a=\sum_{i=0}^{k}a_{i}\left|Q\right|^{i}$, where $0\leq a_{0},a_{1},\ldots,a_{k-1}<\left|Q\right|$ and $0\leq a_{k}<q^{n-km}\leq\left|Q\right|$. Denote $g_{Q}(a)=\left(a_{0},\ldots,a_{k-1},a_{k}\right)$. The time complexity of computing $g_{Q}(a)$ is $O\left(\log_{\left|Q\right|}\left(q^{n-1}-1\right)\right)=O\left(\log\left(q^{n-1}-1\right)/\log\left|Q\right|\right)=O\left(n/m\right)$.
Since $\left|Q\right|\leq q^{m}=O\left(n^{d\log q/2}\right)$, we can efficiently construct the set $Q$ by brute-force. There is a bijection between the set $Q$ and the integer set $\left[0,\left|Q\right|-1\right]$. We can build a lookup table which gives such a bijection. Then for each sequence in $Q$ we can find its corresponding integer in $\left[0,\left|Q\right|-1\right]$ in $O\left(n^{d\log q/2}\right)$ time. Similarly, there is a bijection between $\Sigma_{q}^{n-km}$ and the set $\left[0,q^{n-km}-1\right]$ and for each sequence in $\Sigma_{q}^{n-km}$, we can find its corresponding integer in $\left[0,q^{n-km}-1\right]$ in $O\left(n^{d\log q/2}\right)$ time by a lookup table.
Our encoding algorithm is described as follows. For a sequence $\mathbf{x}\in\Sigma_{q}^{n-1}$, compute $f_{\textup{dec}}(\mathbf{x})$ and $g_{Q}\left(f_{\textup{dec}}(\mathbf{x})\right)$. Suppose that $g_{Q}\left(f_{\textup{dec}}(\mathbf{x})\right)=\left(a_{0},\ldots,a_{k-1},a_{k}\right)$, where $0\leq a_{0},\ldots,a_{k-1}<\left|Q\right|$ and $0\leq a_{k}<q^{n-km}$. As analyzed above, we can map $a_{i}$ to a sequence $\mathbf{y}^{(i)}$ in $Q$ for each $0\leq i\leq k-1$, and map $a_{k}$ to a sequence $\mathbf{y}^{(k+1)}$ in $\Sigma_{q}^{n-km}$. At last, concatenate these sequences and we get the output $\mathbf{y}=\mathbf{y}^{(0)}\cdots\mathbf{y}^{(k-1)}\mathbf{y}^{(k)}$, which is $d$-regular. Clearly, the time complexity of the encoding process is $O\left((k+1)n^{d\log q/2}\right)=O\left(n^{d\log q/2+1}/m\right)$.
Decoding $\mathbf{y}$ to the original sequence is straightforward. And the time complexity is also $O\left(n^{d\log q/2+1}/m\right)$.
B-B The second method
In this subsection, we give a new encoding algorithm based on the sequence replacement technique. This technique has been widely used in the literature to encode a sequence to a sequence without forbidden substrings [29, 44, 36, 38].
Let $m=\left\lfloor\frac{d}{3}\log n\right\rfloor$ (where $d$ will be specified later) and $Q$ be the set defined at the beginning of this section. Then we have that $\left|\Sigma_{q}^{m}\setminus Q\right|\leq(c\cdot q)^{m}$ for some $0<c<1$. Therefore, there is an injective mapping $f_{Q}$ from $\Sigma_{q}^{m}\setminus Q$ to the set $\Sigma_{q}^{\left\lceil m\left(1+\log_{q}c\right)\right\rceil}$. Then the inverse mapping $f_{Q}^{-1}:\left\{f_{Q}:\mathbf{z}\in\Sigma_{q}^{m}\setminus Q\right\}\rightarrow\Sigma_{q}^{m}\setminus Q$ is well-defined. In addition, by building a lookup table, both $f_{Q}$ and $f_{Q}^{-1}$ can be computed in $O\left((cq)^{m}\right)=O\left(n^{\frac{d}{3}\log(cq)}\right)$ time.
In the rest of this subsection, denote $k=\left\lfloor n/m\right\rfloor$. Fix a suitable $d$ such that when $n$ is sufficiently large, we have $m\geq\left\lceil m\left(1+\log_{q}c\right)\right\rceil+\left\lceil\log_{q}k\right\rceil+4$.
For a sequence $\mathbf{x}\in\Sigma_{q}^{n-1}$, we first append a symbol 0 to $\mathbf{x}$ and get the sequence $\tilde{\mathbf{x}}=\mathbf{x}0$. This appended symbol 0 will help to indicate when our decoding algorithm ends. Next, we partition sequence $\tilde{\mathbf{x}}$ as $\tilde{\mathbf{x}}=\tilde{\mathbf{x}}^{(1)}\cdots\tilde{\mathbf{x}}^{(k)}\tilde{\mathbf{x}}^{(k+1)}$, where $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}$ for each $1\leq i\leq k$ and $\tilde{\mathbf{x}}^{(k+1)}\in\Sigma_{q}^{n-km}$. For any symbol $\alpha\in\Sigma_{q}$ and integer $\ell\geq 1$, let $\alpha^{\ell}$ denote the sequence consisting of $\ell$ symbols $\alpha$. For simpler notations, denote $m_{1}=\left\lfloor\frac{m-\left\lceil m\left(1+\log_{q}c\right)\right\rceil-\left\lceil\log_{q}k\right\rceil}{2}\right\rfloor$ and $m_{2}=\left\lceil\frac{m-\left\lceil m\left(1+\log_{q}c\right)\right\rceil-\left\lceil\log_{q}k\right\rceil}{2}\right\rceil$.
The idea behind the encoding algorithm is as follows.
  1. Step 1
    For each $1\leq i\leq k$, check if $\tilde{\mathbf{x}}^{(i)}\in Q$. If not, delete $\tilde{\mathbf{x}}^{(i)}$ from $\tilde{\mathbf{x}}$ and still denote the resulted sequence by $\tilde{\mathbf{x}}$.
  2. Step 2
    Then we append the sequence $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$ to the end of $\tilde{\mathbf{x}}$. Here, we use $b_{q}(i)$ to denote the $q$-ary representation of integer $i$. Clearly, $b_{q}(i)\in\Sigma_{q}^{\left\lceil\log_{q}k\right\rceil}$.
  3. Step 3
    Note that the substring $\tilde{\mathbf{x}}^{(k+1)}$ will never be deleted after Step 1 and Step 2. Continue Step 1 and Step 2 until the left of $\tilde{\mathbf{x}}^{(k+1)}$ is empty or all $\tilde{\mathbf{x}}^{(i)}$ to the left of $\tilde{\mathbf{x}}^{(k+1)}$ is in $Q$.
Since $m\geq\left\lceil m\left(1+\log_{q}c\right)\right\rceil+\left\lceil\log_{q}k\right\rceil+4$, we have $m_{2}\geq m_{1}\geq 2$. Therefore, once Step 2 is executed, a sequence in $Q$ is appended to the end. Note that the length of $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$ is $m$. This implies that the length of $\tilde{\mathbf{x}}$ does not change after Step 2. Denote the sequence at the end of this algorithm by $\mathbf{y}$. The fact that $m=\left\lfloor\frac{d}{3}\log n\right\rfloor$ implies $d\log n>3m-3$. Then it is easy to see that any substring of $\mathbf{y}$ of length at least $d\log n$ either contains a $\tilde{\mathbf{x}}^{(i)}$ (which is on the left of $\tilde{\mathbf{x}}^{(k+1)}$) or contains a substring of the form $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$ (which is on the right of $\tilde{\mathbf{x}}^{(k+1)}$). Therefore, $\mathbf{y}$ is $d$-regular.
The formal algorithm is given in Algorithm 1. Note that in Algorithm 1, it always holds that $\tilde{\mathbf{x}}^{(j+1)}=\tilde{\mathbf{x}}^{(k+1)}$. In each loop, if $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}\setminus Q$, a block of length $m$ to the left of $\tilde{\mathbf{x}}^{(j+1)}$ is deleted. Therefore, in Line 6, the substring between $\tilde{\mathbf{x}}^{(j+1)}$ and $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$ is $\tilde{\mathbf{x}}_{\left[n-(k-j)m+1,n\right]}$, which the suffix of $\tilde{\mathbf{x}}$ (in previous loop) of length (kj)m(k-j)m. There are at most $k$ loops. In each loop, if $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}\setminus Q$, we have to compute $f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)$ and $b_{q}(i)$. Therefore, the overall time complexity is $O\left(kn^{\frac{d}{3}\log(cq)}\right)=O\left(n^{\frac{d}{3}\log(cq)+1}/m\right)$.
Denote the encoder based on Algorithm 1 by $\textup{Enc}(\cdot)$. We present the decoding algorithm in Algorithm 2, where the function $\textup{Dec}(\cdot)$ computes the decimal representation of an input sequence. In the Initialization step of Algorithm 1, a symbol 0 is appended to $\mathbf{x}$. In Line 6 of Algorithm 1, the appended block ends with symbol q1q-1. Therefore, in Algorithm 2, the fact that $x_{n}=q-1$ implies that there is an appended block which has not been deleted and the decoder has to go into another while loop. Suppose that $\mathbf{x}=\tilde{\mathbf{x}}^{(1)}\cdots\tilde{\mathbf{x}}^{(i-1)}\tilde{\mathbf{x}}^{(i+1)}\cdots$ $\tilde{\mathbf{x}}^{(j)}\tilde{\mathbf{x}}^{(j+1)}\tilde{\mathbf{x}}_{\left[n-(k-j)m+1,n\right]}f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}.$ Then it is clear that $\mathbf{x}_{\left[n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil+1,n-m_{1}-m_{2}\right]}=b_{q}(i)$ and $\mathbf{x}_{\left[n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil-\left\lceil m\left(1+\log_{q}c\right)\right\rceil+1,n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil\right]}$. Therefore, we get the index $i$ in Line 4 and $f_{Q}^{-1}$ is the deleted $\tilde{\mathbf{x}}^{(i)}$. When the while loop ends, we get the sequence $\mathbf{x}0$. In Line 8, the last symbol 0 is deleted. Now the correctness of Algorithm 2 is clear. There are at most $k$ while loops and in each loop, we have to compute $\textup{Dec}(\cdot)$ and $f_{Q}^{-1}(\mathbf{u})$. Therefore, the overall time complexity is $O\left(kn^{\frac{d}{3}\log(cq)}\right)=O\left(n^{\frac{d}{3}\log(cq)+1}/m\right)$.
Input: $\mathbf{x}\in\Sigma_{q}^{n-1}$
Output: $\tilde{\mathbf{x}}\in\mathcal{R}_{q,n,d}$
1 Initialization
2 $\tilde{\mathbf{x}}\leftarrow\mathbf{x}0$, write $\tilde{\mathbf{x}}=\tilde{\mathbf{x}}^{(1)}\cdots\tilde{\mathbf{x}}^{(k)}\tilde{\mathbf{x}}^{(k+1)}$, where $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}$ for each $1\leq i\leq k$ and $\tilde{\mathbf{x}}^{(k+1)}\in\Sigma_{q}^{n-km}$
3 $i\leftarrow 1$, $j\leftarrow k$
4 while $i\leq j$ do
5  if $\tilde{\mathbf{x}}^{(i)}\in\Sigma_{q}^{m}\setminus Q$ then
6     $\tilde{\mathbf{x}}\leftarrow\tilde{\mathbf{x}}^{(1)}\cdots\tilde{\mathbf{x}}^{(i-1)}\tilde{\mathbf{x}}^{(i+1)}\cdots\tilde{\mathbf{x}}^{(j)}\tilde{\mathbf{x}}^{(j+1)}\tilde{\mathbf{x}}_{\left[n-(k-j)m+1,n\right]}f_{Q}\left(\tilde{\mathbf{x}}^{(i)}\right)b_{q}(i)0^{m_{1}}(q-1)^{m_{2}}$
7     $j\leftarrow j-1$
8 else
9     $i\leftarrow i+1$
10  end if
11
12 end while
return $\tilde{\mathbf{x}}$
Algorithm 1 Encoding a sequence into a $d$-regular sequence
Input: $\tilde{\mathbf{x}}=\textup{Enc}(\mathbf{x})\in\Sigma_{q}^{n-1}$
Output: $\mathbf{x}\in\Sigma_{q}^{n-1}$
1 Initialization
2 $\mathbf{x}\leftarrow\tilde{\mathbf{x}}$
3 while $x_{n}=q-1$ do
4  $i_{0}\leftarrow\textup{Dec}\left(\mathbf{x}_{\left[n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil+1,n-m_{1}-m_{2}\right]}\right)$
5  $\mathbf{u}\leftarrow\mathbf{x}_{\left[n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil-\left\lceil m\left(1+\log_{q}c\right)\right\rceil+1,n-m_{1}-m_{2}-\left\lceil\log_{q}k\right\rceil\right]}$
6  $\mathbf{x}\leftarrow\mathbf{x}_{\left[1,(i_{0}-1)m\right]}f_{Q}^{-1}\left(\mathbf{u}\right)\mathbf{x}_{\left[(i_{0}-1)m+1,n-m\right]}$
7 end while
[…]
return $\mathbf{x}$
Algorithm 2 Decoding a $d$-regular sequence into the original sequence